### v5

# Geometric interpretation of quantitative instability

Abstract.
Given a real algebraic group action on a linear space $G\lefttorightarrow V$, a vector $v\in V$ is called unstable if $0\in\overline{Gv}-Gv$, where the closure is taken with respect to the Zariski topology. A fundamental theorem of Kempf [21] in geometric invariant theory states that $v$ is unstable if and only if there is a one-parameter subgroup $A$ of $G$ such that $v$ is unstable with respect to it, i.e., $0\in\overline{Av}-Av$. Assuming $G$ is a semisimple real algebraic group defined over $\mathbb{Q}$, we give a new proof to this result using a geometric interpretation of the setting. In the process, we also give a new proof of an effective version of this result by Shah and Yang [29, Prop. 3.2]. Our interpretation involves relating the length of vectors under a linear action to convex functions on certain $\operatorname{CAT}(0)$-spaces, and bound the latter from below by Busemann functions.
1. Introduction
Geometric invariant theory studies the quotient of a variety $V$ by a real algebraic group action $G\lefttorightarrow V$. An important phenomenon occurs when an orbit closure of a point $v$ in $V$ contains a point which is not in the $v$-orbit, i.e, when $\overline{Gv}-Gv\neq\emptyset$, where the the topology considered is the Zariski topology. It is then a natural question to study the behavior of the orbit around such point. In [21] Kempf studied the following question:
Given $v\in V$, $u\in\overline{Gv}-Gv$, for which curves $\{\gamma(t)\}$ in $G$ does the distance $\|u-\gamma(t)v\|$ shrink ‘fastest’?
The phenomenon of non-closed $G$-orbits is called instability. In homogeneous dynamic the $\{0\}$-instability, i.e., the case $u$ is the zero vector in a $G$-representation $V$, is widely used when studying how far an orbit is into a cusp, see [23, 18, 28, 30]. Recently, Kempf’s result has been used in the study of orbits and more specifically orbit closures in homogeneous dynamics. First by Yang [38], which used it to prove not only nondivergence, but also equidistribution of certain measures. Yang’s usage of [21] is extending Shah’s basic lemma [28, Proposition 4.2], which was proved independently. In [17, 29] the method was pushed further, and a quantitative version of Kempf’s result was proved, answering a seemingly different question:
Assuming $0\in\overline{Gv}$, can the size $\left\|gv\right\|$ for all $g\in G$ be controlled?
The solution they provide is using highest weight representations:
Theorem 1.1 ([29, Prop. 3.2]).
Let $\varrho:G\to\operatorname{GL}(V)$ be a $\mathbb{Q}$-representation and $0\neq v\in V(\mathbb{Q})$ be an unstable vector, i.e., $0\in\overline{Gv}-Gv$. Then, there exists a $\mathbb{Q}$-highest weight representation $\tilde{\varrho}:G\rightarrow\operatorname{GL}(W)$ a highest weight vector $w\in W(\mathbb{Q})$, and $a,c>0$ such that such that for all $g\in G$
(1.1) $\log\|\varrho(g)v\|\geq a\log\|\tilde{\varrho}(g)w\|-c.$
This theorem is important for their setting, but it has another application previously unknown, as it answers a question of Weiss, [37, Question 3.4] positively. In this paper we will prove Theorem 1.1 using geometric tools, and give a geometric interpretation to the right hand side of Eq. (1.1).
1.1. Main results
1.1.1. Geometric interpretation
For the left hand side of Eq. (1.1), we analyze the symmetric space of non-positive curvature $M:=K\backslash G$, where $K$ is the maximal compact subgroup of $G$. For every real representation $\varrho:G\rightarrow\operatorname{GL}(V)$, with $V$ equipped with a $K$-invariant norm, the function $g\mapsto\log\|\varrho(g)v\|$ descends to a function $f_{v}:M\to\mathbb{R}$. With the right choice of norm, $f_{v}$ is convex, see Observation 5.3 and Lemma 2.11 for the existence of such a norm. The right hand side of Eq. (1.1) can also be viewed as a multiple of $f_{w}$. However, since $w\in W$ is a highest weight vector, it has an intrinsic description, $f_{w}$ is a constant multiple of a Busemann function on $M$ (See Definition 3.1). We use geometric tools to show the following:
Theorem 1.2.
Let $\varrho:G\rightarrow V$ be a $\mathbb{R}$-representation and assume a $K$-invariant norm on $V$. Then, for every $v\in V$ there exists a Busemann function $\beta$ and constants $a,c>0$ so that we have the following:
  1. (1)
    Any $x\in M$ satisfies
    $f_{v}(x)\geq a\beta(x)-c.$
  2. (2)
    If $v\in V$ is a highest weight vector, then for any $x\in M$
    $f_{v}(x)=a\beta(x)-c.$
Remark 1.3.
In this manuscript we only deal with the real and rational setting (see §6 for the rational setting), but similar results should hold for a general field, using more general theory, see [34, 3].
1.1.2. Algebraic properties
Theorem 1.2 lacks the rational nature of Theorem 1.1. To this end we give various algebraic descriptions for ‘rational’ Busemann functions. The equivalent descriptions for Busemann functions can be found at Theorem 3.5 and the equivalent description for “rational” Busemann functions can be found in Theorem 6.7. Each of these theorems has some equivalent description, and in particular contains the second part of Theorem 1.2. The first is the standard definition of the Busemann function, in Theorem 3.5 and a restriction of the geodesic in Theorem 6.7. The second uses homomorphisms from a parabolic subgroup. The third uses a linear combination of functions of the form $f_{v}$, for $v$ in a special collection of representations, called fundamental representations, (see Definition 2.16). Theorem 6.7 has a forth equivalent description, which is the right hand side of Eq. (1.1).
1.2. On the proof
Analyzing Busemann functions: In §3 we show the connection between the Busemann functions and the fundamental representations. In particular, we prove Theorem 3.5 which implies Theorem 1.2 Part (2).
Convex geometry: In §§4-5 we prove Theorem 5.1 which implies Theorem 1.2 Part (1). We investigate the geodesic paths in $G$ that ‘shrink’ $v$ the most, and prove Theorem 1.2 Part (1). It has two key steps. In §4 we find a ‘fastest shrinking geodesic’ for all ‘nice’ convex functions. In particular, we are able to find the geodesic that shrinks a given vector in a representation the fastest. This part is a geometric analogous to Kempf’s result [21].
In §5 we show that the function defined by the left hand side of (1.1) is indeed ‘nice’, and use some of its algebraic properties to show that it can be bounded below by a (specific) Busemann function. Theorem 5.1 is the main theorem of this part. This part provides a geometric way to view [21], and a good intuition for Theorem 5.1.
Algebraic view point: In the previous part we have constructed a geometric object, the Busemann function, (which is also algebraic by Theorem 6.7) given algebraic data, a vector in an algebraic representation. Next, we wish to say that since the vector is defined over $\mathbb{Q}$, it is invariant under $\operatorname{Gal}(\mathbb{C}/\mathbb{Q})$, hence so does the Busemann function, i.e., it is also defined over $\mathbb{Q}$. This requires us to find an algebraic description for the fastest shrinking geodesic of a vector. Such an algebraic description was defined and proved by Kempf in [21]. In §6 we relate our constructions to Kempf’s construction, and use his result to deduce that our ‘fastest shrinking geodesic’ is indeed algebraic. Thus, the inequality in (1.1) is satisfied for the rational fundamental representations (and not only the real ones).
In §7 we give an alternative proof of Theorem 1.1 using a more precise version of Theorem 1.2.
1.3. Further research
Theorem 5.1 gives a lower bound using a Busemann function for every convex function on a symmetric space of noncompact type $f:M\to\mathbb{R}$ coming from a vector in a representation. This yields the following conjecture:
Conjecture 1.4.
For every convex function $f$ on $M$, there is a Busemann function $\beta:M\to\mathbb{R}$, $a>0,C\in\mathbb{R}$ such that for every $x\in M$
$f(x)>a\beta(x)+C.$
The anlagous conjecture for trees is false: Let $G$ be a leafless $d$-regular tree, $x_{0}\in G$ a vertex and $(x_{i})_{i=0}^{\infty}$ an infinite ray. Let $f:G\to\mathbb{R}$ be the function defined as follows: $f(x_{i})=-\sqrt{i}$ for $i\geq 0$, and for other $x\in G$ let $x_{i_{0}}$ be the closest element to $x$ on the ray. Define $f(x):=-\sqrt{i_{0}}+d(x,x_{i_{0}})(\sqrt{i+1}-\sqrt{i})$. One can verify that it is convex and not bounded from below by any multiple of a Busemann function.
Failures to extend this example to symmetric spaces leads us to believe the conjecture.
Acknowledgment
We would like to thank B. Weiss for suggesting this problem and for familiarizing us with [21]. We would also like to thank P. Yang for bringing our attention to the new results in this topic. The first author would like to thank N. Tur, O. Bojan, and A. Levit for helpful discussions. The second author would also like to thank R. Spatzier for his interest in this project and for several helpful discussions and comments. The first autor is supported by ERC grant HomDyn, ID 833423.
2. Preliminaries
2.1. Homomorphisms
we will use several types of homomorphisms, Algebraic ones and topological ones. The set of continuous homomorphisms will be denoted by $\operatorname{Hom}$. The set of $k$-algebraic homomorphisms will be denoted by $\operatorname{Hom}_{k}$.
2.2. Hadamard spaces
A Hadamard space is defined to be a nonempty complete metric space $(M,d)$ such that, given any points $x,y\in M$, there exists $m\in M$ such that for every $z\in M$ we have
(2.1) $d(z,m)^{2}+{\frac{d(x,y)^{2}}{4}}\leq{\frac{d(z,x)^{2}+d(z,y)^{2}}{2}}.$
The point $m$ is called the midpoint of $x$ and $y$, and it satisfies $d(x,m)=d(y,m)=d(x,y)/2$. we study the behavior of convex functions in such spaces. The main properties of them which are of use to us are listed in Lemma 2.2 below.
Alternatively, a space is Hadamard if it is a complete $\operatorname{CAT}(0)$-space. A metric space $(M,d)$ is a $\operatorname{CAT}(0)$-space if it is geodesic (as defined below) and every ‘small enough’ geodesic triangle satisfies a certain inequality. Such spaces were first defined and studied by Gromov, see [15]. For a more detailed discussion on complete and $\operatorname{CAT}(0)$-spaces see [5].
Definition 2.1 (Geodesic).
Let $(M,d)$ be a metric space. Given $x,y\in M$, a geodesic from $x$ to $y$ is a map $\gamma:I\rightarrow Y$, where $I=[a,b]\subset\mathbb{R}$ is a closed interval, such that $\gamma(a)=x$, $\gamma(b)=y$ and
$d(\gamma(s),\gamma(s^{\prime}))=|s-s^{\prime}|$
for all $s,s^{\prime}\in I$ (in particular, $d(x,y)=b-a$). By abuse of notations, we identify geodesics with their images. A geodesic ray in $M$ is a map $\gamma:[0,\infty)\rightarrow X$ such that $d(\gamma(s),\gamma(s^{\prime}))=|s-s^{\prime}|$ for all $s,s^{\prime}\geq 0$. $(X,d)$ is said to be a geodesic space if every two points in $M$ are joined by a geodesic.
Lemma 2.2.
Let $(M,d)$ be a Hadamard space. Let $\gamma_{1},\gamma_{2}:[0,\infty)\to M$ be geodesic rays such that $\gamma_{1}(0)=\gamma_{2}(0)$. Then,
  1. (1)
    for every $s>t>0$ we have
    $\frac{d(\gamma_{1}(s),\gamma_{2}(s))}{s}\geq\frac{d(\gamma_{1}(t),\gamma_{2}(t))}{t}.$
  2. (2)
    (in particular) for $r:=d(\gamma_{1}(1),\gamma_{2}(1))$ and all $t\geq 1$, we have
    $d(\gamma_{1}(t),\gamma_{2}(t))\geq rt.$
  3. (3)
    the midpoint $m_{t}$ of $\gamma_{1}(t)$ and $\gamma_{2}(t)$ satisfies
    $d(\gamma_{1}(0),m_{t})\leq\sqrt{1-r^{2}/4}\cdot t.$
Proof.
According to [5, Prop. 2.2] the distance function in a $\operatorname{CAT}(0)$-space is convex, i.e., for any $s>t>0$
$d(\gamma_{1}(t),\gamma_{2}(t))$ $\leq\frac{t}{s}d(\gamma_{1}(s),\gamma_{2}(s))+\left(1-\frac{t}{s}\right)d(\gamma_{1}(0),\gamma_{2}(0)).$
Since we assume $\gamma_{1}(0)=\gamma_{2}(0)$, Claim (1) follows. Claim (2) follows directly from Claim (1). By (2.1) and Claim (2) we have
$d(\gamma_{1}(0),m_{t})$ $\leq\sqrt{d(\gamma_{1}(0),\gamma_{1}(t))^{2}/2+d(\gamma_{2}(0),\gamma_{2}(t))^{2}/2-d(\gamma_{1}(t),\gamma_{2}(t))^{2}/4}$
$\leq\sqrt{t^{2}-(rt/2)^{2}},$
proving Claim (3). ∎
2.3. Symmetric spaces of non-compact type
There are many results on symmetric spaces of non-compact type, i.e., Riemannian manifolds of non-positive sectional curvature, whose group of symmetries contains an inversion symmetry about every point. In particular, such spaces are Hadamard spaces. Here we present some of their geometric properties as well as some explicit constructions for future use.
2.3.1. Overview
Let $G$ be an $\mathbb{R}$-reductive group, i.e., connected, linear, algebraic group over $\mathbb{R}$ with a trivial unipotent radical, and $K$ be a maximal compact subgroup of $G$. Let $M:=K\backslash G$, and $\pi:G\to M$ be the projection map. We can define a $G$-invariant Riemannian metric $d_{M}$ on $M$ (See Definition 2.8), which give rise to a metric in the standard way. Then, $(M,d_{M})$ is a symmetric space of non-compact type.
We are also interested in a specific algerbraic subgroup, $A\subset G$, which is a maximal real split torus, i.e., $A\cong(\mathbb{R}^{\times})^{r}$, were $r$ is the real rank of $G$. The subgroup $A$ is called the Cartan torus and have certain compatibility conditions with $K$, see Definition 2.9.
The metric structure on $M$ has some phenomenal properties, which we now describe.
2.3.2. Metric properties
Geodesics in symmetric spaces can be generalized into the higher dimensional concept of flats:
Definition 2.3 (Flats, geometric description).
A map $a:\mathbb{R}^{k}\to M$ is called flat if the restricted metric on $\mathbb{R}^{k}$ is the Euclidean metric. We abuse notations and do not distinguish between $a$ and its image $\operatorname{Im}(a)$. A flat is called maximal if there is no flat which contains it. In particular, in our setting, the dimension of all maximal flats of $M$ is equal to the $\mathbb{R}$-rank of $G$.
Remark 2.4.
Note that it follows from the definition that for any semisimple abelian connected subgroup $T\subset G$, the subgroup $\pi(T)$ is a flat, and vise versa.
Theorem 2.5 (Algebraic description of flats, [5, Remark 10.60(5)]).
For every Cartan torus $A$, all maximal flats can be described as translations of $\pi(A)$. In particular, for every flat $a:\mathbb{R}^{k}\to M$ there is $g\in G$ and a continuous homomorphism $\iota:\mathbb{R}^{k}\to A$ such that $a(x)=\pi(\iota(x)g)$.
Theorem 2.5 also implies an explicit description of the geodesics in $M$, as can be seen in the next chapter.
Claim 2.6.
The group $K$ acts transitively on maximal flats containing $\pi(e)$. Moreover, for every geodesic $\gamma$ containing $\pi(e)$ the subgroup $\operatorname{\mathrm{stab}}_{K}(\gamma)$ acts transitively on the set of maximal flats which contain $\gamma$.
Proof.
The group $G$ acts transitively on the maximal flats by [5, Remark 10.60]. Moreover, since $A$ acts transitively on $\pi(A)$, $G$ acts transitively on pairs $(p,F)$ of a point $p$ in a maximal flat $F$. Since $K$ is the stabilizer of $\pi(e)$, we deduce that it acts transitively on the set of maximal flats containing $\pi(e)$.
As for the second part, assume that $\gamma(t)=\pi(\exp(\mathtt{a}t))$ for some $\mathtt{a}\in\mathfrak{a}$. Note that
$G^{\prime}:=\operatorname{Stab}_{G}(\mathtt{a})=\{g\in G:Ad_{g}(\mathtt{a})=\mathtt{a}\},$
is a reductive subgroup, as defined in [5, Definition 10.56]. We now remark that the first part of the proof, as well as the construction of the symmetric space in [5, Theorem 10.58], follow for reductive subgroups, and not only semisimple groups. As defined in [5, Theorem 10.58], the symmetric space $M^{\prime}$ of $G^{\prime}$ is a subspace of $M$, which contains all maximal flats of $M$ which contains $\gamma$. The second part of the claim now follows from the first part, applied to $M^{\prime}$. ∎
2.3.3. Explicit construction
We follow standard notation and results, see [11, §2] and [16, §IV]. Let $G$ be a semisimple Lie group, , i.e., the Killing form on $\mathfrak{g}=\operatorname{Lie}(G)$ is nondegenerate. Let $K$ be a maximal compact subgroup of $G$. Then, $M:=K\backslash G$ is a manifold. Fixing $o=[K]\in M$ which is stabilized by $K$, define a projection $\pi:G\rightarrow M$ which is given by
$\pi(g)=og,$
for any $g\in G$. Then, $G$ acts on $M$ by right multiplication.
Definition 2.7 (Cartan decomposition of $\mathfrak{g}$).
Recall that we fixed a maximal compact subgroup $K\subseteq G$. Denote its Lie algebra by $\mathfrak{k}=\operatorname{Lie}(K)$. The Killing form $B_{\mathfrak{g}}(\cdot,\cdot)$ is negative definite on $\mathfrak{k}$, and positive definite on its orthogonal complement $\mathfrak{p}=\mathfrak{k}^{\perp}$.
The decomposition $\mathfrak{g}=\mathfrak{k}\oplus\mathfrak{p}$ is the Cartan decomposition of $\mathfrak{g}$. Note that the adjoint action of $K$ on $\mathfrak{g}$ preserves $\mathfrak{k}$ and $B_{\mathfrak{g}}$, and hence preserves $\mathfrak{p}$ as well.
Definition 2.8 (Metric on $M$).
Fixing $o=[K]\in M$ which is stabilized by $K$, note that $T_{o}M=\mathfrak{g}/\mathfrak{k}\cong\mathfrak{p}$, thus define the positive definite bilinear form $B_{o}$ on $T_{o}M$ to be the the restriction of $B_{\mathfrak{g}}$ to $\mathfrak{p}$. Since that $B$ is $K$ invariant, we may use the $G$ action on $M$ and define a Riemannian metric $B_{p}$ on $T_{p}M$ for every $p\in M$. As usual in Riemannian geometry, for every curve segment $\gamma:[0,1]\rightarrow M$ we define the arc length of $\gamma$ by
[…]
and the metric
$d_{M}(p,q):=\inf\{L(\gamma):\gamma:[0,1]\to M,\gamma(0)=p,\gamma(1)=q\}.$
Definition 2.9 (Cartan decomposition of $G$).
Fix a maximal abelian $\mathfrak{a}\subseteq\mathfrak{p}$. Then $A=\exp\mathfrak{a}$ is a maximal split torus in $G$, and one may write (see for example [20, Thm 7.39])
(2.3) $G=KAK.$
The next result follows directly from Theorem 2.5 (see also [11, 2.4.2]).
Corollary 2.10.
All geodesics starting at $\pi(e)$ are of the form $t\mapsto\pi(\exp(t\mathtt{a})g)$ for some $\mathtt{a}\in\mathfrak{a}$ with $B(\mathtt{a},\mathtt{a})=1$ and $g\in K$. Alternatively, such geodesic is of the form $t\mapsto\pi(\exp(t\mathtt{p}))$ for some $\mathtt{p}\in\mathfrak{p}$ with $B(\mathtt{p},\mathtt{p})=1$.
2.3.4. Example: $G=\operatorname{SL}_{n}(\mathbb{R})$
Let $G=\operatorname{SL}_{n}(\mathbb{R})$ and $K=\operatorname{SO}(n)$. Then, $K$ is the maximal compact subgroup of $G$, and $K\backslash G$ is isomorphic to the space of positive-definite symmetric $n\times n$-matrices, with determinant $1$, denoted by $P(n,\mathbb{R})$. $G$ acts on $P(n,\mathbb{R})$ by conjugation. Then, the identity matrix $I$ is stabilized by $K$.
The Cartan decomposition is defined by $\mathfrak{p},\mathfrak{k}$ being the symmetric and antisymmetric matrices in $\mathfrak{sl}_{n}=\operatorname{Lie}(G)$. The tangent space $T_{p}P(n,\mathbb{R})$ at a point $p\in P(n,\mathbb{R})$ is naturally isomorphic (via translation) to $\mathfrak{p}$, and the pseudo-Riemannian structure there is defined by
$B_{p}(Y,Z)=\operatorname{Tr}(p^{-1}Yp^{-1}Z).$
The maximal $\mathbb{R}$-split torus $A$ here is the set of diagonal matrices. For more information see [5, §II.10 The Symmetric Space $P(n,\mathbb{R})$].
2.4. Representations of $G$ and short vectors
We use the following standard notation of arithmetic groups (see [2, 3]). Fix a maximal $\mathbb{R}$-split torus $A$ in $G$, and denote its Lie-algebra by $\mathfrak{a}$. For an $\mathbb{R}$-representation $\varrho:G\rightarrow\operatorname{GL}(V)$ we denote by $\Phi_{\varrho}$ the set of $\mathbb{R}$-weights of $G$, i.e., the set of characters $\lambda\in\mathfrak{a}^{*}$ such that the subspace
$V_{\lambda}=\left\{v\in V:\text{for all }a\in A,\>\varrho(a)v=\lambda(a)v\right\}$
is not trivial. The space $V_{\lambda}$ is called the weight space corresponding to $\lambda$, and elements of $V_{\lambda}$ are called weight vectors corresponding to $\lambda$. Then, there is a decomposition
(2.4) $V=\bigoplus_{\lambda\in\Phi_{\varrho}}V_{\lambda}.$
The following lemma defines a ‘nice’ quadratic form on $V$ which we use to define a norm on it. It is proved in the Appendix.
Lemma 2.11 (Construction of bilinear form).
[…]
2.5. Parabolic subgroups and their properties
Fix a subfield $k\subseteq\mathbb{R}$. We will frequently assume that $G$ is defined over $k$, to use some of the algebraic structure. The reader should note that the Cartan decomposition is not necessarily defined over $k$. We use standard notation and results about parabolic groups in symmetric spaces (see [3, §11], [11, §2.17], or [5, §II.10]). We also prove some properties of them to be used in later chapters.
The classical, algebraic, definition of a parabolic group is a closed subgroup $P$ of $G$ so that $G/P$ is a projective variety. In our view of $G$, as a group acting on a symmetric space, the following, more geometrical, definition of a parabolic group is more informative. See [3, Corollary 11.2] and [24, Proposition 2.6].
Definition 2.12 (Parabolic and Unipotent subgroups).
A subgroup $P$ is called parabolic if it is of the form
$P=P_{\mathtt{a}}:=\left\{g\in G:\lim_{t\rightarrow\infty}\exp(-t\mathtt{a})g\exp(t\mathtt{a})\text{ exists}\right\},$
where $\mathtt{a}\in\mathfrak{p}$. If $G$ is defined over $k$ then the group $P$ is called $k$-parabolic if it is defined over $k$ as an algebraic group. It may be a $k$-parabolic even if $\mathtt{a}$ is $k$-algebraic. There are many different $\mathtt{a}$ defining the same $P_{\mathtt{a}}$.
The unipotent radical of a parabolic group $P$ is the unipotent part of the intersection of all Borel subgroups which are contained in $P$. Explicitly, the unipotent radical of $P_{\mathtt{a}}$ is
$U_{\mathtt{a}}=\left\{g\in G:\lim_{t\rightarrow\infty}\exp(-t\mathtt{a})g\exp(t\mathtt{a})=e\right\}.$
The unipotent radical is always nilpotent, and and if $P$ is a $k$ parabolic then the unipotent radical is also defined over $k$. Moreover, the unipotent radical depends only on $P_{\mathtt{a}}$ and not on $\mathtt{a}$.
The following subset of the semisimple part of the Levi decomposition of $P_{\mathtt{a}}$ is of special interest for us
(2.5) $T_{\mathtt{a}}$ $:=\exp\{\operatorname{\mathrm{stab}}_{\mathfrak{g}}(\mathtt{a})\cap\mathfrak{p}\},$
where $\operatorname{\mathrm{stab}}_{\mathfrak{g}}(\mathtt{a})=\{\mathtt{b}\in\mathfrak{g}:[\mathtt{a},\mathtt{b}]=0\}$. In particular, $T_{\mathtt{a}}\subset\operatorname{\mathrm{stab}}_{G}(\mathtt{a})$. If $\mathtt{a}\in\mathfrak{a}$, then we may further decompose
$\operatorname{Lie}(T_{\mathtt{a}})=\mathfrak{a}\oplus\mathfrak{p}_{\mathtt{a}},$
where $\mathfrak{p}_{\mathtt{a}}$ is orthogonal to $\mathfrak{t}$, with respect to the Killing form (see, for example [11, Prop. 2.17.10 and Prop. 2.17.13] and [19, Prop. 8.1]). Let $K_{\mathtt{a}}:=K\cap P_{\mathtt{a}}$.
Lemma 2.13 (Generalized Iwasawa decomposition, [11, Prop. 2.17.5]).
For any $\mathtt{a}\in\mathfrak{p}$ we have $P_{\mathtt{a}}=K_{\mathtt{a}}\cdot T_{\mathtt{a}}\cdot U_{\mathtt{a}}$ and $G=K\cdot P_{\mathtt{a}}=K\cdot T_{\mathtt{a}}\cdot U_{\mathtt{a}}$. Moreover, in both equations the indicated decomposition is unique.
Remark 2.14 (translation to Eberlein’s notations).
$\mathtt{a},P_{\mathtt{a}},U_{\mathtt{a}},T_{\mathtt{a}},K_{\mathtt{a}}$ corresponds to $X,G_{X},N_{X},A_{X},K_{X}$ in Eberlein’s notoations.
From here until the end of this section $G$ is a $k$-algebraic group.
Definition 2.15 (The Root system and Weyl group).
Let $\mathfrak{a}_{k}$ be a maximal $k$-split torus of $\mathfrak{g}$. We denote by $\Phi_{k}$ the root system of $\mathfrak{a}_{k}$, i.e. the set of non-trivial eigenvalues with respect to the adjoint action of $\mathfrak{a}_{k}$ on $\mathfrak{g}$, and by $W(\Phi_{k})$ the Weyl group of $\Phi_{k}$, i.e. the subgroup generated by reflections through the hyperplanes orthogonal to elements in $\Phi_{k}$.
Recall that we denote $\mathfrak{a}=\mathfrak{a}_{\mathbb{R}}$. For any choice of a positive $\mathbb{R}$-Weyl chamber $\mathfrak{a}^{+}\subset\mathfrak{a}_{k}-\bigcap_{w\in W(\Phi_{\mathbb{R}})}\ker(w)$, there is a maximal nilpotent subgroup $N_{\mathfrak{a}^{+}}$ such that any $\mathtt{a}\in\mathfrak{a}^{+}$ satisfies $U_{\mathtt{a}}=N_{\mathfrak{a}^{+}}$. In particular, as a special case of Lemma 2.13 we have the decomposition
(2.6) $G=K\cdot A\cdot N_{\mathfrak{a}^{+}}.$
[…]
Definition 2.16 (The fundamental weights and representations).
See [4] for the definitions and claim below. Let $\Delta_{k}=\{\alpha_{1},\dots,\alpha_{d}\}$ be a $k$-simple system of $\Phi_{k}$. Fix an order on $\Delta_{k}$. Then, this order determines a partial order on the $k$-weight lattice $\Lambda_{k}:=\operatorname{span}_{\mathbb{Z}}\Delta_{k}$. Note that $\Lambda_{k}$ is the set of all possible $k$-weights of all $k$-representation. Moreover, for any element $\chi\in\operatorname{span}_{\mathbb{N}}\Delta_{k}$ there exists $m\in\mathbb{Z}$ and a $k$-representation such that $m\chi$ is the highest weight of it with respect to the above fixed partial order.
The $k$-fundamental weights are the elements $\chi_{1},\dots,\chi_{r}\in\operatorname{span}_{\mathbb{N}}\Delta_{k}$ which satisfy
(2.7) $\frac{2\langle\alpha_{i},\chi_{j}\rangle}{\langle\alpha_{i},\alpha_{i}\rangle}=m_{i}\delta_{i,j},$
where $\delta_{i,j}$ is Kronecker delta and $m_{i}$ is the smallest integer so that there exist a $k$-representation with a highest weight $\chi_{i}$. We denote this $k$-representation by $\varrho_{i}$ and fix $v_{i}$, a highest weight vector of it. Note that the stabilizer of $v_{i}$ under $\varrho_{i}$ is a maximal $k$-parabolic subgroup, which we denote by $P_{i}$. The intersection $\bigcap_{i=1}^{d}P_{i}$ is a minimal $k$-parabolic subgroup, conjugate to all other minimal $k$-parabolic, and $(P_{i})_{i=1}^{d}$ are all maximal parabolic subgroups containing it.
Definition 2.17 (Attaching a character to a torus element).
For any $\mathtt{a}\in\mathfrak{a}_{k}$ denote by $\chi_{\mathtt{a}}$ the character on $\mathfrak{a}_{k}$ which is defined by
$\mathtt{b}\mapsto\frac{\langle\mathtt{a},\mathtt{b}\rangle}{\langle\mathtt{a},\mathtt{a}\rangle}.$
Similarly, for any character $\chi\in\mathfrak{a}_{k}^{*}$ we denote by $\mathtt{a}_{\chi}$ the element in $\mathfrak{a}_{k}$ which satisfies $\chi=\chi_{\mathtt{a}_{\chi}}$.
Since $\{\chi_{1},\dots,\chi_{r}\}$ spans $\mathfrak{a}_{k}^{*}$, there exist $a_{1},\dots,a_{r}\in\mathbb{R}$ such that
(2.8) $\chi_{\mathtt{a}}=\sum_{j=1}^{r}a_{j}\chi_{j}.$
Moreover, fixing $\Delta_{k}$ determines a positive $k$-Weyl chamber by
$\mathfrak{a}_{k}^{+}:=\{\mathtt{a}:\forall\alpha\in\Delta_{k},\quad\alpha(\mathtt{a})\geq 0\}.$
For any $\mathtt{a}$ there exists a choice of $\Delta_{k}$ so that $\mathtt{a}\in\mathfrak{a}_{k}^{+}$. In that case, the coefficients $a_{1},\dots,a_{r}$ in (2.8) are non-negative.
3. The Busemann functions
In this section we use the notation and results of §2.5.
The main goal of this section is to show equivalent descriptions for Busemann functions. In particular, Theorem 1.2 follows from the last result of this section, Theorem 3.9.
Definition 3.1 (the Busemann function).
Let $d$ be a right-invariant Riemannian metric on $M$. Given a geodesic ray $\gamma:[0,\infty)\rightarrow M$, the function $\beta_{\gamma}:M\rightarrow\mathbb{R}$ defined by
(3.1) $\beta_{\gamma}(x)=\lim_{t\rightarrow\infty}(d(x,\gamma(t))-t)$
is called the Busemann function associated to $\gamma$ (see [5, §II.8.17]).
Example 3.2 (Busemann functions on euclidean spaces).
Let $u\in\mathbb{R}^{k}$ be a vector with $\|u\|=1$. Then, the Busemann function of the geodesic $t\mapsto tu$ is $v\mapsto-\left\langle u,v\right\rangle$.
Definition 3.3 (The modular function).
Given a field $k$ and a $k$-algebraic group $P$, denote by $\delta_{P}:P\to k^{\times}$ the $k$-modular function associated with $P$. That is, assuming $\mu_{P}$ is the right Haar measure on $P$, for every Borel set $S$ and $h\in H$, set
$\delta_{P}(h)=\frac{\mu_{P}(h^{-1}S)}{\mu_{P}(S)}.$
$\delta_{P}$ is a group homomorphism.
Definition 3.4 (Positive homomorphism).
Let $P\subset G$ be a parabolic subgroup. A homomorphism $f:P\to\mathbb{R}$ is called positive if it is a linear combination with positive coefficients of $\log\delta_{P^{\prime}}$ for maximal parabolic subgroups $P^{\prime}$ containing $P$. This definition has an $k$-algebraic version for subfields $k\subseteq\mathbb{R}$. If $G$ is a $k$-algebraic semisimple group and $P$ a $k$-parabolic subgroup in it, then a homomorphism $f:P(\mathbb{R})\to\mathbb{R}$ is called $k$-positive if it is a linear combination with positive coefficients of $\log\delta_{P^{\prime}}$ for $k$-maximal parabolic subgroups $P^{\prime}$ containing $P$.
The main goal of this section is to prove the following theorem.
Theorem 3.5.
The following three classes of functions are equal:
  1. (1)
    Busemann functions: The class of constant shifts of nonnegative multiples of Busemann functions.
  2. (2)
    Homomorphisms from parabolics: The class of functions which satisfies that $\pi(p)\mapsto\chi(p)$ for all $p\in P$,where $P$ is a parabolic subgroup of $G$, $\chi:P\to\mathbb{R}$ is a positive homomorphims.
  3. (3)
    Lengths of highest weight vectors in fundumental representations: The class of functions of the form
    $g\mapsto\sum_{j=1}^{k}\alpha_{j}\log\|\varrho_{j}(g)v_{j}\|+C,$
    where $\varrho_{j}$ are $\mathbb{R}$-fundamental representations and $v_{i}$ are the appropriate highest weight vectors, which correspond to maximal $\mathbb{R}$-parabolic groups, and are defined in §2.5, for some choice of an $\mathbb{R}$-simple system, the $\alpha_{j}$ are non-negative, and $C\in\mathbb{R}$.
For the proof of Theorem 3.5 we need a good description of the modular functions as well as of the kernel of the Busemann functions.
Lemma 3.6 ([25, Lem 1.2]).
Let $\gamma$ be a geodesic defined by $\mathtt{a}\in\mathfrak{a}$, i.e.
(3.2) $\gamma(t)=\pi(\exp(t\mathtt{a})),\quad t\in[0,\infty).$
Let $\mathfrak{a}^{+}$ be a Weyl chamber such that $\mathtt{a}\in\overline{\mathfrak{a}^{+}}$, $N$ be the nilpotent group defined by $\mathfrak{a}^{+}$, and $\mathfrak{a}^{\prime}$ be the orthogonal compliment of $\mathtt{a}$ in $\mathfrak{a}$. Then, the horosphere $\mathcal{H}_{\mathtt{a}}:=\beta_{\gamma}^{-1}(0)\subset M$ is given by
$\mathcal{H}_{\mathtt{a}}=\pi(\exp(\mathfrak{a}^{\prime})\times N).$
Lemma 3.7.
Let $1\leq j\leq d$. Then, for any $h\in P_{j}$ there exists $m_{j}\in\mathbb{N}$ such that
$\varrho_{j}(h)v_{j}=m_{j}\delta_{P_{j}}(h)v_{j}.$
Proof.
First, as stated in §2.5, $v_{j}$ is a highest weight vector in $V_{j}$. Thus, the line spanned by $v_{j}$ is the highest weight eigenspace of $\varrho_{j}$, and is stabilized by $P_{j}$. That is, for any $h\in P_{j}$, the point $\varrho_{j}(h)v_{j}$ is a constant multiple of $v_{j}$.
Second, the modular function of the Lie group $P_{j}$ is given by
$\delta_{P_{j}}(h)=|\det(\operatorname{Ad}_{P_{j}}(h))|,$
where $\operatorname{Ad}_{P_{j}}(h)$ is the adjoint action of $h$ on $\operatorname{Lie}(P_{j})$. Using a standard decomposition of parabolic groups (see [20, §V.7]), one can deduce that $|\det(\operatorname{Ad}_{P_{j}}(h))|$ is equal to the determinant of the adjoint action of $P_{j}$ on $\mathfrak{n}_{j}:=\bigoplus_{\langle\lambda,\chi_{j}\rangle>0}\mathfrak{g}_{\lambda}$. Let $u_{j}$ be the wedge product of a $k$ basis of $\mathfrak{n}_{j}$. Then, $u_{j}$ is a highest weight vector of a positive integer multiple of $\chi_{j}$. The claim now follows. ∎
Proof of Theorem 3.5.
First note that the equivalence between classes (2) and (3) follows directly from Lemma 3.7, and so it is enough to prove the equivalence between classes (1) and (3).
Let $\beta_{\gamma}$ be a Busemann function, i.e., there exist $\mathtt{a}\in\mathfrak{a}$ and $g\in K$ such that $\gamma$ is defined by $t\mapsto\pi(\exp(t\mathtt{a})g)$. By replacing $\mathfrak{a}$ with $\operatorname{Ad}(g)\mathfrak{a}$, another maximal $k$-split torus of $\mathfrak{g}$, we have that $\gamma$ is of the form (3.2) for $\mathtt{a}\in\mathfrak{a}$.
Fix an $\mathbb{R}$-simple system so that $\mathtt{a}\in\mathfrak{a}^{+}$. This determines a set of $\mathbb{R}$-fundamental representations $\varrho_{1},\dots,\varrho_{d}$. We wish to show that for any $g\in G$
(3.3) $\beta_{\gamma}(\pi(g))+C=\sum_{j=1}^{k}a_{j}\log\|\varrho_{j}(g)v_{j}\|,$
where $\{a_{j}\}_{1\leq j\leq k}$ are the constants in the decomposition (2.8) of $\chi_{\mathtt{a}}$, and $C:=-\sum_{j=1}^{k}a_{j}\log\|v_{j}\|$.
Let $N$ and $\mathfrak{a}^{\prime}$ be as in Lemma 3.6. Fix $g\in G$. By (2.6) we may decompose $g=kbn$, where $k\in K$, $n\in N$, and $b=\exp(s\mathtt{a}+\mathtt{b})$ for $s\in\mathbb{R}$, $\mathtt{b}\in\mathfrak{a}^{\prime}$. Then, by Lemma 3.6, the definition of the Busemann function, and since $N$ is stabilized by $A$,
$\beta_{\gamma}(\pi(g))$ $=\beta_{\gamma}(\pi(\exp(s\mathtt{a})))=s$
Since for any $1\leq j\leq r$ the norm on $V_{j}$ is invariant under $K$, and highest weight vectors are invariant under $N$, we also have
$\log\|\rho_{j}(g)v_{j}\|$ $=\log\|\rho_{j}(b)v_{j}\|$
$=\log\|e^{\chi_{j}(b)}v_{j}\|$
$=\chi_{j}(b)+\log\|v_{j}\|.$
Thus, using (2.8) and the definition of $\chi_{\mathtt{a}}$, we arrive at
$\sum_{j=1}^{k}a_{j}\log\|\rho_{j}(g)v_{j}\|$ $=\sum_{j=1}^{k}a_{j}\left(\chi_{j}(b)+\log\|v_{j}\|\right)$
$=\chi_{\mathtt{a}}(b)-C$
$=\frac{\langle\mathtt{a},\mathtt{b}\rangle}{\langle\mathtt{a},\mathtt{a}\rangle}-C$
$=s-C,$
as wanted.
Proving any function in class (3) is a constant shift of non-negative multiple of a Busemann function is similar to the above. ∎
Remark 3.8.
For a Busemann function $\beta_{\gamma}$ of a geodesic ray $\gamma:t\mapsto\pi(\exp(t\mathtt{a}))$, for some $\mathtt{a}\in\mathfrak{p}$, the parabolic subgroup in Class (2) is $P_{\mathtt{a}}$. Indeed, it follows from that proof of Theorem 3.5 that $\beta_{\gamma}$ satisfies (3.3) for the constants in the decomposition (2.8) of $\chi_{\mathtt{a}}$, and that $\chi_{\mathtt{a}}$ is the homomorphism which satisfies the description of $\beta_{\gamma}$ in Class (2). Now, one can see that $\chi_{\mathtt{a}}$ has is positive as a homomorphism from $P_{\mathtt{a}}=\bigcap_{a_{j}\neq 0}P_{j}$.
Theorem 3.9.
[…]
Proof.
Fix a maximal $\mathbb{R}$-split torus $A\subset P$ such that . Since $P$ stabilizes $\mathbb{R}v$, so does $A$, and hence one can deduce that $v$ is a highest weight vector for $\varrho$ with respect to the order implied by some minimal $\mathbb{R}$-parabolic subgroup of $P$. Denote by $\chi$ the weight of $v$. Note that we have $\chi=\sum_{i=1}^{d}\alpha_{i}\chi_{i}$ for some choice of positive $\alpha_{1},\dots,\alpha_{d}$. Since the homomorphisms of $P$ are in one-to-one correspondence with the characters of $A$, and $f$ agrees with $\chi$ on $A$, we may deduce the claim. ∎
4. The fastest shrinking geodesic
To construct the busemann function in Theorem 1.2 Part 1, we use the ‘fastest shrinking geodesic’ of our function of interest. In this section, the fastest shrinking geodesic is constructed for a class of functions in the more general setting of a $\operatorname{CAT}(0)$-space, and the properties of such geodesic are studied in a special case, which is relevant to us, and is used in the next chapter.
Recall the definition (and relevant notations) of a Hadamard space from §2.2. Fix a locally compact, Hadamard space $(M,d)$ and a point $o\in M$.
Definition 4.1 (Convex functions).
A function $f:M\to\mathbb{R}$ is convex if for every geodesic $\gamma:I\to M$ the composition $f\circ\gamma$ is a convex function on the interval $I$.
For $s>0$, $x\in M$ denote by $B(x,s)$ the closed ball of radius $s$ around the point $x$.
Lemma 4.2.
Let $f:M\rightarrow\mathbb{R}$ be a convex function which is unbounded from below. Then, For any $x\in M$, $s>0$, the function $f$ attains a minimum on the closed ball $B(x,s)$ at a unique point on its boundary $\{y\in M:d(y,x)=s\}$.
Proof.
First, assume by contradiction that $f$ attains a minimum on $B(x,s)$ at a point $y_{1}$ with $d(e,y_{1})<s$. In particular, there is a neighbourhood of $y_{1}$ which is contained in $B(x,s)$. Hence, $y_{1}$ is a local minimum of $f$. Since $f$ is convex, this implies that $y_{1}$ is a global minimum. A contradiction to the assumption that $f$ is unbounded from below.
Second, assume by contradiction that $f$ attains its minimum on $B(x,s)$ at two points $y_{1},y_{2}$. Since $M$ is a $CAT(0)$ space, it follows that the midpoint $y_{3}$ of $y_{1}$ and $y_{2}$ satisfies $d(y_{3},x)<s$. Since $f$ is convex it follows that $f(y_{3})\leq f(y_{2})=f(y_{1})$. This implies that $y_{3}$ is another points in which $f$ attains a minimum, with $d(y_{3},x)<s$. A contradiction to the previous discussion. ∎
Fix a convex, unbounded from below, function $f:M\to\mathbb{R}$.
Definition 4.3 (Minimizing points of $f$ in balls).
For any $s>0$ denote by $x_{s}$ the point which minimizes $f$ on $B(o,s)$. By Lemma 4.2 the point $x_{s}$ is uniquely defined and satisfy $d(o,x_{s})=s$. In particular, this definition implies $x_{0}=o$. Denote by $\gamma_{s}:[0,s]\to M$ the geodesic connecting $o$ and $x_{s}$.
Definition 4.4.
We say that $f$ is decreasing linearly if
$\liminf_{s\to\infty}\frac{f(x_{s})}{s}<0.$
Proposition 4.5.
If $f$ is decreasing linearly, then the sequence of geodesics $(\gamma_{s})_{s>0}$ constructed in Definition 4.3 converges to a geodesic $\gamma_{\infty}:[0,\infty)\to M$ pointwise. Moreover, $\gamma_{\infty}:[0,\infty)\to M$ is the unique such geodesic with $\gamma_{\infty}(0)=o$ and it satisfies
$f(\gamma_{\infty}(s))=-as\cdot(1+o(1)),$
for $a:=-\lim_{s\to\infty}\frac{f(x_{s})}{s}>0$.
Proof.
Let
(4.1) $a:=-\liminf_{s\to\infty}\frac{f(x_{s})}{s}.$
Since $f$ is decreasing linearly, we have $a>0$. Since the space of geodesic rays from $o$ is compact, the sequence $\{\gamma_{s}\}$ has a partial limit. Let $\gamma_{\infty}:[0,\infty)\to M$ be such a partial limit. Assume further that $\gamma_{\infty}$ is the limit of a subsequence $\{\gamma_{s_{i}}\}$ which satisfies $a=-\lim_{i\to\infty}\frac{f(x_{s_{i}})}{s_{i}}$.
Since $f$ is convex, for any $s>0$ we have
$f(\gamma_{\infty}(s))$ $=\lim_{i\to\infty}f(\gamma_{s_{i}}(s))$
$\leq\lim_{i\to\infty}\left(\frac{s_{i}-s}{s_{i}}f(x_{0})+\frac{s}{s_{i}}f(x_{s_{i}})\right)$
(4.2) $=f(x_{0})-s\cdot a,$
where the limit in the second line is taken over large enough $i$ so that $s_{i}>s$. Since $\gamma_{\infty}$ starts at $o$, we have that for any $s>0$, $d(\gamma_{\infty}(S),o)=s$. Hence, by Definition 4.3 for any $s>0$ we have
(4.3) $f(\gamma_{\infty}(s))\geq f(x_{s}).$
It now follows from (4.1) that for all $s>0$
$f(\gamma_{\infty}(s))=-as\cdot(1+o(1)).$
Note that (4.2) and (4.3) also imply
$a=\lim_{s\to\infty}-\frac{f(x_{s})}{s},$
i.e., $a$ is the limit of the sequence, and not only a partial limit as in (4.1). Thus, $f\circ\gamma_{\infty}$ of the claimed form.
Let $\gamma_{\infty}^{\prime}:[0,\infty)\to M$ be another geodesic ray with $\gamma_{\infty}^{\prime}(0)=o$ and $f(\gamma_{\infty}^{\prime}(s))=-as\cdot(1+o(1))$ for all $s\geq 0$. Denote by $r$ the distance at time one between the two geodesics, i.e., $r:=d_{M}(\gamma_{\infty}(1),\gamma_{\infty}^{\prime}(1))$. By Lemma 2.2(2), $d_{M}(\gamma_{\infty}(t),\gamma_{\infty}^{\prime}(t))\geq rt$ for every $t\geq 1$. Denote by $m_{t}$ the middle point between $\gamma_{\infty}(t)$ and $\gamma_{\infty}^{\prime}(t)$. Then, by Lemma 2.2(3), we have
(4.4) $d_{M}(x_{0},m_{t})\leq t\sqrt{1-r^{2}/4}.$
On the other hand, since $f$ is convex it follows that $f(m_{t})\leq at(1+o(1))$, and hence
$f(x_{t\sqrt{1-r^{2}/4}})\leq-at(1+o(1)).$
Unless $r=0$, this together with (4.4) contradicts the linear decrease of $f$. Hence $r=0$ and $\gamma_{\infty}=\gamma_{\infty}^{\prime}$. ∎
In view of Proposition 4.5 we define:
Definition 4.6 (Fastest shrinking geodesic).
Let $(M,o,f)$ be a tuple of a locally compact Hadamard space $M$, a point $o$ in $M$, and a convex, unbounded from below, linearly decreasing function $f:M\to\mathbb{R}$. Define the fastest shrinking geodesic of $(M,o,f)$ to be $\gamma=\gamma_{M,o,f}:[0,\infty)\to M$, the limit of $\gamma_{s}$ defined as in Definition 4.3. Then, for $a_{M,o,f}:=-\lim_{s\to\infty}\frac{1}{s}\min_{x\in B(s,o)}f(x)$ we have $f(\gamma(s))=-as\cdot(1+o(1))$ for all $s>0$. The constant $a_{M,o,f}$ is called the shrinking rate of f.
To analyze symmetric spaces (see §2.3) it is useful to consider the maximal flats in them. Hence, we show that restriction to ‘nice’ subspaces does not change the fastest shrinking geodesic or the shrinking rate.
Lemma 4.7 (Restriction).
Let $(M,o,f)$ be a triplet as in Definition 4.6. If $Y\subseteq M$ is a sub Hadamard space which contains the image of $\gamma_{M,o,f}$ then $\gamma_{M,o,f}=\gamma_{Y,o,f|_{Y}}$ and $a_{M,o,f}=a_{Y,o,f|_{Y}}$.
Proof.
We denote $\gamma:=\gamma_{M,o,f}$. By definition,
$a_{M,o,f}\leq a_{Y,o,f|_{Y}}\leq a_{\operatorname{Im}\gamma,o,f\circ\gamma}.$
Proposition 4.5 implies that $a_{\operatorname{Im}\gamma,o,f\circ\gamma}=a_{M,o,f}$. Hence, we deduce $a_{M,o,f}=a_{Y,o,f|_{Y}}$. Then, the uniqueness of the limit of geodesics in Proposition 4.5 implies that $\gamma_{Y,o,f|_{Y}}=\gamma$. ∎
The next claim follows directly from Definition 4.6.
Lemma 4.8 (Bounded shifts).
Let $(M,o,f)$ be a triplet as in Definition 4.6, and $\tilde{f}:M\to\mathbb{R}$ be a convex function such that $|f-\tilde{f}|$ is bounded. Then, $(M,o,f)$ also satisfies the assumptions of Definition 4.6 and
$\gamma_{M,o,f}=\gamma_{M,o,\tilde{f}},\quad a_{M,o,f}=a_{M,o,\tilde{f}}.$
4.1. Piecewise linear convex functions on euclidean spaces
Our goal in defining the fastest shrinking geodesic is to use them to study, and more specifically bound the ‘shrink-rate functions’, which are defined in the next section. The restriction of these functions to maximal flats (see Definition 2.3) are piecewise linear, up to bounded error. Moreover, the metric on the maximal flats is Euclidean. Therefore, we dedicated this subsection to study the fastest shrinking geodesics in this setting.
In this subsection, we assume that $M$ is the Euclidean space $\mathbb{R}^{n}$ with the standard Euclidean metric.
Claim 4.9.
For $v_{1},\dots,v_{n}\in\mathbb{R}^{n}$ and $a_{1},\dots,a_{n}\in\mathbb{R}$ define the function
$f(x):=\max_{1\leq i\leq n}(\left\langle x,v_{i}\right\rangle+a_{i}).$
Then, $f$ is unbounded from below if and only if
(4.5) $0\notin{\rm conv}\{v_{i}:i=1,...,n\}.$
Moreover, assuming (4.5), and denoting by $u$ the closest point to 00 in ${\rm conv}\{v_{i}:i=1,...,n\}$, we have:
  1. (1)
    $f$ decreases linearly,
  2. (2)
    the fastest shrinking geodesic for $f$ is
    $\gamma_{\mathbb{R}^{n},0,f}(t)=-tu/\|u\|,$
  3. (3)
    $a_{M,o,f}=-\|u\|$, and
  4. (4)
    there exists $C\in\mathbb{R}$, such that
    (4.6) $f(x)\geq\left\langle x,u\right\rangle+C.$
Proof.
If $0\in{\rm conv}\{v_{i}:i=1,...,n\}$, then for some non-negative $(\alpha_{i})_{i=1}^{n}$ with $\sum_{i=1}^{n}\alpha_{i}=1$ we have $\sum_{i=1}^{n}\alpha_{i}v_{i}=0$. Then,
$f(x)=\max_{1\leq i\leq n}(\left\langle x,v_{i}\right\rangle+a_{i})\geq\sum_{i=1}^{n}\alpha_{i}(\left\langle x,v_{i}\right\rangle+a_{i})=\sum_{i=1}^{n}\alpha_{i}a_{i}.$
is a bound from below on $f$.
Now, assume $0\notin{\rm conv}\{v_{i}:i=1,...,n\}$. We show that $f$ decreases linearly, which also implies that $f$ is unbounded from below. By Definition 4.3, it is enough to find a direction in which $f$ decreasing linearly, i.e. a vector $v$, such that $f(sv)\leq cs$ for some $c<0$ and all large enough $s$.
Let $u$ be the closest point to 00 in the convex hull ${\rm conv}\{v_{i}:i=1,...,n\}$. Since $u$ is of minimal length, for any $i=1,\dots,n$ and $p\in[0,1]$ we have
$\left\|u\right\|$ $\leq\left\|pv_{i}+(1-p)u\right\|$
$=p^{2}\left\|v_{i}\right\|+2p(1-p)\langle u,v_{i}\rangle+(1-p)^{2}\left\|u\right\|,$
which implies
$\langle u,v_{i}\rangle\geq\left\|u\right\|-\frac{p}{2(1-p)}(\left\|v_{i}\right\|-\left\|u\right\|).$
Thus, we deduce $\langle u,v_{i}\rangle\geq\left\|u\right\|$, for all $i=1,...,n$. We may now compute
$f(-tu)=\max_{1\leq i\leq n}(-t\left\langle u,v_{i}\right\rangle+a_{i})\leq-t\|u\|^{2}+\max_{1\leq i\leq n}a_{i},$
which decreases linearly, proving (1).
According to Proposition 4.5 there is a minimal shrinking geodesic and $a_{\mathbb{R}^{n},0,f}\leq-\|u\|$. On the one hand, for all $x\in\mathbb{R}^{n}$ we have
$f(x)=\max_{1\leq i\leq n}(\left\langle x,v_{i}\right\rangle+a_{i})\geq\sum_{i=1}^{n}\alpha_{i}(\left\langle x,v_{i}\right\rangle+a_{i})=\left\langle x,u\right\rangle+\sum_{i=1}^{n}\alpha_{i}a_{i},$
which implies (4) and $a_{\mathbb{R}^{n},0,f}\geq-\|u\|$. Thus, (2) and (3) are also satisfied. ∎
The following lemma shows that the constant of the lower bound in (4.6) is (in a sense) continuous in the vectors which define $f$ and can be chosen to be uniform on compact sets.
Lemma 4.10.
Let $V_{0}\subseteq\mathbb{R}^{n}$ be a finite set, $K$ be a compact set, and $(r_{v})_{v\in V_{0}}$ be a collection of continuous functions, $r_{v}:K\to[-\infty,\infty)$. Assume that for each $k\in K$ the function
$f_{k}:\mathbb{R}^{n}\to\mathbb{R},\quad f_{k}(x)=\max_{v\in V_{0}}(\left\langle x,v\right\rangle+r_{v}(k))$
is well defined (i.e. $f_{k}$ does not attain the value $\infty$- ∞), and the point of minimal length for each $k$
$u\in{\rm conv}(\{v\in V_{0}:r_{v}(k)>-\infty\})$
is independent of $k$, i.e., a constant vector. Then, there exists $C\in\mathbb{R}$ such that for all $k\in K,x\in\mathbb{R}^{n}$ we have
(4.7) $f_{k}(x)\geq\left\langle x,u\right\rangle+C.$
Proof.
By the definition of $u$, the value
$\xi(k):=\max_{\begin{subarray}{c}V_{1}\subseteq V_{0}\\ u\in{\rm conv}(V_{1})\end{subarray}}\min_{v\in V_{1}}r_{v}(k),$
is not $\infty$- ∞ for every $k\in K$. Since $\xi$ defines a continuous function $\xi:K\to\mathbb{R}$, it attains a minimum, which we denote by $C$.
We now show that for every $k\in K,\,x\in\mathbb{R}^{n}$, Equation (4.7) is satisfied. Indeed, for some $V_{1}\subseteq V_{0}$ with $u\in{\rm conv}(V_{1})$ we have $\min_{v\in V_{1}}r_{v}(k)\geq C$. Since $u\in{\rm conv}(V_{1})$, there exists a convex combination $\sum_{v\in V_{1}}\alpha_{v}v=u$. Then,
$f_{k}(x)$ $=\max_{v\in V_{0}}(\left\langle x,v\right\rangle+r_{v}(k))\geq\sum_{v\in V_{1}}\alpha_{v}(\left\langle x,v\right\rangle+r_{v}(k))$
$=\left\langle x,u\right\rangle+\sum_{v\in V_{1}}\alpha_{v}r_{v}(k)\geq\left\langle x,u\right\rangle+C,$
as wanted. ∎
Remark 4.11.
Lemma 4.10 does not hold if one allows $V_{0}$ to depend on $K$. For example, by taking $K=[0,1]$, and for each $k\in K$,
$V_{k}=\begin{cases}\{(1,-k),(1,1)\}&\text{if }k>0,\\ \{(1,0)\}&\text{if }k=0,\end{cases}$
$r_{0,(1,0)}=0,\quad r_{k,(1,-k)}=0,\quad r_{k,(1,1)}=\begin{cases}\frac{1}{k^{2}}&\text{if }k>0,\\ 0&\text{if }k=0,\end{cases}$
we get that for the sequence of functions
$f_{k}:\mathbb{R}^{2}\to\mathbb{R},\quad f_{k}(x,y)=x+\begin{cases}\max\left(-ky,y-\frac{1}{k^{2}}\right)&\text{if }k>0,\\ 0&\text{if }k=0,\end{cases}$
the result of Lemma 4.10 fails.
The following Claim will be used in section 6.
Claim 4.12.
Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be a piecewise linear function. Then the map $v\mapsto\lim_{t\to\infty}\frac{f(tv)}{t}$ is continuous.
Proof.
Since $f$ is a piecewise linear function, there exist $n\in\mathbb{N}$ and $v_{i},a_{i}$, $1\leq i\leq n$ such that $f(v)=\max_{1\leq i\leq n}(\left\langle v,v_{i}\right\rangle+a_{i})$ for every $v\in\mathbb{R}^{n}$. Then, for any $v\in\mathbb{R}^{n}$
$\lim_{t\to\infty}\frac{f(tv)}{t}=\max_{1\leq i\leq n}\left\langle v,v_{i}\right\rangle,$
which is a continuous function. ∎
5. The shrink-rate function
[…]
Recall the representation theory notations from §2.4, the definition of the Busemann function from §3, and the definition of the fastest shrinking geodesic from §4.
The main goal of this section is to analyse $f_{v}$ assuming $v$ is unstable. We prove Theorem 1.2 Part 1, by proving the next theorem.
Theorem 5.1 (Busemann bounds $f_{v}$ from below).
The tuple $(M,o,f_{v})$ satisfies the assumptions of Definition 4.6. In particular, there exists a fastest shrinking geodesic $\gamma=\gamma_{M,o,f_{v}}:[0,\infty)\to M$ and shrinking rate $a=a_{M,o,f_{v}}>0$ for it. Moreover, there exists $C\in\mathbb{R}$ such that for every $x\in M$, we have
(5.1) $f_{v}(x)\geq a\beta_{\gamma}(x)+C.$
Remark 5.2.
We will distinguish our function $f_{v}$ with the pathological in §1.3 using the fact that our function $f_{v}$ decay rate is almost linear, that is, $f(\gamma_{M,\pi(e),f}(s))=-as+O(1)$.
Since we are interested in a lower bound for $f_{v}$, by the equivalence of norms in finite dimensional vector space, without loss of generality from here on we may assume that the norm on $V$ is the $K$-invariant one which satisfies the assumptions of Lemma 2.11.
Observation 5.3 (Convexity of $f_{v}$).
The choice of quadratic form in Lemma 2.11 implies that $f_{v}$ is convex. Moreover, for every maximal flat $\pi(Ag)$, $g\in K$, we have
$f_{v}(\pi(ag))=\log\sum_{\lambda\in\Phi_{\varrho}}\lambda(a)^{2}\|(gv)_{\lambda}\|^{2}=\max_{\lambda\in\Phi_{\varrho}}\left(\log\lambda(a)+r_{\lambda}\right)+O(1),$
where $r_{\lambda}=\log\|(gv)_{\lambda}\|\in\mathbb{R}\cup\{-\infty\}$ depends continuously on $g$ and $v$.
Recall that $v\in V$ is called unstable if $0\in\overline{Gv}\setminus Gv$, where $\overline{Gv}$ denotes the Zariski-closure of $Gv$.
Claim 5.4.
If $v$ is unstable, then the function $f_{v}$ decreases linearly (see Definition 4.4).
Remark 5.5.
This claim follows from [21], but we prove it differently to illustrate the techniques in a simpler way.
Proof.
Recall that by Definition 2.3 the metric on $M$ (explicitly defined §2.3.3) restricted to any flat is Euclidean. Recall that we denoted by $\mathfrak{a}$ the Lie algebra of $A$.
Assume that $f_{v}$ is unbounded from below on $F_{k}=\pi(Ak)$ for some $k\in K$. The restriction of $f_{v}$ to $F_{k}$ can be described as follows: Since $F_{k}$ is a maximal flat, by Observation 5.3, the explicit construction of the metric in §2.3.3, and the definition of a maximal flat (see Definition 2.3), it follows that for $\mathtt{a}\in\mathfrak{a}$ the difference $f_{v}(\exp(\mathtt{a})k)-\tilde{f}_{k}(\mathtt{a})$ is bounded for
(5.2) $\tilde{f}_{k}(\mathtt{a})=\max_{\lambda\in\Phi_{\varrho}}\left(\left\langle\mathtt{a}_{\lambda},\mathtt{a}\right\rangle+r_{\lambda}(k)\right),$
where $\mathtt{a}_{\lambda}$ is the unique vector such that $\left\langle\mathtt{a}_{\lambda},\mathtt{a}\right\rangle=\lambda(\mathtt{a})$ and $r_{\lambda}(k)=\log\|(kv)_{\lambda}\|\in\mathbb{R}\cup\{-\infty\}$. Here $\lambda$ is viewed as an additive character on $\mathfrak{a}$, in particular, if $\tilde{\lambda}$ is the same character viewed as a multiplicative character on $A$ (i.e., as in Observation 5.3), then $\lambda(\mathtt{a})=\log\tilde{\lambda}(\exp(\mathtt{a}))$. Then, in particular, $\tilde{f}_{k}$ is unbounded from below. Claim 4.9 implies that $\tilde{f}_{k}$ decreases linearly, and hence also $f_{v}$.
Now, assume to the contrary that $f_{v}$ is bounded from below on every maximal flat of the form $F_{k}=\pi(Ak)$ for $k\in K$. It follows that $\tilde{f}_{k}$ are bounded from below as well. By Claim 4.9 we get that
$0\in{\rm conv}(\{v_{\lambda}:\lambda\in\Phi_{\varrho},r_{\lambda}(k)\geq 0\}.$
Lemma 4.10 implies that there is a uniform lower bound on $\tilde{f}_{k}$ for all $k\in K$, and implies that $f_{v}$ has a uniform lower bound on $\bigcup_{k\in K}F_{k}$. Cartan decomposition of $G$, 2.9, implies that $\bigcup_{k\in K}F_{k}=M$. This contradicts the assumption that $v$ is unstable. ∎
Note that Observation 5.3 and Claim 5.4 prove the first part of Theorem 5.1. That is, the tuple $(M,o,f_{v})$ satisfies the assumptions of Definition 4.6. Therefore, from here on we may let $\gamma$ be the fastest shrinking geodesic of $(M,o,f_{v})$, and $a>0$ satisfy the conclusion of Proposition 4.5 for $f_{v}$. In view of Observation 5.3, the study of the restriction of shrink-rate functions to maximal flats is reduced to §4.1.
5.1. Lower bound on maximal flat
The first step in the proof of Theorem 5.1 is showing that Equation (5.1) is satisfied on maximal flats (see §2.3 for the definition of a maximal flat).
Proposition 5.6.
There exists a constant $C>0$, which only depends on $v$, so that for every maximal flat $F_{0}\subset M$ which contains the image of $\gamma$ and every $x\in F_{0}$, inequality (5.1) holds.
Proof.
Let $F_{0}=\pi(Ag)$, $g\in K$, be a maximal flat containing $\gamma$, and $K_{0}\subseteq K$ be the stabilizer of $\gamma$. By Claim 2.6, the group $K_{0}$ acts transitively on the set of maximal flats containing $\gamma$.
As in the proof of Claim 5.4, since $F_{0}$ is a maximal flat, the metric constructed on it in §2.3.3 is Euclidean. By Corollary 2.10 we may assume that for some $\mathtt{a}\in\mathfrak{a}$, $\gamma(t)=\pi(\exp(\mathtt{a})g)$. Moreover, by Example 3.2 the restriction of $\beta_{\gamma}$ to $F_{0}$ is defined by
(5.3) $\beta_{\gamma}(\pi(\exp(\mathtt{b})g))=-\left\langle\mathtt{a},\mathtt{b}\right\rangle.$
Fix $k_{0}\in K_{0}$. We wish to show that each $p\in F_{k_{0}}:=\pi(Agk_{0})$ satisfies
$f_{v}(p)\geq a\beta_{\gamma}(p)+C.$
Write $p=\exp(\mathtt{b})gk_{0}$. Since $K_{0}$ preserves $\beta_{\gamma}$, by (5.3), this is equivalent to showing that
$f_{\varrho(k_{0})v}(p)\geq-\left\langle\mathtt{a},\mathtt{b}\right\rangle+C.$
Since $f_{\varrho(k_{0})v}$ is a shift (by $k_{0}$) of the function $f_{v}$, by Lemma 4.7, the fastest shrinking geodesic of the function $f_{\varrho(k_{0})v}\mid_{F_{k_{0}}}$ is $\gamma_{gk_{0}}(t)=\pi(\exp(\mathtt{a})gk_{0})$ and the shrinking rate of it is $a$. By Observation 5.3 it follows that for any $p=\exp(\mathtt{b})gk_{0}\in F_{k_{0}}$ $f_{\varrho(k_{0})v}(p)-\tilde{f}_{gk_{0}}(\mathtt{b})$ is bounded, where for $k:=gk_{0}$ the function $\tilde{f}_{k}$ is defined as in (5.2). It follows from Lemma 4.8 that the fastest shrinking geodesic of $\tilde{f}_{k}$ is $\gamma_{k}$ and the shrinking rate of it is $a$. The result now follows from Lemma 4.10. ∎
5.2. proof of the second part of Theorem 5.1
As in the proof of Proposition 5.6, there exists $\mathtt{a}\in\mathfrak{p}$ such that for any $t>0$
$\gamma(t)=\pi(\exp(t\mathtt{a})).$
In particular, we may extend the domain of $\gamma$ to be $\mathbb{R}$.
Let $C\in\mathbb{R}$ be the constant which satisfies the conclusion of proposition 5.6. Fix $x_{0}=\pi(g_{0})\in M$ for some $g_{0}\in G$. We will show that $f_{v}(x_{0})-a\beta_{\gamma}(x_{0})\geq C$
By Lemma 2.13, the element $g_{0}$ can be decomposed as $g_{0}=kp_{1}p_{2}$, where $p_{1}\in U_{\mathtt{a}}$, $p_{2}\in T_{\mathtt{a}}$. Let $\gamma_{x_{0}}$ be the geodesic defined by $t\mapsto\pi(\exp(t\mathtt{a})p_{1}p_{2})$. Let $\tilde{f}:[0,\infty)\rightarrow\mathbb{R}$ be defined by
$\tilde{f}(t):=f_{v}(\gamma_{x_{0}}(t))-a\beta_{\gamma}(\gamma_{x_{0}}(t)).$
By definition, $p_{1}p_{2}\in P_{\mathtt{a}}$ and hence $\beta_{\gamma}(\gamma_{x_{0}}(t))=\beta_{\gamma}(\pi(\exp(t\mathtt{a})p_{1}p_{2}))$ is linear in $t$, as $\mathtt{a}\in\operatorname{Lie}(P_{\mathtt{a}})$ and Remark 3.8. Hence $\tilde{f}$ is convex.
We will prove that
$\lim_{t\to\infty}\tilde{f}(t)\geq C,$
and then Eq. (5.1) follows using the next observation.
Observation 5.7.
Let $f:\mathbb{R}\to\mathbb{R}$ be a convex function. Assume $\lim_{t\to\infty}f(t)$ is well defined. Then, $f$ is monotone decreasing, and
$f(0)\geq\lim_{t\to\infty}f(t).$
Since $p_{2}$ commutes with $\exp(\mathtt{a})$, it follows that $\pi(p_{2})$ and $\pi(\exp(\mathtt{a}))$ lie on the same maximal flat (see Remark 2.4). Hence, Inequality (5.1) holds for any point in $\{\pi(p_{2}\exp(t\mathtt{a})):t\in\mathbb{R}\}$. Since $p_{1}\in U_{\mathtt{a}}$ we get that
(5.4) $d(\pi(\exp(t\mathtt{a})p_{1}p_{2}),\pi(\exp(t\mathtt{a})p_{2}))=d(\pi(\exp(t\mathtt{a})p_{1}\exp(-t\mathtt{a})),o)\xrightarrow{t\to\infty}0$
Since $f_{v}$ and $\beta_{\gamma}$ are Lipschitz continuous it follows that
$\lim_{t\to\infty}\tilde{f}(t)$ $=\lim_{t\to\infty}f_{v}(\pi(\exp(t\mathtt{a})p_{1}p_{2}))+a\beta_{\gamma}(\pi(\exp(t\mathtt{a})p_{1}p_{2}))$
$=\lim_{t\to\infty}f_{v}(\pi(\exp(t\mathtt{a})p_{2}))+a\beta_{\gamma}(\pi(\exp(t\mathtt{a})p_{2}))$
$\geq C,$
where the first equality is by definition, the second one is by Eq. (5.4) and the Lipschitz continuity of $f_{v}$ and $\beta_{\gamma}$, and the inequality at the end is by Proposition 5.6 and the fact that $\pi(\exp(t\mathtt{a})p_{2}$ lies in a flat containing $\gamma$. ∎
6. An Algebraic interpretation and a result by Kempf
Inspecting our setting from an algebraic point of view, one can use Kempf [21] to obtain Theorem 5.1, which is a more explicit version of Theorem 1.2 Part 1.
6.1. algebraic analogous of geometric claims
Kempf studied group homomorphisms in $\operatorname{Hom}_{k}(\mathbb{G}_{m},G)$, for $k\subset\mathbb{R}$, which are algebraic analogous of the geometric notion of geodesics. The following claim describes the connection between them:
Claim 6.1 (Connection between geodesics and group homomorphism).
Let $G$ an $\mathbb{R}$-algebraic semisimple group (as in §2.3). For every nontrivial homomorphism $\tau:\mathbb{G}_{m}\to G$ there is $\mathtt{a}\in\mathfrak{p}$ such that the homomorphism $t\mapsto\tau(\exp(t))$ is a conjugate, by an element in $P_{\mathtt{a}}$ to the homomorphism $t\mapsto\exp(t\mathtt{a})$ (see §2.3.3 for the definition of $\mathfrak{p}$ and §2.5 for the definition of $P_{\mathtt{a}}$).
Moreover, $\mathtt{a}$ is uniquely determined as a function of $\tau$.
Remark 6.2.
A similar claim could be made for a general homomorphism in $\operatorname{Hom}(\mathbb{R},G)$, but since the adjoint action of the homomorphism may have non-real eigenvalues, the claim and its proof are more elaborate.
Remark 6.3.
In the setting of Claim 6.1, $P_{\mathtt{a}}$ can be defined using $\tau$ by
$P_{\mathtt{a}}=P_{\tau}:=\{g\in G:\tau(t)g\tau(t^{-1})\text{ converges as }t\to 0\}.$
Proof.
Let $\tau:\mathbb{G}_{m}\to G$ be a homomorphism. First note that by [24, Proposition 2.6], $P_{\tau}$, as defined in Remark 6.3, is parabolic.
Now, $\tau((0,\infty))$ is a one parameter torus of $G$. In particular, there exists $g\in G$ so that $\tau((0,\infty))=\{g\exp(\mathtt{a}\mathbb{R})g^{-1}\}$ for some $\mathtt{a}\in\mathfrak{p}$. Since $\mathfrak{p}$ is invariant under $K$ and by Lemma 2.13, $G=P_{\tau}K$, we may assume that $g\in P_{\tau}$. Then, the element $\mathtt{a}\in\mathfrak{p}$ satisfies $\tau(\exp(s))=g\exp(s\mathtt{a})g^{-1}$.
Note that $\gamma:s\mapsto\pi(\tau^{\prime}(\exp(-s/\|\mathtt{a}\|)))$ is a geodesic. The definition of the parabolic subgroup implies that $d(\gamma(s),\pi(\tau(\exp(-s/\|\mathtt{a}\|))))$ is bounded as $s\to\infty$. Lemma 2.2 implies that this defines $\gamma$ uniquely. ∎
Definition 6.4.
[…]
Remark 6.5.
The other direction of Claim 6.1 is not generally true. That is, there exist geodesic rays which are not corresponding to any algebraic geodesic rays.
For example, take $G=\operatorname{SL}_{3}(\mathbb{R}),\,K=SO(3)$, and
$\gamma(t)=\pi(\exp(t\operatorname{\mathrm{diag}}(1,\sqrt{2},-1-\sqrt{2}))),$
then $\gamma$ does arise from any algebraic homomorphism in $\operatorname{Hom}_{\mathbb{R}}(\mathbb{G}_{m},G)$.
From here on we assume that $G$ is defined over a fix field $k\subseteq\mathbb{R}$.
Definition 6.6 ($k$-algebraic geodesics and Busemann functions).
A geodesic ray $\gamma:[0,\infty)\to M$ is called $k$-algebraic if it associated to a $k$-algebraic group homomorphism $\tau\in\operatorname{Hom}_{k}(\mathbb{G}_{m},G)$. The shrinking constant $a_{\tau}$ of $\tau$ depends on $\gamma$ up to a multiplication by a rational constant and is called the shrinking constant of $\gamma$. In this case let $\beta_{\gamma}$ be the Busemann function defined by $\gamma$ (as in (3.1)). For every $q\in\mathbb{Q}^{>0}$ the function $qa_{\tau}\beta_{\gamma}$ is called the $k$-renormalized Busemann function.
We can now provide an algebraic analogous of Theorem 3.5.
Theorem 6.7.
The following three classes of functions are equal:
  1. (1)
    Busemann functions: The class of constant shifts of $k$-renormalized Busemann functions.
  2. (2)
    Homomorphisms from parabolics: The class of functions which are projections of $k$-positive homomorphisms $P\to\mathbb{R}$, where $P$ is a $k$-parabolic subgroup of $G$, see Definition 3.4 of $k$-positive homomorphism.
  3. (3)
    Lengths of highest weight vectors in fundumental representations: The class of functions of the form
    $g\mapsto\sum_{j=1}^{d}\alpha_{j}\log\|\varrho_{j}(g)v_{j}\|+C,$
    where $\varrho_{j}$ are $k$-fundamental representations and $v_{i}$ are the appropriate highest weight vectors, which correspond to maximal $k$-parabolic groups, and are defined in §2.5, for some choice of an $k$-simple system, the $\alpha_{j}$ are non-negative, and $C\in\mathbb{R}$.
  4. (4)
    Length of a single parabolic equivariant vector: The class of functions of the form
    $g\mapsto\alpha\log\|\varrho(g)v\|+C,$
    where $\varrho$ is a $k$-algebraic representations with a vector $v$ satisfying that the ray $kv$ is stabilized by a parabolic subgroup.
Proof.
The proof of the equivalence between the first three classes follows in a very similar way to the proof of Theorem 3.5, using the more general theory which is presented in §2.5. The implication of Part (4) $\implies$ the rest, is similar to Theorem 3.9. We are left to show Part (4) from the other three.
Assume a function $f$ is in the third class, i.e.,
$f(\pi(g))=\sum_{i=1}^{d}\alpha_{i}\log\|\varrho_{i}(g)v_{i}\|+C,$
where $C\in\mathbb{R}$, for any $1\leq i\leq d$, $\alpha_{i}$ is rational and nonnegative, and the vector $v_{i}$ is a highest weight vector in $V_{i}$. In particular, the line $kv_{i}$ is $P_{i}$-invariant for a $k$-maximal parabolic subgroup $P_{i}$, and there exists a single (minimal) $k$-parabolic subgroup $P$ contained in each $P_{i}$. Also, up to multiplication by an integer, we may assume that the $\alpha_{i}$’s are integers.
Let
$v:=v_{1}^{\otimes\alpha_{1}}\otimes v_{2}^{\otimes\alpha_{2}}\otimes\cdots\otimes v_{k}^{\otimes\alpha_{k}}\in V_{1}^{\otimes\alpha_{1}}\otimes V_{2}^{\otimes\alpha_{2}}\otimes\cdots\otimes V_{k}^{\otimes\alpha_{k}}.$
Then, the line $kv$ is $P$-invariant via the representation $\varrho_{1}^{\otimes\alpha_{1}}\otimes\varrho_{2}^{\otimes\alpha_{2}}\otimes\cdots\otimes\varrho_{k}^{\otimes\alpha_{k}}$, and this vector in this representation represents the function as in class (4).
6.2. Kempf’s result
The following definition is a normalized sense of how fast an algebraic torus shrinks a vector.
Definition 6.8 (Shrinking rate of an algebraic torus).
Let $\tau:\mathbb{G}_{m}\to G$ be an algebraic nontrivial homomorphism, $\varrho:G\rightarrow\operatorname{GL}(V)$ be an algebraic representation, and $v\in V$.
$V$ can be decomposed into its eigenspaces with respect to the $\mathbb{G}_{m}$ action $\varrho\circ\tau$
(6.1) $V=\bigoplus_{n\in\mathbb{Z}}V_{n},$
where for any $n\in\mathbb{Z}$
(6.2) $V_{n}:=\left\{w\in V:\forall t\in\mathbb{G}_{m},\,\varrho\circ\tau(t)w=t^{n}w\right\}.$
Let $m(v,\tau)$ be the maximal integer for which $v\in\bigoplus_{n\geq m(v,\tau)}V_{n}$.
Note that for every $n\in\mathbb{N}$ we have
$m(v,\tau^{n})=nm(v,\tau).$
Using $e\in T_{1}\mathbb{G}_{m}$, we define a norm on $\operatorname{Hom}_{\mathbb{Q}}(\mathbb{G}_{m},G)$
(6.3) $\|\tau\|:=\sqrt{B(D_{e}\tau(e),D_{e}\tau(e))},$
where $B=B_{e}$ is the killing form (see §2.3.3) and $D_{e}\tau:T_{1}\mathbb{G}_{m}\to T_{1}G\cong\mathfrak{g}$ is the differential of $\tau$ at $e$.
Remark 6.9.
Note that for $\mathtt{a}\in\mathfrak{p}$ which satisfies the conclusion of Claim 6.1 for $\tau$, we have $\left\|\tau\right\|=\left\|\mathtt{a}\right\|$, where the norm on $\mathfrak{p}$ is as defined in §2.3.3. In particular, the norm on $\operatorname{Hom}_{\mathbb{Q}}$ is well defined. Moreover, since the derivative of the action of $\tau$ on $\mathfrak{g}$ has integer, it follows that for any $\tau\in\operatorname{Hom}_{\mathbb{Q}}(\mathbb{G}_{m},G)$ the value in (6.3) is a square root of an integer (see also the discussion in [21, §2]).
Lemma 6.10 ( [21, Lemma 3.2](c)).
The quantity $m(v,\tau)$ can be described as follows: Consider the map $f=f_{\tau,\varrho,v}:\mathbb{G}_{m}\to V$ defined by
$f(t)=\varrho\circ\tau(t)v.$
For every functional $\varphi\in V^{*}$ we can consider the function $\varphi\circ f$ and compute its valuation at $t=0$. $m(v,\tau)$ is the minimal such valuation.
Remark 6.11.
The alternative description of $m(v,\tau)$ is the algebraic analogous for the shrinking rate of $\|\varrho(\tau(t))v\|$ as $t\to 0$.
Recall the definition of a unstable vector from §5.
Theorem 6.12 ([21, Theorem 4.2]).
Let $G$ be a semisimple $k$-algebraic group, $\varrho:G\to\operatorname{GL}(V)$ an algebraic $G$-representation, and $v\in V$ a unstable vector. Then, there exists $\tau\in\operatorname{Hom}_{\mathbb{Q}}(\mathbb{G}_{m},G)$ which maximize $\frac{m(v,\tau)}{\|\tau\|}$ and for which $m(v,\tau)>0$. This subgroup $\tau$ is unique up to taking power and conjugating by elements in $P_{\tau}$. We call such $\tau$ Kempf’s homomorphism of $v$.
Recall the definition of the shrink-rate function of a vector from §5, as well as the definitions of the fastest shrinking geodesic of a function and the shrinking rate of it from Definition 4.6.
The following claim identifies the fastest shrinking geodesic of the shrink-rate function of a vector and Kempf’s homomorphism of it.
Lemma 6.13.
Let $\varrho:G\to\operatorname{GL}(V)$ be a $k$-algebraic representation, and $v\in V(k)$ be an unstable vector. Then, the fastest shrinking geodesic of $f_{v}$ (as defined in §5) is of the form $\gamma_{\tau}$ where $\tau\in\operatorname{Hom}_{k}(\mathbb{G}_{m},G)$ is Kemp’s homomorphism of $v$. In addition, the shrinking rate satisfies $a_{f_{v},M,o}=\frac{m(v,\tau)}{\|\tau\|}$.
Proof.
Let $\varsigma\in\operatorname{Hom}(\mathbb{G}_{m},G)$ be nontrivial. Then, by Lemma 6.10 and the definitions of $f_{v}$, we have
$\frac{m(v,\varsigma)}{\|\varsigma\|}=\lim_{s\to\infty}\frac{-f_{v}(\gamma_{\varsigma}(s))}{s}.$
Using Claim 6.1, this implies that Kempf’s homomorphism is the algebraic geodesic rays that shrinks $f_{v}$ the fastest. In particular, $a_{f_{v},M,o}\geq\frac{m(v,\varsigma)}{\|\varsigma\|}$ for all $\varsigma\in\operatorname{Hom}(\mathbb{G}_{m},G)$ and equality occurs if and only if the fastest shrinking geodesic is algebraic.
Without loss of generality, assume that the fastest shrinking geodesic is $\gamma(s)=\exp(s\mathtt{a})$ for some $\mathtt{a}\in\mathfrak{a}$. Note that the set of algebraic geodesic rays is dense in the set of all rays, and that by Observation 5.3 and Claim 4.12, the function
$\gamma\mapsto\lim_{s\to\infty}\frac{-f_{v}(\gamma(s))}{s},$
defined for all geodesic rays $\gamma$ in $\pi(A)$, is continuous. It follows that $a_{f_{v},M,o}$ is at most the supremum of $\lim_{s\to\infty}\frac{-f_{v}(\gamma(s))}{s}$ over all algebraic geodesic rays $\gamma$ from $o$ in $\pi(A)$. This concludes the proof. ∎
7. Alternative proof of Theorem 1.1
Let $\varrho:G\to\operatorname{GL}(V)$ be a $\mathbb{Q}$-representation, $V$ be equipped with a $K$-invariant norm, and $v$ be an unstable non-zero vector in $V(\mathbb{Q})$.
Define the shrink-rate function $f_{v}:G\to\mathbb{R}$ of $v$ as in §5, i.e.,
$f_{v}(\pi(g))=\log\|gv\|.$
Note that the function $f_{v}$ is the quantity on the left hand side of (1.1), which we wish to control. By Theorem 5.1 there exists a fastest shrinking geodesic for $f_{v}$ (see §4 for the definition of a fastest shrinking geodesic), denote it by $\gamma$, and $C\in\mathbb{R}$ such that for every $x\in M:=K\backslash G$, we have
(7.1) $f_{v}(x)\geq\tilde{a}\beta_{\gamma}(x)+C,$
where $\beta_{\gamma}$ is the Busemann function associated to $\gamma$ (defined in §3) and $\tilde{a}$ is the shrinking rate of $f_{v}$ on $\gamma$, that is, $\tilde{a}=-\lim_{t\to\infty}\frac{f_{v}(\gamma(t))}{t}$. Moreover, since $v\in V(\mathbb{Q})$, by Lemma 6.13 $\gamma$ is $\mathbb{Q}$-algebraic, and hence $\tilde{a}\beta_{\gamma}$ is a $\mathbb{Q}$-renormalized Busemann function (see Definition 6.6).
Now, Theorem 6.7 connects the above Busemann function to a $\mathbb{Q}$-highest weight representation, $\varrho^{\prime}:G\rightarrow\operatorname{GL}(W)$ implying that there exist non-negative $\tilde{a}$ and $\tilde{C}$ such that for any $g\in G$
(7.2) $\beta_{\gamma}(\pi(g))=\tilde{\alpha}\log\|\varrho^{\prime}(g)w\|-\tilde{C}.$
Combining (7.1) and (7.2), one may deduce that for some non-negative $a,c$ and all $g\in G$, we have
$f_{v}(\pi(g))\geq a\log\|\varrho^{\prime}(g)w\|-c.$
The claim now follows from the definition of $f_{v}$.
8. Appendix
Here we prove Lemma 2.11. We restate it below for the convenience of the reader.
Lemma 8.1 (Construction of bilinear form).
[…]
For the proof we need some definitions regarding Galois group actions.
Definition 8.2 (Galois Action).
Denote by $\operatorname{Gal}(\mathbb{C}/\mathbb{R})$ the Galois group, i.e., the group of automorphisms of $\mathbb{C}$ which fix $\mathbb{R}$ pointwise. Then, the only elements in $\operatorname{Gal}(\mathbb{C}/\mathbb{R})$ are the identity map and the complex conjugation, which we denote by $\operatorname{{\mathrm{conj}}}$. Moreover, there is an action of $\operatorname{Gal}(\mathbb{C}/\mathbb{R})$ on:
  1. (i)
    complexifications of $\mathbb{R}$-algebraic objects, such as $G^{\mathbb{C}}$ and $V\otimes\mathbb{C}$, in a natural way.
  2. (ii)
    the category of complex vector spaces in the following way: Given a vector space $U$, let $U^{\operatorname{{\mathrm{conj}}}}$ be a vector space with the same set of points as $U$, let the addition on $U^{\operatorname{{\mathrm{conj}}}}$ be the same as on $U$, and the multiplication by a scalar $c\in\mathbb{C}$ on $U^{\operatorname{{\mathrm{conj}}}}$ be the multiplication by the conjugate of $c$ on $U$. We denote the topological map from $U$ to $U^{\operatorname{{\mathrm{conj}}}}$ by $\operatorname{{\mathrm{conj}}}$.
  3. (iii)
    the category of complex representations of $G^{\mathbb{C}}$ as follows: For $\sigma:G^{\mathbb{C}}\to\operatorname{GL}(U)$ we can set a (canonical) representation $\sigma^{\operatorname{{\mathrm{conj}}}}:G^{\mathbb{C}}\to\operatorname{GL}(U^{\operatorname{{\mathrm{conj}}}})$ which acts by
    $\sigma^{\operatorname{{\mathrm{conj}}}}(g)(u)=\operatorname{{\mathrm{conj}}}(\sigma({\operatorname{{\mathrm{conj}}}}(g))({\operatorname{{\mathrm{conj}}}(u)})).$
  4. (iv)
    $G$ actions have the following phenomenon: Assume $G$ is a group acting on a space $X$ via $\sigma:G\to\operatorname{Aut}(X)$. An action $\operatorname{Gal}(\mathbb{C}/\mathbb{R})\lefttorightarrow\varrho$ is a tuple of actions $\operatorname{Gal}(\mathbb{C}/\mathbb{R})\lefttorightarrow G,\operatorname{Gal}(\mathbb{C}/\mathbb{R})\lefttorightarrow X$ such that
    $\operatorname{{\mathrm{conj}}}(\sigma(g)(x))=\sigma(\operatorname{{\mathrm{conj}}}(g))(\operatorname{{\mathrm{conj}}}(x)),$
    for every $g\in G,x\in X$. The above is equivalent to an extension of $\sigma$ to an action of the semidirect product $\tilde{\sigma}:\operatorname{Gal}(\mathbb{C}/\mathbb{R})\ltimes G\to\operatorname{Aut}(X)$.
Proof of Lemma 8.1.
To show the claim, we begin by extending the maximal compact subgroup $K$ to the complexification $G^{\mathbb{C}}$. As stated in Definition 8.2, The action of $G$ on $V$ can be extended to $V\otimes\mathbb{C}$. Then, there exists a positive bilinear form on $V^{\mathbb{C}}$ so that the linear spaces $V_{\lambda}^{\mathbb{C}}$ are orthogonal with respect to it. Using the unitarian trick, we can deduce that this bilinear form on is $K$-invariant.
To find an explicit description of $K^{\mathbb{C}}$, the maximal compact subgroup of $G^{\mathbb{C}}$, we use standard constructions, see[20, §VI,2]. Recall that $\theta:\mathfrak{g}\to\mathfrak{g}$ is the Cartan involution antihomomorphism, it acts on the lie algebra $\mathfrak{g}=\mathfrak{k}\oplus\mathfrak{p}$ by inverting $\mathfrak{k}$ and preserving $\mathfrak{p}$. Since $\mathfrak{g}^{\mathbb{C}}=\mathfrak{g}\otimes\mathbb{C}$, we may extend $\theta$ to a real invaluation $\theta^{\mathbb{C}}:\mathfrak{g}^{\mathbb{C}}\to\mathfrak{g}^{\mathbb{C}}$ by defining
$\theta^{\mathbb{C}}=\theta\otimes\operatorname{{\mathrm{conj}}}.$
Since $\mathfrak{g}\otimes\mathbb{C}=\mathfrak{k}\otimes\mathbb{C}\oplus\mathfrak{p}\otimes\mathbb{C}$, the above defines a decomposition $\mathfrak{g}^{\mathbb{C}}=\mathfrak{k}^{\mathbb{C}}\oplus\mathfrak{p}^{\mathbb{C}}$, viewed as real vector spaces, for
$\mathfrak{k}^{\mathbb{C}}:=$ $\mathfrak{k}\oplus i\mathfrak{p},$
$\mathfrak{p}^{\mathbb{C}}:=$ $i\mathfrak{k}\oplus\mathfrak{p}.$
Moreover, $\mathfrak{k}^{\mathbb{C}}$ is the $1$ eigenspace of $\theta^{\mathbb{C}}$ and $\mathfrak{p}^{\mathbb{C}}$ is the $-1$ eigenspace.
Since $\theta^{\mathbb{C}}$ is a lie algebra homomorphism, we deduce that its fixed points, i.e., $\mathfrak{k}^{\mathbb{C}}$, is also a lie algebra. Direct computation shows that the killing form of the group $G^{\mathbb{C}}$, viewed as a real algebraic group, is positive on $\mathfrak{p}^{\mathbb{C}}$ and negative on $\mathfrak{k}^{\mathbb{C}}$. This implies that this is indeed the cartan decomposition of $G^{\mathbb{C}}$, thought of as a real Lie group. In particular, $K^{\mathbb{C}}:=\exp\mathfrak{k}^{\mathbb{C}}$ is a maximal compact subgroup of $G^{\mathbb{C}}$.
The unitarian trick states that since $K^{\mathbb{C}}$ is a real Lie group whose complexification is again $G^{\mathbb{C}}$, the restriction of representations of $G^{\mathbb{C}}$ to $K^{\mathbb{C}}$ induces an isomorphism of the category of complex representations, see [36, Theorem 4.11.14].
We assume without loss of generality that $\varrho$ is irreducible. Then, one of the following holds:
  1. (1)
    $\varrho\otimes\mathbb{C}$ is an irreducible representation of $G^{\mathbb{C}}$.
  2. (2)
    There is a decomposition $\varrho\otimes\mathbb{C}\cong\sigma\oplus\sigma^{\operatorname{{\mathrm{conj}}}}$.
We will first prove the lemma for $\varrho$ as in the case (1). In that case, $\varrho|_{K^{\mathbb{C}}}$ is also irreducible.
Claim 8.3.
On every irreducible complex $K^{\mathbb{C}}$-representation there is exactly one $K^{\mathbb{C}}$ invariant, positive definite Hermitian form up to multiplication by a positive scalar.
Proof.
Let $\sigma:K^{\mathbb{C}}\to\operatorname{GL}(U)$ be an irreducible representation. There is at least one such invariant positive definite Hermitian form, as we can average a non-invariant positive definite Hermitian form along the $K^{\mathbb{C}}$ action.
If there are two such positive definite Hermitian forms, $H_{1},H_{2}$, consider the infimum
$\alpha:=\inf\{\alpha^{\prime}>0:H_{1}-\alpha^{\prime}H_{2}\text{ is not positive definite}\}.$
Then, the hermitian form $H_{1}-\alpha H_{2}$ is non-negative definite and $K^{\mathbb{C}}$-invariant. Moreover, the set
$W=\{w\in U:(H_{1}-\alpha H_{2})(w,w)=0\}$
is nonempty, not equal to $U$, and is $K^{\mathbb{C}}$-invariant. This contradicts the irreducability of $\sigma$ as a $K^{\mathbb{C}}$ representation. ∎
Denote the unique invariant, positive definite Hermitian form on $V^{\mathbb{C}}=V\otimes\mathbb{C}$ by $H^{\mathbb{C}}$. Since the the Galois group $\operatorname{Gal}(\mathbb{C}/\mathbb{R})$ acts on the $K^{\mathbb{C}}$, $V^{\mathbb{C}}$, and the $K^{\mathbb{C}}$-action $\varrho\otimes\mathbb{C}$ (i.e., satisfies property (iv) of Definition 8.2), we deduce that $H^{\mathbb{C}}$ is Galois invariant. Hence, $H^{\mathbb{C}}$ is induced from a positive definite symmetric billinear form $H$ on $V$.
We are left to show that $H$ satisfies the desired property, that is, the linear spaces $V_{\lambda}$ are orthogonal with respect to it. Assume that $V\cong\mathbb{R}^{n}$, $V\otimes\mathbb{C}\cong\mathbb{C}^{n}$, and $H^{\mathbb{C}}$ is the standard hermitian form on $\mathbb{C}^{n}$. Then, it is enough to show that $A$ is sent by $\varrho^{\mathbb{C}}$ to a group of Hermitian matrices.
As in Definition 8.2, the representation $\varrho$ defines a map $\varrho^{\mathbb{C}}:G^{\mathbb{C}}\to\operatorname{GL}_{n}(\mathbb{C})$. Define $\theta_{n}:\operatorname{GL}_{n}(\mathbb{C})\to\operatorname{GL}_{n}(\mathbb{C})$ by
$\theta_{n}(M)=\operatorname{{\mathrm{conj}}}(M^{-t}).$
Since $\varrho^{\mathbb{C}}(A)$ is invariant under $\theta^{\mathbb{C}}$, in order to show that $\varrho^{\mathbb{C}}(A)$ is also invariant under $\theta_{n}$, it suffices to prove that $\theta_{n}\circ\varrho^{\mathbb{C}}\circ\theta^{\mathbb{C}}=\varrho^{\mathbb{C}}$. Since $K^{\mathbb{C}}$ preserves the Hermitian form $H^{\mathbb{C}}$, it follows that $\varrho^{\mathbb{C}}(K_{\mathbb{C}})\subseteq U(n)$. In particular, $\theta_{n}\circ\varrho^{\mathbb{C}}\circ\theta^{\mathbb{C}}|_{K^{\mathbb{C}}}=\rho|_{K^{\mathbb{C}}}$. Since $K^{\mathbb{C}}$ is Zariski dense in $G^{\mathbb{C}}$, it follows that $\theta_{n}\circ\varrho^{\mathbb{C}}\circ\theta^{\mathbb{C}}=\varrho^{\mathbb{C}}$, as desired.
Now, assume we are in case (2), i.e., $\varrho\otimes\mathbb{C}\cong\sigma\oplus\sigma^{\operatorname{{\mathrm{conj}}}}$. for some $\sigma:G^{\mathbb{C}}\to\operatorname{GL}(U)$. The $\mathbb{R}$-linear map $\operatorname{{\mathrm{conj}}}$ on $V\otimes\mathbb{C}\cong U\oplus U^{\operatorname{{\mathrm{conj}}}}$ defines an action on the space $\operatorname{Pos}(V\otimes\mathbb{C})$ of positive definite Hermitian forms $H$ on $V\otimes\mathbb{C}$ by
$\operatorname{{\mathrm{conj}}}(H)(v_{1},v_{2})=\operatorname{{\mathrm{conj}}}(H(\operatorname{{\mathrm{conj}}}(v_{1}),\operatorname{{\mathrm{conj}}}(v_{2}))).$
The group $G^{\mathbb{C}}$ acts on $\operatorname{Pos}(V\otimes\mathbb{C})$ as well, and the $K^{\mathbb{C}}$ invariant hermitian forms
$\operatorname{Pos}(V\otimes\mathbb{C})^{K^{\mathbb{C}}}=\operatorname{Pos}(U)^{K^{\mathbb{C}}}\oplus\operatorname{Pos}(U^{\operatorname{{\mathrm{conj}}}})^{K^{\mathbb{C}}}.$
Each of these components is isomorphic to $\mathbb{R}^{>0}$ by Claim 8.3. Note that $\operatorname{Gal}(\mathbb{C}/\mathbb{R})$ acts on the action $K^{\mathbb{C}}\lefttorightarrow\operatorname{Pos}(V\otimes\mathbb{C})$, and when restricting this action to $\operatorname{Pos}(V\otimes\mathbb{C})^{K^{\mathbb{C}}}$ we get that $\operatorname{{\mathrm{conj}}}$ replaces the two isomorphic components. As an action of actions we can think of it as an action of the semidirect product, $\operatorname{Gal}(\mathbb{C}/\mathbb{R})\ltimes K^{\mathbb{C}}\lefttorightarrow\operatorname{Pos}(V\otimes\mathbb{C})$. The space of invariants of this action is $\mathbb{R}^{>0}$, and hence up to multiplication by positive scalar there is again only one such positive quadratic form. Since $\operatorname{{\mathrm{conj}}}$ preserves $K^{\mathbb{C}}$ in $G^{\mathbb{C}}$ we get an action of the compact group $\operatorname{Gal}(\mathbb{C}/\mathbb{R})\ltimes K^{\mathbb{C}}$ on $V^{\mathbb{C}}$. The rest of the proof is similar to case (1). ∎
The next example shows that case (2) in the proof of Lemma 8.1 can occur.
Example 8.4.
Let $G=\operatorname{SU}(3)$, $V=\mathbb{C}^{3}$, and $\varrho:G\rightarrow\operatorname{GL}(V)$ be the standard complex representation. Since $G$ acts transitively on the unit sphere in $V$, it is irreducible. Moreover, restriction of scalars $\varrho^{\prime}=\operatorname{Res}_{\mathbb{C}/\mathbb{R}}(\varrho)\in{\rm Rep}_{\mathbb{R}}(G)$ is also irreducible. However, the complexification of $\varrho^{\prime}$ is reducible, as $\varrho^{\prime}\otimes_{\mathbb{R}}\mathbb{C}=\varrho\otimes_{\mathbb{R}}\mathbb{C}$ is of complex dimension $6$ and has the $\mathbb{C}$-linear multiplication map to $\varrho=\varrho\otimes_{\mathbb{C}}\mathbb{C}$.