### v6

# Geometric interpretation of quantitative instability

Abstract.
Given a real algebraic group action on a linear space $G\lefttorightarrow V$, a vector $v\in V$ is called unstable if $0\in\overline{Gv}-Gv$, where the closure is taken with respect to the Zariski topology. A fundamental theorem of Kempf [21] in geometric invariant theory states that $v$ is unstable if and only if there is a one-parameter subgroup $A$ of $G$ such that $v$ is unstable with respect to it, i.e., $0\in\overline{Av}-Av$. Assuming $G$ is a semisimple real algebraic group defined over $\mathbb{Q}$, we give a new proof to this result using a geometric interpretation of the setting. In the process, we also give a new proof of an effective version of this result by Shah and Yang [30, Prop. 2.2]. Our interpretation involves relating the length of vectors under a linear action to convex functions on certain $\operatorname{CAT}(0)$-spaces, and bound the latter from below by Busemann functions.
1. Introduction
Geometric invariant theory studies the quotient of a variety $V$ by a real algebraic group action $G\lefttorightarrow V$. An important phenomenon occurs when an orbit closure of a point $v$ in $V$ contains a point which is not in the $v$-orbit, i.e, when $\overline{Gv}-Gv\neq\emptyset$, where the the topology considered is the Zariski topology. It is then a natural question to study the behavior of the orbit around such point. In [21] Kempf studied the following question:
Given $v\in V$, $u\in\overline{Gv}-Gv$, for which curves $\{\gamma(t)\}$ in $G$ does the distance $\|u-\gamma(t)v\|$ shrink ‘fastest’?
The phenomenon of non-closed $G$-orbits is called instability. In homogeneous dynamic the $\{0\}$-instability, i.e., the case $u$ is the zero vector in a $G$-representation $V$, is widely used when studying how far an orbit is into a cusp, see [24, 19, 29, 31]. Recently, Kempf’s result has been used in the study of orbits and more specifically orbit closures in homogeneous dynamics. First by Yang [39], which used it to prove not only nondivergence, but also equidistribution of certain measures. Yang’s usage of [21] is extending Shah’s basic lemma [29, Prop. 4.2], which was proved independently of Kempf’s result. In [18, 30] the method was pushed further, and a quantitative version of Kempf’s result was proved, answering a seemingly different question:
Assuming $0\in\overline{Gv}$, can the size $\left\|gv\right\|$ for all $g\in G$ be controlled?
The solution they provide is using highest weight representations:
Theorem 1.1 ([30, Prop. 2.2]).
Assume $G$ is semisimple and let $\varrho:G\to\operatorname{GL}(V)$ be a $\mathbb{Q}$-representation and $0\neq v\in V(\mathbb{Q})$ be an unstable vector, i.e., $0\in\overline{Gv}-Gv$. Then, there exists a $\mathbb{Q}$-highest weight representation $\tilde{\varrho}:G\rightarrow\operatorname{GL}(W)$, a highest weight vector $w\in W(\mathbb{Q})$, and $a,c>0$ such that such that for all $g\in G$
(1.1) $\log\|\varrho(g)v\|\geq a\log\|\tilde{\varrho}(g)w\|-c.$
Remark 1.2.
Note that the highest weight in Theorem 1.1 is chosen with respect to some choice a maximal $\mathbb{Q}$ split torus and a choice of $\Delta_{\mathbb{Q}}$ (see §2.6).
This theorem is important for their setting, but it has another application previously unknown, as it answers a question of Weiss, [38, Question 3.4] positively. In this paper we will prove Theorem 1.1 using geometric tools, and give a geometric interpretation to the right hand side of Eq. (1.1).
1.1. Main results
1.1.1. Geometric interpretation
[…]
Theorem 1.3.
Let $\varrho:G\rightarrow V$ be a $\mathbb{R}$-representation and assume a $K$-invariant norm on $V$. Then, for every $v\in V$ there exists a Busemann function $\beta$ and constants $a,c>0$ so that we have the following:
  1. (1)
    Any $x\in M$ satisfies
    $f_{v}(x)\geq a\beta(x)-c.$
  2. (2)
    If $v\in V$ is a highest weight vector, then for any $x\in M$
    $f_{v}(x)=a\beta(x)-c.$
Remark 1.4.
In this manuscript we only deal with the real and rational setting (see §6 for the rational setting), but similar results should hold for a general field, using more general theory, see [35, 4].
1.1.2. Algebraic properties
Theorem 1.3 lacks the rational nature of Theorem 1.1. To this end we give various algebraic descriptions for ‘rational’ Busemann functions. The equivalent descriptions for Busemann functions can be found at Theorem 3.7 and the equivalent description for ‘rational’ Busemann functions can be found in Theorem 6.7. Each of these theorems has some equivalent description, and in particular contains the second part of Theorem 1.3. The first is the standard definition of the Busemann function, in Theorem 3.7 and a restriction of the geodesic in Theorem 6.7. The second uses homomorphisms from a parabolic subgroup. The third uses a linear combination of functions of the form $f_{v}$, for $v$ in a special collection of representations, called fundamental representations, (see Definition 2.21). Theorem 6.7 has a forth equivalent description, which is the right hand side of Eq. (1.1).
1.2. On the proof
Analyzing Busemann functions: In §3 we show the connection between the Busemann functions and the fundamental representations. In particular, we prove Theorem 3.7 which implies Theorem 1.3 Part (2).
Convex geometry: In §4-5 we prove Theorem 5.1 which implies Theorem 1.3 Part (1). We investigate the geodesic paths in $G$ that ‘shrink’ $v$ the most, and prove Theorem 1.3 Part (1). It has two key steps. In §4 we find a ‘fastest shrinking geodesic’ for all ‘nice’ convex functions. In particular, we are able to find the geodesic that shrinks a given vector in a representation the fastest. This part is a geometric analogous to Kempf’s result [21].
In §5 we show that the function defined by the left hand side of (1.1) is indeed ‘nice’, and use some of its algebraic properties to show that it can be bounded below by a (specific) Busemann function. Theorem 5.1 is the main theorem of this part. This part provides a geometric way to view [21], and a good intuition for Theorem 5.1.
Algebraic view point: In the previous part we have constructed a geometric object, the Busemann function, (which is also algebraic by Theorem 6.7) given algebraic data, a vector in an algebraic representation. Next, we wish to say that since the vector is defined over $\mathbb{Q}$, it is invariant under $\operatorname{Gal}(\mathbb{C}/\mathbb{Q})$, hence so does the Busemann function, i.e., it is also defined over $\mathbb{Q}$. This requires us to find an algebraic description for the fastest shrinking geodesic of a vector. Such an algebraic description was defined and proved by Kempf in [21]. In §6 we relate our constructions to Kempf’s construction, and use his result to deduce that our ‘fastest shrinking geodesic’ is indeed algebraic. Thus, the inequality in (1.1) is satisfied for the rational fundamental representations (and not only the real ones).
In §7 we give an alternative proof of Theorem 1.1 using a more precise version of Theorem 1.3.
1.3. Further research
Theorem 5.1 gives a lower bound using a Busemann function for every convex function on a symmetric space of noncompact type $f:M\to\mathbb{R}$ coming from a vector in a representation. This yields the following conjecture:
Conjecture 1.5.
For every non-constant convex function $f$ on $M$, there is a Busemann function $\beta:M\to\mathbb{R}$, $a>0,C\in\mathbb{R}$ such that for every $x\in M$
$f(x)>a\beta(x)+C.$
The anlagous conjecture for trees is false: Let $G$ be a leafless $d$-regular tree, $x_{0}\in G$ a vertex and $(x_{i})_{i=0}^{\infty}$ an infinite ray. Let $f:G\to\mathbb{R}$ be the function defined as follows: $f(x_{i})=-\sqrt{i}$ for $i\geq 0$, and for other $x\in G$ let $x_{i_{0}}$ be the closest element to $x$ on the ray. Define $f(x):=-\sqrt{i_{0}}+d(x,x_{i_{0}})(\sqrt{i+1}-\sqrt{i})$. One can verify that it is convex and not bounded from below by any multiple of a Busemann function.
Failures to extend this example to symmetric spaces leads us to believe the conjecture.
Acknowledgment
We would like to thank B. Weiss for suggesting this problem and for familiarizing us with [21]. We would also like to thank P. Yang for bringing our attention to the new results in this topic. The first author would like to thank N. Tur, O. Bojan, and A. Levit for helpful discussions. The second author would also like to thank R. Spatzier for his interest in this project and for several helpful discussions and comments.
2. Preliminaries
2.1. Homomorphisms
We will use several types of homomorphisms, Algebraic ones and topological ones. The set of continuous homomorphisms between topological groups will be denoted by $\operatorname{Hom}$. The set of $\ell$-algebraic homomorphisms between $\ell$-algebraic groups will be denoted by $\operatorname{Hom}_{\ell}$.
2.2. Hadamard spaces
A Hadamard space is defined to be a nonempty complete metric space $(M,d)$ such that, given any points $x,y\in M$, there exists $m\in M$ such that for every $z\in M$ we have
(2.1) $d(z,m)^{2}+{\frac{d(x,y)^{2}}{4}}\leq{\frac{d(z,x)^{2}+d(z,y)^{2}}{2}}.$
The point $m$ is called the midpoint of $x$ and $y$, and it satisfies $d(x,m)=d(y,m)=d(x,y)/2$. we study the behavior of convex functions in such spaces. The main properties of them which are of use to us are listed in Lemma 2.2 below.
Alternatively, a space is Hadamard if it is a complete $\operatorname{CAT}(0)$-space. A metric space $(M,d)$ is a $\operatorname{CAT}(0)$-space if it is geodesic (as defined below) and every ‘small enough’ geodesic triangle satisfies a certain inequality. Such spaces were first defined and studied by Gromov, see [16]. For a more detailed discussion on complete and $\operatorname{CAT}(0)$-spaces see [6].
Definition 2.1 (Geodesic).
Let $(M,d)$ be a metric space. Given $x,y\in M$, a geodesic from $x$ to $y$ is a map $\gamma:I\rightarrow Y$, where $I=[a,b]\subset\mathbb{R}$ is a closed interval, such that $\gamma(a)=x$, $\gamma(b)=y$ and
$d(\gamma(s),\gamma(s^{\prime}))=|s-s^{\prime}|$
for all $s,s^{\prime}\in I$ (in particular, d(x,y)=bad(x,y)=b-a). By abuse of notations, we identify geodesics with their images. A geodesic ray in $M$ is a map $\gamma:[0,\infty)\rightarrow M$ such that $d(\gamma(s),\gamma(s^{\prime}))=|s-s^{\prime}|$ for all $s,s^{\prime}\geq 0$. $(M,d)$ is said to be a geodesic space if every two points in $M$ are joined by a geodesic.
Lemma 2.2.
Let $(M,d)$ be a Hadamard space. Let $\gamma_{1},\gamma_{2}:[0,\infty)\to M$ be geodesic rays such that $\gamma_{1}(0)=\gamma_{2}(0)$. Then:
  1. (1)
    For every $s>t>0$ we have
    $\frac{d(\gamma_{1}(s),\gamma_{2}(s))}{s}\geq\frac{d(\gamma_{1}(t),\gamma_{2}(t))}{t}.$
  2. (2)
    In particular, for $r:=d(\gamma_{1}(1),\gamma_{2}(1))$ and all $t\geq 1$, we have
    $d(\gamma_{1}(t),\gamma_{2}(t))\geq rt.$
  3. (3)
    For every $t\geq 1$, the midpoint $m_{t}$ of $\gamma_{1}(t)$ and $\gamma_{2}(t)$ satisfies
    $d(\gamma_{1}(0),m_{t})\leq\sqrt{1-r^{2}/4}\cdot t.$
Proof.
According to [6, Prop. 2.2] the distance function in a $\operatorname{CAT}(0)$-space is convex, i.e., for any $s>t>0$
$d(\gamma_{1}(t),\gamma_{2}(t))$ $\leq\frac{t}{s}d(\gamma_{1}(s),\gamma_{2}(s))+\left(1-\frac{t}{s}\right)d(\gamma_{1}(0),\gamma_{2}(0)).$
Since we assume $\gamma_{1}(0)=\gamma_{2}(0)$, Claim (1) follows. Claim (2) follows directly from Claim (1). By (2.1) and Claim (2) we have
$d(\gamma_{1}(0),m_{t})$ $\leq\sqrt{d(\gamma_{1}(0),\gamma_{1}(t))^{2}/2+d(\gamma_{2}(0),\gamma_{2}(t))^{2}/2-d(\gamma_{1}(t),\gamma_{2}(t))^{2}/4}$
$\leq\sqrt{t^{2}-(rt/2)^{2}},$
proving Claim (3). ∎
Definition 2.3 (Convex functions).
A function $f:M\to\mathbb{R}$ is convex if for every geodesic $\gamma:I\to M$ the composition $f\circ\gamma$ is a convex function on the interval $I$.
2.3. Symmetric spaces of non-compact type
There are many results on symmetric spaces of non-compact type, i.e., Riemannian manifolds of non-positive sectional curvature, whose group of symmetries contains an inversion symmetry about every point. In particular, such spaces are Hadamard spaces. Here we present some of their geometric properties as well as some explicit constructions for future use.
2.3.1. Overview
[…]
We are also interested in a specific algerbraic subgroup, $A\subset G$, which is a maximal real split torus, i.e., $A\cong(\mathbb{R}^{\times})^{r}$, were $r$ is the real rank of $G$. The subgroup $A$ is called the Cartan torus and has certain compatibility conditions with $K$, see Definition 2.10. Let $\mathfrak{a}\subset\mathfrak{g}$ denote the Lie algebra of $A$.
The metric structure on $M$ has some phenomenal properties, which we now describe.
2.3.2. Metric properties
Geodesics in symmetric spaces can be generalized into the higher dimensional concept of flats:
Definition 2.4 (Flats, geometric description).
A map $a:\mathbb{R}^{k}\to M$ is called flat if the pullback metric on $\mathbb{R}^{k}$ is the Euclidean metric. We abuse notations and do not distinguish between $a$ and its image $\operatorname{Im}(a)$. A flat is called maximal if no other flat properly contains it. In our setting, the dimension of all maximal flats of $M$ is equal to the $\mathbb{R}$-rank of $G$.
Theorem 2.5 (Algebraic description of flats, [6, Rem. 10.60(5)]).
For every Cartan torus $A$, all maximal flats can be described as translations of $\pi(A)$. In particular, for every flat $a:\mathbb{R}^{k}\to M$ there is $g\in G$ and a continuous homomorphism $\iota:\mathbb{R}^{k}\to A$ such that $a(x)=\pi(\iota(x)g)$.
Theorem 2.5 also implies an explicit description of the geodesics in $M$, as can be seen in the next chapter.
Claim 2.6.
The group $K$ acts transitively on maximal flats containing $\pi(e)$. Moreover, for every geodesic $\gamma$ containing $\pi(e)$ the subgroup $\operatorname{\mathrm{stab}}_{K}(\gamma)$ acts transitively on the set of maximal flats which contain $\gamma$.
Proof.
The group $G$ acts transitively on the maximal flats by [6, Rem. 10.60]. Moreover, since $A$ acts transitively on $\pi(A)$, $G$ acts transitively on pairs $(p,F)$ of a point $p$ in a maximal flat $F$. Since $K$ is the stabilizer of $\pi(e)$, we deduce that it acts transitively on the set of maximal flats containing $\pi(e)$.
As for the second part, assume that $\gamma(t)=\pi(\exp(\mathtt{a}t))$ for some $\mathtt{a}\in\mathfrak{a}$. Note that
$G^{\prime}:=\operatorname{Stab}_{G}(\mathtt{a})=\{g\in G:Ad_{g}(\mathtt{a})=\mathtt{a}\},$
is a reductive subgroup, as defined in [6, Def. 10.56]. We now remark that the first part of the proof, as well as the construction of the symmetric space in [6, Thm. 10.58], follow for reductive subgroups, and not only semisimple groups. As defined in [6, Thm. 10.58], the symmetric space $M^{\prime}$ of $G^{\prime}$ is a subspace of $M$, which contains all maximal flats of $M$ which contains $\gamma$. The second part of the claim now follows from the first part, applied to $M^{\prime}$. ∎
The following corollary is an algebraic analog to Claim 2.6. Recall that a Cartan algebra is a maximal Abelian subalgebra.
Corollary 2.7.
The group $K$ acts transitively on the collection of Cartan algebras. For every $\mathtt{a}\in\mathfrak{g}$ the stabilizer $\operatorname{\mathrm{stab}}_{K}(\mathtt{a})$ acts transitively on the collection of Cartan Algebras containing $\mathtt{a}$. ∎
2.3.3. Explicit construction
[…]
$\pi(g)=og,$
for any $g\in G$. Then, $G$ acts on $M$ by right multiplication.
Definition 2.8 (Cartan decomposition of $\mathfrak{g}$).
Recall that we fixed a maximal compact subgroup $K\subseteq G$. Denote its Lie algebra by $\mathfrak{k}=\operatorname{Lie}(K)$. The Killing form $B_{\mathfrak{g}}(\cdot,\cdot)$ is negative definite on $\mathfrak{k}$, and positive definite on its orthogonal complement $\mathfrak{p}=\mathfrak{k}^{\perp}$.
The decomposition $\mathfrak{g}=\mathfrak{k}\oplus\mathfrak{p}$ is the Cartan decomposition of $\mathfrak{g}$. Note that the adjoint action of $K$ on $\mathfrak{g}$ preserves $\mathfrak{k}$ and $B_{\mathfrak{g}}$, and hence preserves $\mathfrak{p}$ as well.
Definition 2.9 (Metric on $M$).
Fixing $o=[K]\in M$ which is stabilized by $K$, note that $T_{o}M=\mathfrak{g}/\mathfrak{k}\cong\mathfrak{p}$, thus define the positive definite bilinear form $B_{o}$ on $T_{o}M$ to be the the restriction of $B_{\mathfrak{g}}$ to $\mathfrak{p}$. Since $B$ is $K$ invariant, we may use the $G$ action on $M$ and define a Riemannian metric $B_{p}$ on $T_{p}M$ for every $p\in M$. As usual in Riemannian geometry, for every curve segment $\gamma:[0,1]\rightarrow M$ we define the arc length of $\gamma$ by
(2.2) $L(\gamma)=\int_{0}^{1}B_{\gamma(t)}(\dot{\gamma}(t),\dot{\gamma}(t))^{1/2}dt,$
and the metric
$d_{M}(p,q):=\inf\{L(\gamma)\ \ |\ \ \gamma:[0,1]\to M,\gamma(0)=p,\gamma(1)=q\}.$
Definition 2.10 (Cartan decomposition of $G$).
Fix a maximal abelian $\mathfrak{a}\subseteq\mathfrak{p}$. Then $A=\exp\mathfrak{a}$ is a maximal split torus in $G$, and one may write (see for example [22, Thm. 7.39])
(2.3) $G=KAK.$
In this case $A$ is called Cartan torus.
The next result follows directly from Theorem 2.5 (see also [12, 2.4.2]).
Corollary 2.11.
All geodesics starting at $\pi(e)$ are of the form $t\mapsto\pi(\exp(t\mathtt{a})g)$ for some $\mathtt{a}\in\mathfrak{a}$ with $B(\mathtt{a},\mathtt{a})=1$ and $g\in K$. Alternatively, such geodesic is of the form $t\mapsto\pi(\exp(t\mathtt{p}))$ for some $\mathtt{p}\in\mathfrak{p}$ with $B(\mathtt{p},\mathtt{p})=1$.
2.3.4. Example: $G=\operatorname{SL}_{n}(\mathbb{R})$
[…]
The Cartan decomposition is defined by $\mathfrak{p},\mathfrak{k}$ being the symmetric and antisymmetric matrices in $\mathfrak{sl}_{n}=\operatorname{Lie}(G)$. The tangent space $T_{p}P(n,\mathbb{R})$ at a point $p\in P(n,\mathbb{R})$ is naturally isomorphic (via translation) to $\mathfrak{p}$, and the pseudo-Riemannian structure there is defined by
$B_{p}(Y,Z)=\operatorname{Tr}(p^{-1}Yp^{-1}Z).$
The maximal $\mathbb{R}$-split torus $A$ here is the set of diagonal matrices. For more information see [6, §II.10].
2.4. Representations of $G$ and short vectors
Fix a subfield $\ell\subseteq\mathbb{R}$. We assume that $G$ is defined over $\ell$, and frequently use some of the algebraic structure. The reader should note that the Cartan decomposition is not necessarily defined over $\ell$. We use the following standard notation of arithmetic groups (see [2, 4]). Fix a subfield $\ell\subseteq\mathbb{R}$ and a maximal $\ell$-split torus $A_{\ell}$ in $G$, and denote its Lie-algebra by $\mathfrak{a}_{\ell}$. We can conjugate $K$ and obtain that $\mathfrak{a}_{\ell}\subset\mathfrak{p}$. Note that $\mathfrak{p}$ is not necessarily defined over $\ell$.
Given an $\ell$-representation $\varrho:G\rightarrow\operatorname{GL}(V)$, we denote by $\Phi_{\varrho}$ the set of $\ell$-weights of $G$, i.e., the set of characters $\lambda\in\mathfrak{a}^{*}_{\ell}$ such that the subspace
$V_{\lambda}=\left\{v\in V:\text{for all }a=\exp(\mathtt{a})\in A_{\ell},\>\varrho(a)v=\exp(\lambda(\mathtt{a}))v\right\}$
is not trivial. The space $V_{\lambda}$ is called the weight space corresponding to $\lambda$, and elements of $V_{\lambda}$ are called weight vectors corresponding to $\lambda$. Then, there is a decomposition
(2.4) $V=\bigoplus_{\lambda\in\Phi_{\varrho}}V_{\lambda}.$
The following lemma defines a ‘nice’ quadratic form on $V$ which we use to define a norm on it. It is proved in the Appendix.
Lemma 2.12 (Construction of bilinear form).
If $\ell=\mathbb{R}$ and $A$ is a cartan torus, then there is a $K$-invariant positive bilinear form $\left\langle\cdot,\cdot\right\rangle$ on $V$ so that the linear spaces $V_{\lambda}$ are orthogonal with respect to it.
Since $\mathfrak{a}_{\ell}\subset\mathfrak{p}$, we deduce that $A_{\ell}$ is contained in some Cartain torus, and hence we may apply Lemma 2.12 also for $A_{\ell}$.
2.5. Parabolic subgroups and their properties
We continue to use the notation of §2.4 We use standard notation and results about parabolic groups in symmetric spaces (see [4, §11], [12, §2.17], or [6, §II.10]). We also prove some properties of them to be used in later chapters.
The classical, algebraic, definition of a parabolic group is a closed subgroup $P$ of $G$ so that $G/P$ is a projective variety. In our view of $G$, as a group acting on a symmetric space, the following, more geometrical, definition of a parabolic group is more informative. See [4, Cor. 11.2] and [25, Prop. 2.6].
Definition 2.13 (Parabolic and Unipotent subgroups).
A subgroup $P$ is called parabolic if it is of the form
$P=P_{\mathtt{a}}:=\left\{g\in G:\lim_{t\rightarrow\infty}\exp(-t\mathtt{a})g\exp(t\mathtt{a})\text{ exists}\right\},$
where $\mathtt{a}\in\mathfrak{p}$. If $G$ is defined over $\ell$ then the group $P$ is called $\ell$-parabolic if it is defined over $\ell$ as an algebraic group. It may be an $\ell$-parabolic even if $\mathtt{a}$ is not $\ell$-algebraic. There are many different elements $\mathtt{a}$ defining the same $P_{\mathtt{a}}$.
A Borel subgroup of $G$ is a maximal closed, connected, solvable, subgroup of $G$. The unipotent radical of a parabolic group $P$ is the unipotent part of the intersection of all Borel subgroups which are contained in $P$. Explicitly, the unipotent radical of $P_{\mathtt{a}}$ is
$U_{\mathtt{a}}=\left\{g\in G:\lim_{t\rightarrow\infty}\exp(-t\mathtt{a})g\exp(t\mathtt{a})=e\right\}.$
The unipotent radical is always nilpotent, and if $P$ is an $\ell$ parabolic then the unipotent radical is also defined over $\ell$. Moreover, the unipotent radical depends only on $P_{\mathtt{a}}$ and not on $\mathtt{a}$.
The following subset of the semisimple part of the Levi decomposition of $P_{\mathtt{a}}$ is of special interest for us
(2.5) $T_{\mathtt{a}}$ $:=\exp\{\operatorname{\mathrm{stab}}_{\mathfrak{g}}(\mathtt{a})\cap\mathfrak{p}\},$
where $\operatorname{\mathrm{stab}}_{\mathfrak{g}}(\mathtt{a})=\{\mathtt{b}\in\mathfrak{g}:[\mathtt{a},\mathtt{b}]=0\}$. In particular, $T_{\mathtt{a}}\subset\operatorname{\mathrm{stab}}_{G}(\mathtt{a})$. Note that $T_{\mathtt{a}}$ is not a group. Let $K_{\mathtt{a}}:=K\cap P_{\mathtt{a}}$.
Lemma 2.14 (Generalized Iwasawa decomposition, [12, Prop. 2.17.5]).
For any $\mathtt{a}\in\mathfrak{p}$ we have $P_{\mathtt{a}}=K_{\mathtt{a}}\cdot T_{\mathtt{a}}\cdot U_{\mathtt{a}}$ and $G=K\cdot T_{\mathtt{a}}\cdot U_{\mathtt{a}}$. Moreover, in both equations the indicated decomposition is unique. It follows that $G=KP_{\mathtt{a}}$.
Remark 2.15.
Our symbols $\mathtt{a},P_{\mathtt{a}},U_{\mathtt{a}},T_{\mathtt{a}},K_{\mathtt{a}}$ correspond to the symbols $X,G_{X},N_{X},A_{X},K_{X}$ in [12].
Claim 2.16.
Let $P=P_{\mathtt{a}}$ for some $\mathtt{a}\in\mathfrak{p}$. Let $G\lefttorightarrow V$ be a representation of $G$ and $v\in V$ be a vector such that $\mathbb{R}v$ is $P$-invariant. Then $v$ is $U_{\mathtt{a}}$-invaraint.
Proof.
Since $\exp(t\mathtt{a})\in P$ for every $t\in\mathbb{R}$, we deduce that $\exp(t\mathtt{a})v=\lambda_{t}v$ for some $\lambda_{t}\in\mathbb{R}$. Hence, for every $u\in U_{\mathtt{a}}$ we have
$uv=\exp(-t\mathtt{a})u\exp(t\mathtt{a})v\xrightarrow{t\to\infty}v.$
Example 2.17 (Example 2.3.4 continued).
Assuming $G=\operatorname{SL}_{n}(\mathbb{R})$, every parabolic subgroup is the stabilizer of a flag $0<V_{1}<\cdots<V_{s-1}<V_{s}=\mathbb{R}^{n}$ in the space of flags. Thus, every parabolic subgroup is conjugated to a group consisting of all block upper triangular matrices.
For example, when $n=3$, there are, up to conjugation, three proper parabolic subgroups:
[…]
where the leftmost one is a minimal parabolic.
2.6. The fundamental representation
In this section, we keep the setting and notation of §2.4 and §2.5. We also follow standard notation and results (see [14, 22, 5]).
The Killing form defines an inner product on $\mathfrak{a}$ (and so also on $\mathfrak{a}^{*}$), which we denote by $\langle\cdot,\cdot\rangle$.
Definition 2.18 (The Root system and Weyl group).
Let $\mathfrak{a}_{\ell}$ be a maximal $\ell$-split torus of $\mathfrak{g}$. We denote by $\Phi_{\ell}$ the root system of $\mathfrak{a}_{\ell}$, i.e. the set of non-trivial eigenvalues with respect to the adjoint action of $\mathfrak{a}_{\ell}$ on $\mathfrak{g}$, and by $W(\Phi_{\ell})$ the Weyl group of $\Phi_{\ell}$, i.e. the group generated by the reflections $w_{\lambda}$, $\lambda\in\Phi_{\ell}$, defined by
(2.6) $w_{\lambda}(\chi)=\chi-2\frac{\langle\chi,\lambda\rangle}{\langle\lambda,\lambda\rangle}\lambda,$
for any characters $\chi\in\mathfrak{a}^{*}_{\ell}$.
Lemma 2.19 ([22, Prop. 2.68]).
For any $\chi\in\mathfrak{a}^{*}_{\ell}$ there exists an element $w\in W_{\ell}$ so that for every $\alpha\in\Delta_{\ell}$
$\langle w(\chi),\alpha\rangle\geq 0.$
Remark 2.20.
By [4, §11.19], the Weyl group can be realized as
$N_{G}(A_{\ell})/Z_{G}(A_{\ell}),$
where $N_{G}(A_{\ell})$ is the normalizer of $A_{\ell}$ in $G$ and $Z_{G}(A_{\ell})$ is the centralizer of $A_{\ell}$ in $G$. Moreover, if $\ell=\mathbb{R}$ and $A_{\ell}$ is the Cartan torus, [22, §IV.6] shows that the representatives of $W_{\ell}$ in $N_{G}(A_{\ell})/Z_{G}(A_{\ell})$ can be chosen to be from $K$.
Recall that we denote $\mathfrak{a}=\mathfrak{a}_{\mathbb{R}}$. For every $\mathbb{R}$-Weyl chamber $\mathfrak{a}^{\circ}\subset\mathfrak{a}_{\ell}-\bigcap_{w\in W(\Phi_{\mathbb{R}})}\ker(w)$, and any $\mathtt{a}\in\mathfrak{a}^{\circ}$, the group $U_{\mathtt{a}}$ depends only on $\mathfrak{a}^{\circ}$ and not on $\mathtt{a}$. It is denoted $N_{\mathfrak{a}^{\circ}}$. In particular, as a special case of Lemma 2.14 we have the decomposition
(2.7) $G=K\cdot A\cdot N_{\mathfrak{a}^{\circ}}.$
Definition 2.21 (The simple system and the highest weight).
See [5] for the definitions and claim below. Let $\Delta_{\ell}=\{\alpha_{1},\dots,\alpha_{r}\}$ be an $\ell$-simple system of $\Phi_{\ell}$. Given an $\ell$-representation $\varrho:G\rightarrow\operatorname{GL}(V)$ and $\lambda\in\Phi_{\varrho}$, we say that $\lambda$ is the $\ell$-highest weight of $\varrho$ if for any $\lambda^{\prime}\in\Phi_{\varrho}$ we have
$\lambda-\lambda^{\prime}\in\operatorname{span}_{\mathbb{N}\cup\{0\}}\Delta_{\ell}.$
Note that a different choice of simple system yields a different highest weight.
The next result follows from the construction in [3, §7].
Lemma 2.22.
For any irreducible $\ell$-representation $\varrho$ there exists an $\ell$-highest weight.
Lemma 2.23.
Let $\varrho:G\rightarrow\operatorname{GL}(V)$ be an $\ell$-representation of $G$, $P$ be a parabolic subgroup of $G$ and $v\in V$ satisfy that $\varrho(P)v\subseteq\mathbb{R}v$. Then, there exists an irreducible sub-representation $\varrho^{\prime}:G\rightarrow\operatorname{GL}(V^{\prime})$ of $\varrho$ for some $V^{\prime}<V$ so that $v$ is its $\ell$-highest weight vector with respect to some choice of a simple system.
Proof.
Let $V^{\prime}$ be the subspace of $V$ spanned by $\varrho(G)v$ and $\varrho^{\prime}:G\rightarrow\operatorname{GL}(V^{\prime})$ be the implied sub-representation of $\varrho$. If we show that $\varrho^{\prime}$ is irreducible, then the claim will follow from [3, §7].
Let $G^{\mathbb{C}}$ be the complexification of $G$, $\varrho^{\mathbb{C}}:G^{\mathbb{C}}\to GL(V^{\mathbb{C}})$ be the complexification of $\varrho$. It is enough to show that $\varrho^{\mathbb{C}}(G^{\mathbb{C}}).v$ generate an irreducible representation in $V^{\mathbb{C}}$. Let $B<P^{\mathbb{C}}$ be a Borel subgroup of $G^{\mathbb{C}}$. Let $\sigma:B\to\mathbb{C}^{\times}$ be the character of $B$ using its action on $\mathbb{C}v$. The construction of the Verma-module (see [22, §V.3]) provides us with the unique irreducible representation of $G^{\mathbb{C}}$ with a vector $v^{\prime}$, on which $B$ acts by $\sigma$, and concludes the proof. ∎
Let us look at the adjoint representation $\operatorname{Ad}:G\rightarrow\operatorname{Aut}(\mathfrak{g})$. As described in §2.4, it has the decomposition
$\mathfrak{g}=\bigoplus_{\lambda\in\Phi_{\ell}=\Phi_{\operatorname{Ad}}}\mathfrak{g}_{\lambda}.$
Definition 2.24.
For any $1\leq i\leq r$ let $P_{i}$ be the $\ell$-parabolic subgroup of $G$ with Lie algebra
$\operatorname{Lie}(P_{i})=\bigoplus_{\lambda\not\leq-\alpha_{i}}\mathfrak{g}_{\lambda}$
(see [22, §V.7]).
Remark 2.25.
For any $1\leq i\leq r$ the set $\{\lambda\in\Phi_{\ell}:\lambda\not\leq-\alpha_{i}\}$ is equal to the collection of $\lambda\in\Psi$ such that in the representation $\lambda=\sum_{j=1}^{r}c_{j}\alpha_{j}$ we have $c_{i}\geq 0$.
For any $1\leq i\leq r$ let
$\chi_{i}:=\sum_{\lambda\not\leq-\alpha_{i}}\lambda\dim\mathfrak{g}_{\lambda},\quad d_{i}:=\sum_{\lambda\not\leq-\alpha_{i}}\dim\mathfrak{g}_{\lambda}=\dim\operatorname{Lie}(P_{i}).$
Let $\bigwedge^{d_{i}}\mathfrak{g}$ be the implied wedge representation such that for any $g\in G$ and $\bigwedge_{j=1}^{d_{i}}v_{j}\in\bigwedge^{d_{i}}\mathfrak{g}$,
$\hat{\varrho}_{i}(g)\left(\bigwedge_{j=1}^{d_{i}}v_{j}\right)=\bigwedge_{j=1}^{d_{i}}\operatorname{Ad}(g)(v_{j}).$
Let $(v_{j,i})_{j=1}^{d_{i}}$ be a basis of $\operatorname{Lie}(P_{i})$. Then $v_{i}=\bigwedge_{j=1}^{d_{i}}v_{j,i}$ satisfies that $\mathbb{R}v_{i}$ is $P_{i}$ invariant. The $G$-sub-representations $\varrho_{i}:G\to\operatorname{GL}(V_{i})$ generated by $v_{i}$ is irreducible by Lemma 2.23, and is termed fundamental representations. The vector $v_{i}$ is of highest weight $\chi_{i}\in\Phi_{\varrho_{i}}$ in $V_{i}$.
The following result is known to experts, but as we did not find it in the literature, we prove it here. It shows that the representations constructed in this section are ‘almost’ the $\ell$-fundamental representations, which are the ones who satisfy the conclusion of Proposition 2.26 with minimal $m_{i}$.
Proposition 2.26.
For any $1\leq i\leq r$ the highest weight satisfies
(2.8) $\langle\alpha_{j},\chi_{i}\rangle=m_{i}\delta_{i,j},$
where $m_{i}$ is a positive integer and $\delta_{i,j}$ is Kronecker delta, for all $1\leq j\leq r$.
Proof.
In a similar fashion to [34, Lemma 5.1], for any $j\neq i$, $\chi_{i}$ is invariant under the action of $w_{\alpha_{j}}$ (see Definition 2.18), which implies (2.8) when $j\neq i$. Since $\chi_{i}$ is a non-zero, non-negative integer combination of the $\ell$-simple system, (2.8) must hold for $j=i$ as well. ∎
Definition 2.27 (Attaching a character to a torus element).
For any $\mathtt{a}\in\mathfrak{a}_{\ell}$ denote by $\chi_{\mathtt{a}}$ the character on $\mathfrak{a}_{\ell}$ which is defined by
$\mathtt{b}\mapsto\frac{\langle\mathtt{a},\mathtt{b}\rangle}{\langle\mathtt{a},\mathtt{a}\rangle}.$
Similarly, for any character $\chi\in\mathfrak{a}_{\ell}^{*}$ we denote by $\mathtt{a}_{\chi}$ the element in $\mathfrak{a}_{\ell}$ which satisfies $\chi=\chi_{\mathtt{a}_{\chi}}$.
Since $\{\chi_{1},\dots,\chi_{r}\}$ spans $\mathfrak{a}_{\ell}^{*}$, there exist $a_{1},\dots,a_{r}\in\mathbb{R}$ such that
(2.9) $\chi_{\mathtt{a}}=\sum_{i=1}^{r}a_{i}\chi_{i}.$
Moreover, fixing $\Delta_{\ell}$ determines a positive $\ell$-Weyl chamber by
$\mathfrak{a}_{\ell}^{+}:=\{\mathtt{a}:\forall\alpha\in\Delta_{\ell},\quad\alpha(\mathtt{a})>0\}.$
For any $\mathtt{a}$ there exists a choice of $\Delta_{\ell}$ so that $\mathtt{a}\in\mathfrak{a}_{\ell}^{+}$. In that case, the coefficients $a_{1},\dots,a_{r}$ in (2.9) are non-negative.
The next property of $\mathfrak{a}_{\ell}^{+}$ follows from Definition 2.21, (2.6), as the set of weights of a given representation is preserved under the action of the Weyl group.
Lemma 2.28.
If $\lambda\in\mathfrak{a}_{\ell}^{*}$ is the $\ell$-highest weight of some irreducible representation of $G$, then $\lambda\in\overline{\mathfrak{a}_{\ell}^{+}}$.
Claim 2.29.
For every $\mathtt{a}\in\overline{\mathfrak{a}^{+}_{\ell}}$, one can write $\mathtt{a}=\sum_{i=1}^{r}c_{i}\mathtt{a}_{\chi_{i}}$, with $c_{i}\geq 0$. Then $P_{\mathtt{a}}=\bigcap_{c_{i}>0}P_{i}$. In particular, $P_{i}=P_{\mathtt{a}_{\chi_{i}}}$.
Proof.
We prove the claim by showing the equality of the Lie algebras.
First, we show $\bigcap_{c_{i}>0}\operatorname{Lie}(P_{i})\subseteq\operatorname{Lie}(P_{\mathtt{a}})$. Let
$I_{\mathtt{a}}=\{\lambda\in\Phi_{\ell}:\forall c_{i}>0,\lambda\not\leq-\alpha_{i}\}.$
Then, by (2.8), for any $\lambda\in I_{\mathtt{a}}$ we have $\langle\lambda,\chi_{i}\rangle\geq 0$, which implies
(2.10) $\lambda(\mathtt{a})=\sum_{i=1}^{r}c_{i}\langle\lambda,\chi_{i}\rangle\geq 0$
Let $X\in\bigcap_{c_{i}>0}\operatorname{Lie}P_{i}$. Since $\bigcap_{c_{i}>0}\operatorname{Lie}P_{i}=\bigoplus_{\lambda\in I_{\mathtt{a}}}\mathfrak{g}_{\lambda}$, we may write $X=\sum_{\lambda\in I_{\mathtt{a}}}X_{\lambda}$, where for any $\lambda$ we have $X_{\lambda}\in\mathfrak{g}_{\lambda}$. Then, we can compute
(2.11) $\operatorname{Ad}(\exp(-t\mathtt{a}))(X)$ $=\sum_{\lambda\in I_{\mathtt{a}}}\exp(-t\lambda(\mathtt{a}))X_{\lambda}.$
By (2.10), for any $\lambda\in I_{\mathtt{a}}$, the value of $\lim\exp(-t\lambda(\mathtt{a}))$ is either 0 or 11, and so the above converges, implying $X\in\operatorname{Lie}(P_{\mathtt{a}})$.
To see that $\bigcap_{c_{i}>0}\operatorname{Lie}(P_{i})\supseteq\operatorname{Lie}(P_{\mathtt{a}})$, we note that by (2.8) for any $\lambda\in\Phi_{\operatorname{Ad}}\setminus I_{\mathtt{a}}$ we have $\langle\lambda,\chi_{i}\rangle<0$, which implies
$\lambda(\mathtt{a})=\sum_{i=1}^{r}c_{i}\langle\lambda,\chi_{i}\rangle<0.$
Hence, for a point $X$ with a non-trivial factor in $\bigoplus_{\lambda\in\Phi_{\operatorname{Ad}}\setminus I_{\mathtt{a}}}\mathfrak{g}_{\lambda}$, the sum in (2.11) does not converge. ∎
3. Busemann functions
In this section we use the notation and results of §2.3 and §2.5. We study a semisimple Lie group $G$ of noncompact type and the corresponding symmetric space $M$.
The main goal of this section is to show equivalent descriptions for Busemann functions. In particular, the second part of Theorem 1.3 follows from the last result of this section, Theorem 3.10.
Definition 3.1 (Busemann functions).
Let $d$ be a right-invariant Riemannian metric on $M$. Given a geodesic ray $\gamma:[0,\infty)\rightarrow M$, the function $\beta_{\gamma}:M\rightarrow\mathbb{R}$ defined by
(3.1) $\beta_{\gamma}(x)=\lim_{t\rightarrow\infty}(d(x,\gamma(t))-t)$
is called the Busemann function associated to $\gamma$ (see [6, §II.8.17]).
Example 3.2 (Busemann functions on Euclidean spaces).
Let $u\in\mathbb{R}^{\ell}$ be a vector with $\|u\|=1$. Then, the Busemann function of the geodesic $t\mapsto tu$ is $v\mapsto-\left\langle u,v\right\rangle$.
Definition 3.3 (The modular function).
Given a field $\ell$ and an $\ell$-algebraic group $P$, denote by $\delta_{P}:P\to\ell^{\times}$ the $\ell$-modular function associated with $P$. That is, assuming $\mu_{P}$ is the left Haar measure on $P$, for a Borel subset $S\subseteq P$ with positive $\mu_{P}$-measure and $h\in P$, we have
$\delta_{P}(h)=\frac{\mu_{P}(Sh^{-1})}{\mu_{P}(S)}$
(see[22, §VIII.2]). Recall that $\delta_{P}$ is a group homomorphism.
The following known result gives a simple description of the modular function associated with $P$, if $P$ is a Lie group.
Lemma 3.4 ([22, Prop. 8.27]).
If $P$ is a Lie group, then the modular function $\delta_{P}$ of $P$ is given by
$\delta_{P}(h)=|\det(\operatorname{Ad}_{P}(h))|,$
where $\operatorname{Ad}_{P}(h)$ is the adjoint action of $h$ on $\operatorname{Lie}(P)$.
Definition 3.5 (Positive homomorphism).
Let $P\subset G$ be a parabolic subgroup. A homomorphism $f:P\to\mathbb{R}$ is called positive if it is a linear combination with nonnegative coefficients of $\log\delta_{P^{\prime}}$ for maximal parabolic subgroups $P^{\prime}$ containing $P$.
This definition has an $\ell$-algebraic version for subfields $\ell\subseteq\mathbb{R}$. If $G$ is an $\ell$-algebraic semisimple group and $P$ an $\ell$-parabolic subgroup in it, then a homomorphism $f:P(\mathbb{R})\to\mathbb{R}$ is called $\ell$-positive if it is a linear combination with nonnegative and rational coefficients of $\log\delta_{P^{\prime}}$ for $\ell$-maximal parabolic subgroups $P^{\prime}$ containing $P$.
Example 3.6 (Example 2.3.4 continued).
For $1\leq m\leq n-1$ let us denote by $P_{m}$ the parabolic subgroup of $\operatorname{SL}_{n}(\mathbb{R})$ of block upper triangular matrices with block sizes $m$ and nmn-m. That is
$P_{m}=\left\{\begin{pmatrix}A&B\\ 0&C\end{pmatrix}:A\in M_{m,m},\>B\in M_{m,n-m},\>C\in M_{n-m,n-m}\right\},$
where $M_{k,l}$ denotes the set of $k\times l$ matrices with entries in $\mathbb{R}$. Then, the modular function $\delta_{P_{m}}$ satisfies
$\delta_{P_{m}}\left(\begin{pmatrix}A&B\\ 0&C\end{pmatrix}\right)=\frac{\det A^{n-m}}{\det C^{m}}=\det A^{n}.$
Hence, if $n=n_{1}+n_{2}+\cdots+n_{s}$ and $P$ is the parabolic subgroup of all upper diagonal block matrices with respect to the $(n_{i})_{i=1}^{s}$ block structure, then the positive homomorphisms consist of all maps $\sigma:P\to\mathbb{R}$ such that for some $c_{1}\geq c_{2}\geq\cdots\geq c_{s}$,
$\sigma(A)=\sum_{i=1}^{s}c_{i}\log|\det A_{i}|,$
where $A_{i}$ are the blocks of $A$ on the diagonal.
The main goal of this section is to prove the following theorem.
Theorem 3.7.
The following three classes of functions are equivalent:
  1. (1)
    Busemann functions: The class of constant shifts of nonnegative multiples of Busemann functions.
  2. (2)
    Homomorphisms from parabolics: The class of functions which satisfies that $\pi(p)\mapsto\chi(p)$ for all $p\in P$, where $P$ is a parabolic subgroup of $G$, $\chi:P\to\mathbb{R}$ is a positive homomorphims.
  3. (3)
    Lengths of highest weight vectors in fundumental representations: The class of functions of the form
    $\pi(g)\mapsto\sum_{i=1}^{r}c_{i}\log\|\varrho_{i}(gg_{0})v_{i}\|+C,$
    where $\varrho_{i}$ are $\mathbb{R}$-fundamental representations, $g_{0}\in G$, and $v_{i}$ are the appropriate highest weight vectors, which correspond to maximal $\mathbb{R}$-parabolic groups, and are defined in §2.6, for some choice of an $\mathbb{R}$-simple system, the $c_{i}$ are non-negative, and $C\in\mathbb{R}$.
For the proof of Theorem 3.7 we need a good description of the modular functions as well as of the kernel of the Busemann functions.
Lemma 3.8.
Let $1\leq i\leq r$. Then, for any $h\in P_{i}$,
$\varrho_{i}(h)v_{i}=\delta_{P_{i}}(h)v_{i}.$
Proof.
First, let $v_{i}$ be as in §2.6, a highest weight vector. Thus, the line spanned by $v_{i}$ is the highest weight eigenspace of $\varrho_{i}$, and is stabilized by $P_{i}$. That is, for any $h\in P_{i}$, the point $\varrho_{i}(h)v_{i}$ is a constant multiple of $v_{i}$.
Second, using Lemma 3.4 and a standard decomposition of parabolic groups (see [22, §V.7]), one can deduce that $|\det(\operatorname{Ad}_{P_{i}}(h))|$ is equal to the determinant of the adjoint action of $P_{i}$ on $\mathfrak{n}_{i}:=\bigoplus_{\langle\lambda,\chi_{i}\rangle>0}\mathfrak{g}_{\lambda}$. As the vector $v_{i}$ (as in §2.6) is defined to be the wedge product of a basis of $\mathfrak{n}_{i}$, the claim follows. ∎
Proof of Theorem 3.7.
First note that the equivalence between classes (2) and (3) follows directly from Lemma 3.8, and so it is enough to prove the equivalence between classes (1) and (3).
Let $\beta_{\gamma}$ be a Busemann function, i.e., there exist $\mathtt{a}\in\mathfrak{a}$ and $g\in K$ such that $\gamma$ is defined by $t\mapsto\pi(\exp(t\mathtt{a})g)$. Since both classes (1) and (3) are invariant under the $G$ action, we may apply $\operatorname{Ad}(g)^{-1}$ to replace the geodesic with $t\mapsto\pi(\exp(t\mathtt{a}))$, and assume that $g=I$. Moreover, by Lemma 2.19 and Remark 2.20, for some element $k\in K$ we may apply $\operatorname{Ad}(k)^{-1}$ to assume that $\mathtt{a}\in-\overline{\mathfrak{a}^{+}}$. Hence, one can represent $\mathtt{a}=\sum_{i=1}^{r}c_{i}\mathtt{a}_{\chi_{i}}$, with $\mathtt{a}_{\chi_{i}}$ as in Definition 2.27.
Let $C:=-\sum_{i=1}^{r}c_{i}\log\|v_{i}\|$, so that the desired equality
(3.2) $\beta_{\gamma}(\pi(g))=\sum_{i=1}^{r}c_{i}\log\|\varrho_{i}(g)v_{i}\|+C,$
holds for $g=I$. Since $\pi(A)$ is a flat, Example 3.2 implies that (3.2) holds for all $g\in A$.
[…]
$\beta_{\gamma}(xu)=\lim_{t\rightarrow\infty}(d_{M}(xu,\gamma(t))-t)=\lim_{t\rightarrow\infty}(d_{M}(x,\gamma(t)u^{-1})-t).$
Since
$\lim_{t\to\infty}d_{M}(\gamma(t)u^{-1},\gamma(t))=d_{M}(\pi(\exp(t\mathtt{a})u^{-1}\exp(-t\mathtt{a})),\pi(I))\xrightarrow{t\to\infty}0,$
we deduce that $\beta_{\gamma}$ is $U$ invariant. Similarly, by Claim 2.16, the right $U$-action preserves the right-hand side of Eq. (3.2).
Let $k\in\operatorname{\mathrm{stab}}_{K}\gamma$. Since $k$ preserves $\gamma$, we deduce that $k$ preserves $\beta_{\gamma}$. Since $\operatorname{\mathrm{stab}}_{K}\gamma\subseteq P$ we deduce that $k\in P$ and hence the right $k$ action preserves the right-hand side of (3.2).
Since the norms on $V_{i}$ are $K$ invariant, Eq. (3.2) depends only on $\pi(g)\in M$, and not on $g$.
Let $g\in G$ and represent $g=kup$, where $k\in K$, $u\in U_{-\mathtt{a}}$ and $p\in T_{-\mathtt{a}}$. Since $p\in T_{-\mathtt{a}}$, we deduce that $p=\exp(\mathtt{a}^{\prime})$ and $\mathtt{a}^{\prime}\in\mathfrak{p}$ commutes with $\mathtt{a}$. Let $\mathfrak{a}^{\prime}\subseteq\mathfrak{p}$ be a maximal abelian subalgebra containing $\mathtt{a},\mathtt{a}^{\prime}$. By 2.7, there is an element $k\in\operatorname{\mathrm{stab}}_{K}(\mathtt{a})$ such that $Ad_{k}(\mathfrak{a})=\mathfrak{a}^{\prime}$. Since (3.2) holds on $A=\exp(\mathtt{a})$, and is preserved by $\operatorname{\mathrm{stab}}_{K}(\mathtt{a})$, we deduce that (3.2) holds on $p$. Since (3.2) is preserved by multiplication from the left by $K$, and from the right by $p^{-1}up\in U$, we deduce that it holds on $g=kpp^{-1}up$. The other direction, if a function is of class (3) then it is also of class (1) follows from the same computation. ∎
Remark 3.9.
For a Busemann function $\beta_{\gamma}$ of a geodesic ray $\gamma:t\mapsto\pi(\exp(t\mathtt{a}))$, for some $\mathtt{a}\in\mathfrak{p}$, the parabolic subgroup in Class (2) is $P_{-\mathtt{a}}$.
Theorem 3.10.
[…]
Proof.
Fix a maximal $\mathbb{R}$-split torus $A\subset P$. By Lemma 2.23, since $P$ stabilizes $\mathbb{R}v$, $v$ is an $\mathbb{R}$-highest weight vector for $\varrho$ with respect to some choice of a simple system. Without loss of generality, we may assume that, in the notation of §2.6, the minimal $\mathbb{R}$-parabolic subgroup $P_{0}=\bigcap_{i=1}^{r}P_{i}$ is contained in $P$ (as otherwise it can be replaces with its conjugated to the one chosen here). Denote this $\mathbb{R}$-highest weight by $\chi$. Without loss of generality, we may assume $A$ is the one chosen in §2.6 (as it is conjugated to the one chosen here).
As defined in §2.4, $\chi$ is a character on $\mathfrak{a}$. Equivalently, $\chi$ can be viewed as a character on $A$. Recall that for any $a\in A$ we have $\varrho(a)v=e^{\chi(a)}v$. Since $\varrho(p_{1}p_{2})v=\varrho(p_{1})\varrho(p_{2})v$ and $\varrho(P)v=\mathbb{R}v$, we can expend $\chi$ to a character on $P$ which satisfies
$\varrho(p)v=e^{\chi(p)}v.$
By (2.9) and Lemma 2.28, we have $\chi=\sum_{i=1}^{r}c_{i}\chi_{i}$ for some choice of non-negative $c_{1},\dots,c_{d}$. That is,
$\varrho(p)v=e^{\chi(p)}v=e^{\sum_{i=1}^{r}c_{i}\chi_{i}(p)}v=\prod_{i=1}^{r}\delta_{P_{i}}^{c_{i}}(p)v,$
where the last equality follows from Lemma 3.8. ∎
Example 3.11 (Example 2.3.4 continued).
Assuming $G=\operatorname{SL}_{n}(\mathbb{R})$, for any $1\leq i\leq n-1$ the fundamental weight $\chi_{i}$ is a scalar multiple of the highest weight of the exterior product representation (on $\bigwedge_{i}\mathbb{R}^{n}$). In particular, one may use theis representation instead of $\varrho_{i}$, and obtain a simplification of Class (3) in Theorem 3.7: Let $u_{1},...,u_{n}$ be a basis of $\mathbb{R}^{n}$, $c_{1},...,c_{n-1}\geq 0$ and $C>0$. Consider the class of functions of the form
$\pi(g)\mapsto\sum_{i=1}^{n-1}c_{i}\log\|(gu_{1})\wedge(gu_{2})\wedge\cdots\wedge(gu_{i})\|+C.$
4. The fastest shrinking geodesic
To construct the Busemann function in Theorem 1.31), we use the ‘fastest shrinking geodesic’ of our function of interest. In this section, the fastest shrinking geodesic is constructed for a class of functions in the more general setting of a $\operatorname{CAT}(0)$-space, and the properties of such geodesic are studied in a special case, which is relevant to us, and is used in the next chapter.
Recall the definition (and relevant notations) of a Hadamard space from §2.2. Fix a locally compact, Hadamard space $(M,d)$ and a point $o\in M$.
For $s>0$, $x\in M$ denote by $B(x,s)$ the closed ball of radius $s$ around the point $x$.
Lemma 4.1.
Let $f:M\rightarrow\mathbb{R}$ be a convex function which is unbounded from below. Then, for any $x\in M$, $s>0$, the function $f$ attains a minimum on the closed ball $B(x,s)$ at a unique point on its boundary $\{y\in M:d(y,x)=s\}$.
Proof.
First, assume by contradiction that $f$ attains a minimum on $B(x,s)$ at a point $y_{1}$ with $d(e,y_{1})<s$. In particular, there is a neighbourhood of $y_{1}$ which is contained in $B(x,s)$. Hence, $y_{1}$ is a local minimum of $f$. Since $f$ is convex, this implies that $y_{1}$ is a global minimum. A contradiction to the assumption that $f$ is unbounded from below.
Second, assume by contradiction that $f$ attains its minimum on $B(x,s)$ at two points $y_{1},y_{2}$. Since $M$ is a $CAT(0)$ space, it follows that the midpoint $y_{3}$ of $y_{1}$ and $y_{2}$ satisfies $d(y_{3},x)<s$. Since $f$ is convex it follows that $f(y_{3})\leq f(y_{2})=f(y_{1})$. This implies that $y_{3}$ is another points in which $f$ attains a minimum, with $d(y_{3},x)<s$. A contradiction to the previous discussion. ∎
Fix a convex, unbounded from below, function $f:M\to\mathbb{R}$.
Definition 4.2 (Minimizing points of $f$ in balls).
For any $s>0$ denote by $x_{s}$ the point which minimizes $f$ on $B(o,s)$. By Lemma 4.1 the point $x_{s}$ is uniquely defined and satisfy $d(o,x_{s})=s$. In particular, this definition implies $x_{0}=o$. Denote by $\gamma_{s}:[0,s]\to M$ the geodesic connecting $o$ and $x_{s}$.
Lemma 4.3.
The sequence $\frac{f(x_{s})}{s}$ has a limit.
Proof.
Let us note that by the definition of $x_{s}$ and the convexity of $f$, for any $s<t$ we have
$\frac{f(x_{s})}{s}$ $\leq\frac{f(\gamma_{t}(s))}{s}\leq\frac{t-s}{t}\frac{f(\gamma_{t}(0))}{s}+\frac{s}{t}\frac{f(\gamma_{t}(t))}{s}$
$=\left(\frac{1}{s}-\frac{1}{t}\right)f(o)+\frac{f(x_{t})}{t}.$
That is, the shift g(x)=f(x)f(o)g(x)=f(x)-f(o) is a convex function, so that $\frac{g(x_{s})}{s}$ is negative and non-decreasing. Hence, the sequence $\frac{g(x_{s})}{s}$ has a limit, which implies that the sequence $\frac{f(x_{s})}{s}$ also has a limit. ∎
Definition 4.4.
We say that $f$ is decreasing linearly if
$\lim_{s\to\infty}\frac{f(x_{s})}{s}<0.$
Proposition 4.5.
If $f$ is decreasing linearly, then the sequence of geodesics $(\gamma_{s})_{s>0}$ constructed in Definition 4.2 converges to a geodesic $\gamma_{\infty}:[0,\infty)\to M$ pointwise. Moreover, setting $a:=-\lim_{s\to\infty}\frac{f(x_{s})}{s}>0$, the geodesic $\gamma_{\infty}$ is the unique geodesic which satisfies
$\gamma_{\infty}(0)=o\quad\text{and}\quad f(\gamma_{\infty}(s))=-as\cdot(1+o(1)).$
Proof.
Let
(4.1) $a:=-\lim_{s\to\infty}\frac{f(x_{s})}{s}.$
Since $f$ is decreasing linearly, we have $a>0$. Since the space of geodesic rays from $o$ is compact, the sequence $\{\gamma_{s}\}$ has a partial limit. Let $\gamma_{\infty}:[0,\infty)\to M$ be such a partial limit.
Since $f$ is convex, for any $s>0$ we have
$f(\gamma_{\infty}(s))$ $=\lim_{i\to\infty}f(\gamma_{s_{i}}(s))$
$\leq\lim_{i\to\infty}\left(\frac{s_{i}-s}{s_{i}}f(x_{0})+\frac{s}{s_{i}}f(x_{s_{i}})\right)$
$=f(x_{0})-s\cdot a,$
where the limit in the second line is taken over large enough $i$ so that $s_{i}>s$. Since $\gamma_{\infty}$ starts at $o$, we have that for any $s>0$, $d(\gamma_{\infty}(s),o)=s$. Hence, by Definition 4.2 for any $s>0$ we have
(4.2) $f(\gamma_{\infty}(s))\geq f(x_{s}).$
It now follows from (4.1) that for all $s>0$
$f(\gamma_{\infty}(s))=-as\cdot(1+o(1)).$
Thus, $f\circ\gamma_{\infty}$ is of the claimed form.
Let $\gamma_{\infty}^{\prime}:[0,\infty)\to M$ be another geodesic ray with $\gamma_{\infty}^{\prime}(0)=o$ and $f(\gamma_{\infty}^{\prime}(s))=-as\cdot(1+o(1))$ for all $s\geq 0$. Denote by $r$ the distance at time one between the two geodesics, i.e., $r:=d_{M}(\gamma_{\infty}(1),\gamma_{\infty}^{\prime}(1))$. By Lemma 2.2(2), $d_{M}(\gamma_{\infty}(t),\gamma_{\infty}^{\prime}(t))\geq rt$ for every $t\geq 1$. Denote by $m_{t}$ the middle point between $\gamma_{\infty}(t)$ and $\gamma_{\infty}^{\prime}(t)$. Then, by Lemma 2.2(3), we have
(4.3) $d_{M}(x_{0},m_{t})\leq t\sqrt{1-r^{2}/4}.$
On the other hand, since $f$ is convex it follows that $f(m_{t})\leq at(1+o(1))$, and hence
$f(x_{t\sqrt{1-r^{2}/4}})\leq-at(1+o(1)).$
Unless $r=0$, this together with (4.3) contradicts the definition of $a$. Hence $r=0$ and $\gamma_{\infty}=\gamma_{\infty}^{\prime}$. ∎
In view of Proposition 4.5 we define:
Definition 4.6 (Fastest shrinking geodesic).
Let $(M,o,f)$ be a tuple of a locally compact Hadamard space $M$, a point $o$ in $M$, and a convex, linearly decreasing function $f:M\to\mathbb{R}$. Define the fastest shrinking geodesic of $(M,o,f)$ to be $\gamma=\gamma_{M,o,f}:[0,\infty)\to M$, the limit of $\gamma_{s}$ defined as in Definition 4.2. Then, for $a_{M,o,f}:=-\lim_{s\to\infty}\frac{1}{s}\min_{x\in B(s,o)}f(x)$ we have $f(\gamma(s))=-a_{M,o,f}s\cdot(1+o(1))$ for all $s>0$. The constant $a_{M,o,f}$ is called the shrinking rate of f.
To analyze symmetric spaces (see §2.3) it is useful to consider the maximal flats in them. Hence, we show that restriction to ‘nice’ subspaces does not change the fastest shrinking geodesic or the shrinking rate.
Lemma 4.7 (Restriction).
Let $(M,o,f)$ be a triplet as in Definition 4.6. If $Y\subseteq M$ is a sub Hadamard space which contains the image of $\gamma_{M,o,f}$ then $\gamma_{M,o,f}=\gamma_{Y,o,f|_{Y}}$ and $a_{M,o,f}=a_{Y,o,f|_{Y}}$.
Proof.
We denote $\gamma:=\gamma_{M,o,f}$. By definition,
$a_{M,o,f}\leq a_{Y,o,f|_{Y}}\leq a_{\operatorname{Im}\gamma,o,f\circ\gamma}.$
Proposition 4.5 implies that $a_{\operatorname{Im}\gamma,o,f\circ\gamma}=a_{M,o,f}$. Hence, we deduce $a_{M,o,f}=a_{Y,o,f|_{Y}}$. Then, the uniqueness of the limit of geodesics in Proposition 4.5 implies that $\gamma_{Y,o,f|_{Y}}=\gamma$. ∎
The next result follows directly from Proposition 4.5.
Lemma 4.8 (Bounded shifts).
Let $(M,o,f)$ be a triplet as in Definition 4.6, and $\tilde{f}:M\to\mathbb{R}$ be a convex function such that |ff~||f-$\tilde${f}| is bounded. Then $(M,o,f)$ satisfies the assumptions of Definition 4.6 and
$\gamma_{M,o,f}=\gamma_{M,o,\tilde{f}},\quad a_{M,o,f}=a_{M,o,\tilde{f}}$.
4.1. Piecewise linear convex functions on Euclidean spaces
Our goal in defining the fastest shrinking geodesic is to use them to study, and more specifically bound the ‘shrink-rate functions’, which are defined in the next section. The restriction of these functions to maximal flats (see Definition 2.4) are piecewise linear, up to bounded error. Moreover, the metric on the maximal flats is Euclidean. Therefore, we dedicated this subsection to study the fastest shrinking geodesics in this setting.
In this subsection, we assume that $M$ is the Euclidean space $\mathbb{R}^{n}$ with the standard Euclidean metric.
Claim 4.9.
Let $V_{0}\subseteq\mathbb{R}^{n}$ and $(a_{v})_{v\in V_{0}}\subseteq\mathbb{R}^{n}$ be a collection of real numbers. Define the function
$f(x):=\max_{v\in V_{0}}(\left\langle x,v\right\rangle+a_{v}).$
Then, $f$ is unbounded from below if and only if
(4.4) $0\notin{\rm conv}(V_{0}).$
Moreover, assuming (4.4), and denoting by $u$ the closest point to 0 in ${\rm conv}(V_{0})$, we have:
  1. (1)
    $f$ decreases linearly,
  2. (2)
    The fastest shrinking geodesic for $f$ is
    $\gamma_{\mathbb{R}^{n},0,f}(t)=-tu/\|u\|,$
  3. (3)
    $a_{M,o,f}=-\|u\|$, and
  4. (4)
    There exists $C\in\mathbb{R}$, such that
    (4.5) $f(x)\geq\left\langle x,u\right\rangle+C.$
Proof.
If $0\in{\rm conv}(V_{0})$, then for some non-negative $(c_{v})_{v\in V_{0}}$ with $\sum_{v\in V_{0}}c_{v}=1$ we have $\sum_{v\in V_{0}}^{n}c_{v}v=0$. Then,
$f(x)=\max_{v\in V_{0}}(\left\langle x,v\right\rangle+a_{v})\geq\sum_{v\in V_{0}}^{n}c_{v}(\left\langle x,v\right\rangle+a_{v})=\sum_{v\in V_{0}}c_{v}a_{v}.$
is a bound from below on $f$.
Now, assume $0\notin{\rm conv}(V_{0})$. We show that $f$ decreases linearly, which also implies that $f$ is unbounded from below. By Definition 4.2, it is enough to find a direction in which $f$ decreasing linearly, i.e. a vector $v$, such that $f(sv)\leq cs$ for some $c<0$ and all large enough $s$.
Let $u$ be the closest point to 0 in the convex hull ${\rm conv}(V_{0})$. Since $u$ is of minimal length, for any $v\in V_{0}$ and $p\in(0,1)$ we have
$\left\|u\right\|^{2}$ $\leq\left\|pv+(1-p)u\right\|^{2}$
$=p^{2}\left\|v\right\|^{2}+2p(1-p)\langle u,v\rangle+(1-p)^{2}\left\|u\right\|^{2},$
which implies
$\langle u,v\rangle\geq\left\|u\right\|^{2}-\frac{p}{2(1-p)}(\left\|v\right\|^{2}-\left\|u\right\|^{2}).$
Thus, we deduce $\langle u,v\rangle\geq\left\|u\right\|^{2}$, for all $v\in V_{0}$. We may now compute
$f(-tu)=\max_{1\leq i\leq n}(-t\left\langle u,v\right\rangle+a_{v})\leq-t\|u\|^{2}+\max_{v\in V_{0}}a_{v},$
which decreases linearly, proving (1).
According to Proposition 4.5 there is a minimal shrinking geodesic and $a_{\mathbb{R}^{n},0,f}\leq-\|u\|$. On the one hand, for all $x\in\mathbb{R}^{n}$ we have
$f(x)=\max_{v\in V_{0}}(\left\langle x,v\right\rangle+a_{v})\geq\sum_{v\in V_{0}}c_{v}(\left\langle x,v\right\rangle+a_{v})=\left\langle x,u\right\rangle+\sum_{v\in V_{0}}c_{v}a_{v},$
which implies (4) and $a_{\mathbb{R}^{n},0,f}\geq-\|u\|$. Thus, (2) and (3) are also satisfied. ∎
The following lemma shows that the constant of the lower bound in (4.5) is (in a sense) continuous in the vectors which define $f$ and can be chosen to be uniform on compact sets.
Lemma 4.10.
Let $V_{0}\subseteq\mathbb{R}^{n}$ be a finite set, $K$ be a compact set, and $(r_{v})_{v\in V_{0}}$ be a collection of continuous functions, $r_{v}:K\to\mathbb{R}\cup\{-\infty\}$. Assume that for each $k\in K$ the function
$f_{k}:\mathbb{R}^{n}\to\mathbb{R},\quad f_{k}(x)=\max_{v\in V_{0}}(\left\langle x,v\right\rangle+r_{v}(k))$
is well defined (i.e. $f_{k}$ does not attain the value -$\infty$), and the point of minimal length for each $k$
$u\in{\rm conv}(\{v\in V_{0}:r_{v}(k)>-\infty\})$
is independent of $k$, i.e., a constant vector. Then, there exists $C\in\mathbb{R}$ such that for all $k\in K,x\in\mathbb{R}^{n}$ we have
(4.6) $f_{k}(x)\geq\left\langle x,u\right\rangle+C.$
Proof.
By the definition of $u$, the value
$\xi(k):=\max_{\begin{subarray}{c}V_{1}\subseteq V_{0}\\ u\in{\rm conv}(V_{1})\end{subarray}}\min_{v\in V_{1}}r_{v}(k),$
is not -$\infty$ for every $k\in K$. Since $\xi$ defines a continuous function $\xi:K\to\mathbb{R}$, it attains a minimum, which we denote by $C$.
We now show that for every $k\in K,\,x\in\mathbb{R}^{n}$, Equation (4.6) is satisfied. Indeed, for some $V_{1}\subseteq V_{0}$ with $u\in{\rm conv}(V_{1})$ we have $\min_{v\in V_{1}}r_{v}(k)\geq C$. Since $u\in{\rm conv}(V_{1})$, there exists a convex combination $\sum_{v\in V_{1}}c_{v}v=u$. Then,
$f_{k}(x)$ $=\max_{v\in V_{0}}(\left\langle x,v\right\rangle+r_{v}(k))\geq\sum_{v\in V_{1}}c_{v}(\left\langle x,v\right\rangle+r_{v}(k))$
$=\left\langle x,u\right\rangle+\sum_{v\in V_{1}}c_{v}r_{v}(k)\geq\left\langle x,u\right\rangle+C,$
as wanted. ∎
Remark 4.11.
Lemma 4.10 does not hold if one allows $V_{0}$ to depend on $K$. For example, by taking $K=[0,1]$, and for each $k\in K$,
$V_{k}=\begin{cases}\{(1,-k),(1,1)\}&\text{if }k>0,\\ \{(1,0)\}&\text{if }k=0,\end{cases}$
$r_{0,(1,0)}=0,\quad r_{k,(1,-k)}=0,\quad r_{k,(1,1)}=\begin{cases}\frac{1}{k^{2}}&\text{if }k>0,\\ 0&\text{if }k=0,\end{cases}$
we get that for the sequence of functions
$f_{k}:\mathbb{R}^{2}\to\mathbb{R},\quad f_{k}(x,y)=x+\begin{cases}\max\left(-ky,y-\frac{1}{k^{2}}\right)&\text{if }k>0,\\ 0&\text{if }k=0,\end{cases}$
the result of Lemma 4.10 fails.
The following Claim will be used in §6.
Claim 4.12.
Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be a convex and piecewise linear function with finitely many slopes. Then the map $v\mapsto\lim_{t\to\infty}\frac{f(tv)}{t}$ is continuous.
Proof.
Since $f$ is a piecewise linear function, there exist a finite set $V_{0}\subseteq\mathbb{R}^{n}$ and a collection of real numbers $(a_{v})_{v\in V_{0}}$ such that $f(x)=\max_{v\in V_{0}}(\left\langle x,v\right\rangle+a_{v})$ for every $v\in\mathbb{R}^{n}$. Then, for any $x\in\mathbb{R}^{n}$
$\lim_{t\to\infty}\frac{f(tx)}{t}=\max_{v\in V_{0}}\left\langle x,v\right\rangle,$
which is a continuous function. ∎
5. The shrink-rate function
Let $\varrho:G\rightarrow\operatorname{GL}(V)$ be an $\mathbb{R}$-representation, $v\in V$, and $\|\cdot\|$ be a norm on $V$ as in Lemma 2.12. The map $G\to\mathbb{R}$ defined by $g\mapsto\log\|\varrho(g)v\|$ is invariant under the right action of $K$. Hence, it defines a map $f_{v}:M\to\mathbb{R}$. The function $f_{v}$ is called the shrink-rate function of $v$.
Recall the representation theory notations from §2.4, the definition of the Busemann function from §3, and the definition of the fastest shrinking geodesic from §4.
The main goal of this section is to analyse $f_{v}$ assuming $v$ is unstable. We prove Theorem 1.3 Part 1, by proving the next theorem.
Theorem 5.1 (Busemann bounds $f_{v}$ from below).
The tuple $(M,o,f_{v})$ satisfies the assumptions of Definition 4.6. In particular, there exists a fastest shrinking geodesic $\gamma=\gamma_{M,o,f_{v}}:[0,\infty)\to M$ and shrinking rate $a=a_{M,o,f_{v}}>0$ for it. Moreover, there exists $C\in\mathbb{R}$ such that for every $x\in M$, we have
(5.1) $f_{v}(x)\geq a\beta_{\gamma}(x)+C.$
Remark 5.2.
We distinguish our function $f_{v}$ with the pathological in §1.3 using the fact that our function $f_{v}$ decay rate is almost linear, that is, $f(\gamma_{M,\pi(e),f}(s))=-as+O(1)$.
Remark 5.3.
We use norms on $V$ that satisfy Lemma 2.12 to obtain the convexity of $f_{v}$. If we did not assume that the norm satisfies this condition and is only $K$-invariant, (5.1) would still hold with a different constant $C$, as all norms on $V$ are equivalent.
Observation 5.4 (Convexity of $f_{v}$).
The choice of quadratic form in Lemma 2.12 implies that $f_{v}$ is convex. Moreover, for every maximal flat $\pi(Ag)$, $g\in K$, we have
$f_{v}(\pi(ag))=\frac{1}{2}\log\sum_{\lambda\in\Phi_{\varrho}}\lambda(a)^{2}\|(gv)_{\lambda}\|^{2}=\max_{\lambda\in\Phi_{\varrho}}\left(\log\lambda(a)+r_{\lambda}\right)+O(1),$
where $r_{\lambda}=\log\|(gv)_{\lambda}\|\in\mathbb{R}\cup\{-\infty\}$ depends continuously on $g$ and $v$.
Recall that $v\in V$ is called unstable if $0\in\overline{Gv}\setminus Gv$, where $\overline{Gv}$ denotes the Zariski-closure of $Gv$.
Claim 5.5.
If $v$ is unstable, then the function $f_{v}$ decreases linearly (see Definition 4.4).
Remark 5.6.
This claim follows from [21], but we prove it differently to illustrate the techniques in a simpler way.
Proof.
Recall that by Definition 2.4 the metric on $M$ (explicitly defined in §2.3.3) restricted to any flat is Euclidean. Recall that we denoted by $\mathfrak{a}$ the Lie algebra of $A$.
First, note that every maximal flat of $M$ is of form $F_{k}:=\pi(Ak)$, for some $k\in K$. Moreover, the restriction of $f_{v}$ to $F_{k}$ can be approximated by
(5.2) $\tilde{f}_{k}(\mathtt{a})=\max_{\lambda\in\Phi_{\varrho}}\left(\left\langle\mathtt{a}_{\lambda},\mathtt{a}\right\rangle+r_{\lambda}(k)\right),$
where $\mathtt{a}_{\lambda}$ is the unique vector such that $\left\langle\mathtt{a}_{\lambda},\mathtt{a}\right\rangle=\lambda(\mathtt{a})$ and $r_{\lambda}(k)=\log\|(kv)_{\lambda}\|\in\mathbb{R}\cup\{-\infty\}$. By approximated we mean that for $\mathtt{a}\in\mathfrak{a}$ the difference $f_{v}(\exp(\mathtt{a})k)-\tilde{f}_{k}(\mathtt{a})$ is bounded. This approximation holds since $F_{k}$ is a maximal flat, by Observation 5.4, the explicit construction of the metric in §2.3.3, and the definition of a maximal flat (see Definition 2.4). Note also that here $\lambda$ is viewed as an additive character on $\mathfrak{a}$. In particular, if $\tilde{\lambda}$ is the same character viewed as a multiplicative character on $A$ (i.e., as in Observation 5.4), then $\lambda(\mathtt{a})=\log\tilde{\lambda}(\exp(\mathtt{a}))$.
Next, assume that $f_{v}$ is unbounded from below on $F_{k}$ for some $k\in K$. Then, in particular, $\tilde{f}_{k}$ is unbounded from below. Claim 4.9 implies that $\tilde{f}_{k}$ decreases linearly, and hence also $f_{v}$.
Last, we assume that $f_{v}$ is bounded from below on all maximal flats $F_{k}$, $k\in K$ and show that it implies that $v$ is stable, i.e., $0\notin\overline{Gv}\setminus Gv$. It follows from the assumption that $\tilde{f}_{k}$ are bounded from below as well. By Claim 4.9 for all $k\in K$,
$0\in{\rm conv}(\{v_{\lambda}:\lambda\in\Phi_{\varrho},r_{\lambda}(k)\geq 0\}.$
Lemma 4.10 implies that there is a uniform lower bound on $\tilde{f}_{k}$ for all $k\in K$, and implies that $f_{v}$ has a uniform lower bound on $\bigcup_{k\in K}F_{k}$. Using the Cartan decomposition of $G$ (see Definition 2.10), we have $\bigcup_{k\in K}F_{k}=M$, we conclude that $v$ is stable. ∎
Note that Observation 5.4 and Claim 5.5 prove the first part of Theorem 5.1. That is, the tuple $(M,o,f_{v})$ satisfies the assumptions of Definition 4.6. Therefore, set $\gamma$ to be the fastest shrinking geodesic of $(M,o,f_{v})$, and $a>0$ satisfy the conclusion of Proposition 4.5 for $f_{v}$. In view of Observation 5.4, the study of the restriction of shrink-rate functions to maximal flats is reduced to §4.1.
5.1. Lower bound on maximal flat
The first step in the proof of Theorem 5.1 is showing that Equation (5.1) is satisfied on maximal flats (see §2.3 for the definition of a maximal flat).
Proposition 5.7.
There exists a constant $C>0$, which only depends on $v$, so that for every maximal flat $\gamma\subseteq F\subset M$ and for every $x\in F$, inequality (5.1) holds.
Proof.
Let $F:=\pi(Ag)$, $g\in K$, be a maximal flat containing $\gamma$, and $K_{0}\subseteq K$ be the stabilizer of $\gamma$. By Claim 2.6, the group $K_{0}$ acts transitively on the set of maximal flats containing $\gamma$.
As in the proof of Claim 5.5, since $F$ is a maximal flat, the metric constructed on it in §2.3.3 is Euclidean. By Corollary 2.11 we may assume that for some $\mathtt{a}\in\mathfrak{a}$, $\gamma(t)=\pi(\exp(\mathtt{a})g)$. Moreover, by Example 3.2 the restriction of $\beta_{\gamma}$ to $F$ is defined by
(5.3) $\beta_{\gamma}(\pi(\exp(\mathtt{b})g))=-\left\langle\mathtt{a},\mathtt{b}\right\rangle.$
Fix $k_{0}\in K_{0}$. We wish to show that each $p\in F_{k_{0}}:=\pi(Agk_{0})$ satisfies
$f_{v}(p)\geq a\beta_{\gamma}(p)+C.$
Write $p=\exp(\mathtt{b})gk_{0}$ for $\mathtt{b}\in\mathfrak{a}$. Since $K_{0}$ preserves $\beta_{\gamma}$, by (5.3), this is equivalent to showing that
$f_{\varrho(k_{0})v}(p)\geq-\left\langle\mathtt{a},\mathtt{b}\right\rangle+C.$
Since $f_{\varrho(k_{0})v}$ is a shift (by $k_{0}$) of the function $f_{v}$, by Lemma 4.7, the fastest shrinking geodesic of the function $f_{\varrho(k_{0})v}\mid_{F_{k_{0}}}$ is $\gamma_{gk_{0}}(t)=\pi(\exp(\mathtt{a})gk_{0})$ and the shrinking rate of it is $a$. By Observation 5.4 it follows that for any $p=\exp(\mathtt{b})gk_{0}\in F_{k_{0}}$ $f_{\varrho(k_{0})v}(p)-\tilde{f}_{gk_{0}}(\mathtt{b})$ is bounded, where for $k:=gk_{0}$ the function $\tilde{f}_{k}$ is defined as in (5.2). It follows from Lemma 4.8 that the fastest shrinking geodesic of $\tilde{f}_{k}$ is $\gamma_{k}$ and the shrinking rate of it is $a$. The result now follows from Lemma 4.10. ∎
5.2. Proof of the second part of Theorem 5.1
The following proposition is a generalization of the fact that in a rank-$1$ space, for every geodesic ray $\gamma$ and a point $p$ not in the geodesic, there is another geodesic $\gamma^{\prime}$ through $p$ such that $d(\gamma^{\prime}(t),\gamma)\xrightarrow{t\to\infty}0$.
Proposition 5.8.
For every geodesic ray $\gamma:[0,\infty)\to M$ and a point $x\in M$, there exists a maximal flat $\gamma\subseteq F\subseteq M$ and two geodesics $\gamma_{1}:\mathbb{R}\to F$ and $\gamma_{2}:\mathbb{R}\to M$ such that $\gamma_{1}$ is parallel to $\gamma$ in $F$, $\lim_{t\to\infty}d_{M}(\gamma_{1}(t),\gamma_{2}(t))=0$, and $\gamma_{2}(0)=x$.
Proof.
Up to translation of the we may assume that $\gamma$ is defined by
$\gamma(t)=\pi(\exp(t\mathtt{a})),\quad\text{for}\quad t\in\mathbb{R},$
for some $\mathtt{a}\in\mathfrak{p}$. Write $x=\pi(g_{0})\in M$ for some $g_{0}\in G$. By Lemma 2.14, the element $g_{0}$ can be decomposed as $g_{0}=ktu$, $k\in K$, $t\in T_{-\mathtt{a}}$, $u\in U_{-\mathtt{a}}$. Since $U_{-\mathtt{a}}$ is stabilized by $T_{-\mathtt{a}}$, this is equivalent to $g_{0}=kp_{2}p_{1}$, where $k\in K$, $p_{2}\in U_{-\mathtt{a}}$, $p_{1}\in T_{-\mathtt{a}}$. Let $\gamma_{1}$ be the geodesic defined by $t\mapsto\pi(\exp(t\mathtt{a})p_{1})$. Let $\gamma_{2}$ be the geodesic defined by $t\mapsto\pi(\exp(t\mathtt{a})p_{2}p_{1})$. Since $p_{1}=\exp(\mathtt{p}_{1})$ for some $\mathtt{p}_{1}\in\mathfrak{p}$ that commutes with $\exp(\mathtt{a})$, it follows that $\mathtt{p}_{1}$ and $\mathtt{a}$ lie in some maximal abelian $\mathfrak{a}^{\prime}<\mathfrak{p}$. Then, $A^{\prime}=\exp(\mathfrak{a}^{\prime})$ is a Cartan torus such that $\pi(A^{\prime})$ contains $\gamma$ and $\gamma_{1}$. By Theorem 2.5, $F=\pi(A^{\prime})$ is maximal flat containing the geodesic $\gamma_{1}$. It is left to prove the limit,
$d_{M}(\gamma_{1}(t),\gamma_{2}(t))$ $=d_{M}(\pi(\exp(t\mathtt{a})),\pi(\exp(t\mathtt{a})p_{2}))$
$=d_{M}\left(\pi(I),\pi(\exp(t\mathtt{a})p_{2}\exp(-t\mathtt{a}))\right)\xrightarrow{t\to\infty}0.$
The second equality holds because the right $G$ action on $M$ is by isometries, and the limit is true since $p_{2}\in U_{-\mathtt{a}}$
Let $C\in\mathbb{R}$ be the constant that satisfies the conclusion of Proposition 5.7. Fix $x\in M$. We will show that $f_{v}(x_{0})\geq C+a\beta_{\gamma}(x_{0})$.
Let $\gamma_{1},\gamma_{2},F$ be as in Proposition 5.8. Let us study $\beta_{\gamma}$ on $\gamma_{1}$ and $\gamma_{2}$. Let $i=1,2$. Since $\beta_{\gamma}$ is convex, we deduce that $t\mapsto\beta_{\gamma}(\gamma_{i}(t))$ is convex. Since $\beta_{\gamma}$ is $1$-Lipschitz, we deduce that $t\mapsto\beta_{\gamma}(\gamma_{i}(t))$ is $1$-Lipschitz. Since $\beta_{\gamma}$ is Lipschitz-continuous, and $d_{M}(\gamma_{i}(t),\gamma(t))$ id bounded for $t\geq 0$, we deduce that
$\beta_{\gamma}(\gamma_{i}(t))+t=\beta_{\gamma}(\gamma_{i}(t))-\beta_{\gamma}(\gamma(t))$
is bounded for $t\geq 0$.
Observation 5.9.
If $f:\mathbb{R}\to\mathbb{R}$ is a convex, such that $\lim_{t\to\infty}\frac{f(t)}{t}=b$ exists. f(t)tbf(t)-tb is decreasing.
Since $\beta_{\gamma}(\gamma_{i}(t))+t$ is bounded for $t\geq 0$, observation 5.9 implies that $\beta_{\gamma}(\gamma_{i}(t))+t$ is decreasing. Since $\beta_{\gamma}(\gamma_{i}(t))$ is $1$-Lipschitz, we deduce that $\beta_{\gamma}(\gamma_{i}(t))+t$ is constant. Since $d_{M}(\gamma_{1}(t),\gamma_{2}(t))\xrightarrow{t\to\infty}0$, and $\beta_{\gamma}(\gamma_{i}(t))$ is Lipschitz-continuous, we deduce that $\beta_{\gamma}(\gamma_{1}(t))+t=\beta_{\gamma}(\gamma_{2}(t))+t$, and denote this constant by $C^{\prime}$
We turn to study $f_{v}$. By Claim 5.7, we have that
(5.4) $f_{v}(\gamma_{1}(t))\geq a\beta_{\gamma}(\gamma_{1}(t))+C=a(C^{\prime}-t)+C.$
By Proposition 4.5, we have that $\lim_{t\to\infty}f_{v}(\gamma(t))/t=-a$. For $i=1,2$, and $t\geq\infty$, since $d_{M}(\gamma_{i}(t),\gamma(t))$ is bounded and $f_{v}$ is Lipschitz-continuous, we deduce that $\lim_{t\to\infty}f_{v}(\gamma_{i}(t))/t=-a$. Observation 5.9 implies that
$f_{v}(\gamma_{i}(t))+at$
is decreasing. By (5.4), we deduce that $f_{v}(\gamma_{1}(t))+at\geq aC^{\prime}+C$, and in particular, $\lim_{t\to\infty}f_{v}(\gamma_{1}(t))+at$ exists, and is greater than $aC^{\prime}+C$. Since $d_{M}(\gamma_{1}(t),\gamma_{2}(t))\xrightarrow{t\to\infty}0$, we deduce that $\lim_{t\to\infty}f_{v}(\gamma_{2}(t))+at=\lim_{t\to\infty}f_{v}(\gamma_{1}(t))+at$ exists. Since $f_{v}(\gamma_{2}(t))+at$ is decreasing, we deduce that $f_{v}(\gamma_{2}(t))+at\geq aC^{\prime}+C$. This implies that
$f_{v}(\gamma_{2}(t))\geq a(C^{\prime}-t)+C=a\beta_{\gamma}(\gamma_{2}(t))+C,$
Substituting $t=0$, we obtain Eq. (5.1) for $x$.
6. An algebraic interpretation and a result by Kempf
Inspecting our setting from an algebraic point of view, one can use Kempf [21] to obtain Theorem 5.1, which is a more explicit version of Theorem 1.3 Part 1. In this section, we discuss the algebraic analog of the geometric notions and results presented in previous sections. Specifically, we show Theorem 6.7, which is an algebraic analog of Theorem 3.7, and Theorem 6.17, which is an algebraic amplification of the notion of the fastest shrinking geodesic.
6.1. Algebraic analogous of geometric claims
Kempf studied group homomorphisms in $\operatorname{Hom}_{\ell}(\mathbb{G}_{m},G)$, for $\ell\subseteq\mathbb{R}$, which are algebraic analogous of the geometric notion of geodesics. The following claim describes the connection between them:
Claim 6.1 (Connection between geodesics and group homomorphism).
Let $G$ an $\mathbb{R}$-algebraic semisimple group (as in §2.3). For every nontrivial homomorphism $\tau:\mathbb{G}_{m}\to G$ there is a unique $\mathtt{a}\in\mathfrak{p}$ such that the homomorphism $t\mapsto\tau(\exp(t))$ is a conjugate, by an element in $P_{\mathtt{a}}$ to the homomorphism $t\mapsto\exp(t\mathtt{a})$ (see §2.3.3 for the definition of $\mathfrak{p}$ and §2.5 for the definition of $P_{\mathtt{a}}$).
Remark 6.2.
A similar claim could be made for a general homomorphism in $\operatorname{Hom}(\mathbb{R},G)$, but since the adjoint action of the homomorphism may have non-real eigenvalues, the claim and its proof are more elaborate.
Remark 6.3.
In the setting of Claim 6.1, $P_{\mathtt{a}}$ can be defined using $\tau$ by
$P_{\mathtt{a}}=P_{\tau}:=\{g\in G:\tau(t)g\tau(t^{-1})\text{ converges as }t\to 0\}.$
Proof of Claim 6.1.
Let $\tau:\mathbb{G}_{m}\to G$ be a nontrivial homomorphism. First note that by [25, Prop. 2.6], $P_{\tau}$, as defined in Remark 6.3, is parabolic.
Now, $\tau(\mathbb{G}_{m})$ is a torus of $G$. By [3, Thm 15.13] there exists $g\in G$ so that $g\tau(\mathbb{G}_{m})g^{-1}\subseteq A=\exp\mathfrak{a}$ for some Cartan torus $\mathfrak{a}<\mathfrak{p}$. Hence there is $\mathtt{a}^{\prime}\in\mathfrak{a}$ such that $g\tau(\exp(s))g^{-1}=\exp(s\mathtt{a}^{\prime})$. By Lemma 2.14 we obtain that $G=KP_{\tau}$. Decompose $g=kp$ with $k\in K$ and $p\in P_{\tau}$.
Since $\mathfrak{p}$ is invariant under $K$ we deduce that $\mathtt{a}=\operatorname{Ad}_{k}(\mathtt{a}^{\prime})\in\mathfrak{p}$ satisfies $p\tau(\exp(s))p^{-1}=\exp(s\mathtt{a})$ for all $s\in\mathbb{R}$. Since $\tau$ is nontrivial, $\mathtt{a}\neq 0$.
Note that $\gamma:s\mapsto\pi(\exp(s\mathtt{a}/\|\mathtt{a}\|))$ is a geodesic. The definition of $P_{\tau}$ implies that
$d(\gamma(s),\pi(\tau(\exp(s/\|\mathtt{a}\|))))$ $=d(\pi(\exp(t\mathtt{a}/\|\mathtt{a}\|)),\pi(\tau(\exp(s/\|\mathtt{a}\|))))$
$=d(\pi(p\tau(\exp(s/\|\mathtt{a}\|))p^{-1}\tau(\exp(-s/\|\mathtt{a}\|))),\pi(1))$
is bounded as $s\to$-$\infty$. Lemma 2.2 implies that this defines $\gamma$ uniquely. ∎
Definition 6.4.
Let $\tau\in\operatorname{Hom}_{\mathbb{R}}(\mathbb{G}_{m},G)$ be a non-trivial homomorphism and $\mathtt{a}$ be as in Claim 6.1. The geodesic ray $\gamma_{\tau}$ defined by $\gamma_{\tau}(t):=\pi(\exp(t\mathtt{a}/\|\mathtt{a}\|))$ is called the geodesic ray associated to $\tau$ (where $\left\|\cdot\right\|$ is as defined in §2.3.3), and $\|\mathtt{a}\|$ is called the renormalization constant of $\tau$. Remark 6.9 below shows that $\|\mathtt{a}\|^{2}$ is an integer.
Remark 6.5.
Note that not all geodesic rays are associated to a homomorphism. For example, take $G=\operatorname{SL}_{3}(\mathbb{R}),\,K=SO(3)$, and
$\gamma(t)=\pi(\exp(t\operatorname{\mathrm{diag}}(1,\sqrt{2},-1-\sqrt{2}))),$
then $\gamma$ does arise from any algebraic homomorphism in $\operatorname{Hom}_{\mathbb{R}}(\mathbb{G}_{m},G)$.
From here on we assume that $G$ is defined over a fixed field $\ell\subseteq\mathbb{R}$.
Definition 6.6 ($\ell$-algebraic geodesics and Busemann functions).
A geodesic ray $\gamma:[0,\infty)\to M$ is called $\ell$-algebraic if it is associated to an $\ell$-algebraic group homomorphism $\tau\in\operatorname{Hom}_{\ell}(\mathbb{G}_{m},G)$. In this case, let $\beta_{\gamma}$ be the Busemann function defined by $\gamma$ (as in (3.1)) and $a_{\tau}$ be the renormalization constant of $\tau$. For every $q\in\mathbb{Q}^{>0}$ the function $qa_{\tau}\beta_{\gamma}$ is called the $\ell$-renormalized Busemann function.
We can now provide an algebraic analogous of Theorem 3.7.
Theorem 6.7.
The following three classes of functions are equal:
  1. (1)
    Busemann functions: The class of constant shifts of $\ell$-renormalized Busemann functions.
  2. (2)
    Homomorphisms from parabolics: The class of functions which are projections of $\ell$-positive homomorphisms $P\to\mathbb{R}$, where $P$ is an $\ell$-parabolic subgroup of $G$, see Definition 3.5 of $\ell$-positive homomorphism.
  3. (3)
    Lengths of highest weight vectors in fundumental representations: The class of functions of the form
    $\pi(g)\mapsto\sum_{i=1}^{r}c_{i}\log\|\varrho_{i}(gg_{0})v_{i}\|+C,$
    where $\varrho_{i}$ are $\ell$-fundamental representations, $g_{0}\in G(\ell)$, and $v_{i}$ are the appropriate highest weight vectors, which correspond to maximal $\ell$-parabolic groups, and are defined in §2.5 for some choice of an $\ell$-simple system, the $c_{i}$ are non-negative rational numbers, and $C\in\mathbb{R}$.
  4. (4)
    Length of a single parabolic equivariant vector: The class of functions of the form
    $\pi(g)\mapsto\alpha\log\|\varrho(g)v\|,$
    where $\varrho:G\to\operatorname{GL}(V)$ is an $\ell$-algebraic representations with a vector $v\in V(\ell)$ such that the ray $\ell v$ is stabilized by a parabolic subgroup, and $\alpha>0$ is rational.
Proof.
The proof of the equivalence between the first three classes follows in a very similar way to the proof of Theorem 3.7, using the more general theory which is presented in §2.5. The implication of first three classes from Part (4) is similar to Theorem 3.10. We are left to show Part (4) assuming the other three.
Assume a function $f$ is in the third class, i.e.,
$f(\pi(g))=\sum_{i=1}^{r}c_{i}\log\|\varrho_{i}(gg_{0})v_{i}\|+C,$
where $C\in\mathbb{R}$, for any $1\leq i\leq d$, $c_{i}$ is rational and nonnegative, $g_{0}\in G$, and the vector $v_{i}$ is the previously chosen highest weight vector in $V_{i}(\ell)$. In particular, the line $\ell v_{i}$ is $P_{i}$-invariant for an $\ell$-maximal parabolic subgroup $P_{i}$. Set $P=\bigcap_{i}P_{i}$. Then, $P$ is a parabolic subgroup by Claim 2.29. Applying $g_{0}$, the line $\ell\varrho_{i}(g_{0})v_{i}$ is $g_{0}P_{i}g_{0}^{-1}$-invariant, and so $g_{0}Pg_{0}^{-1}$-invariant. Also, up to multiplication by an integer, we may assume that the $c_{i}$’s are integers.
Let
$v^{\prime}:=v_{1}^{\otimes c_{1}}\otimes v_{2}^{\otimes c_{2}}\otimes\cdots\otimes v_{r}^{\otimes c_{r}}\in V_{1}^{\otimes c_{1}}\otimes V_{2}^{\otimes c_{2}}\otimes\cdots\otimes V_{r}^{\otimes c_{r}}$
and $v=\varrho(g_{0})v^{\prime}$, where $\varrho=\varrho_{1}^{\otimes c_{1}}\otimes\varrho_{2}^{\otimes c_{2}}\otimes\cdots\otimes\varrho_{r}^{\otimes c_{r}}$. Then, the line $\ell v$ is $P$-invariant via the representation $\varrho$ and this vector in this representation represents the function as in class (4).
6.2. Kempf’s result
The following definition is a normalized sense of how fast an algebraic torus shrinks a vector.
Definition 6.8 (Shrinking rate of an algebraic torus).
Let $\tau:\mathbb{G}_{m}\to G$ be an algebraic nontrivial homomorphism, $\varrho:G\rightarrow\operatorname{GL}(V)$ be an algebraic representation, and $v\in V$.
$V$ can be decomposed into its eigenspaces with respect to the $\mathbb{G}_{m}$ action $\varrho\circ\tau$
(6.1) $V=\bigoplus_{n\in\mathbb{Z}}V_{n},$
where for any $n\in\mathbb{Z}$
(6.2) $V_{n}:=\left\{w\in V:\forall t\in\mathbb{G}_{m},\,\varrho\circ\tau(t)w=t^{n}w\right\}.$
Let $m(v,\tau)$ be the maximal integer for which $v\in\bigoplus_{n\geq m(v,\tau)}V_{n}$.
Note that for every $n\in\mathbb{N}$ we have
$m(v,\tau^{n})=nm(v,\tau).$
Using $e=\frac{\mathbf{d}}{\mathbf{d}s}\in T_{1}\mathbb{G}_{m}$, we define a “norm” on $\operatorname{Hom}_{\mathbb{Q}}(\mathbb{G}_{m},G)$
(6.3) $\|\tau\|:=\sqrt{B(D_{1}\tau(e),D_{1}\tau(e))}=\sqrt{\sum_{n\in\mathbb{Z}}n^{2}\dim V_{n}},$
where $B=B_{e}$ is the killing form (see §2.3.3) and $D_{1}\tau:T_{1}\mathbb{G}_{m}\to T_{1}G\cong\mathfrak{g}$ is the differential of $\tau$ at $1$.
Remark 6.9.
Note that for $\mathtt{a}\in\mathfrak{p}$ which satisfies the conclusion of Claim 6.1 for $\tau$, we have $\left\|\tau\right\|=\left\|\mathtt{a}\right\|$, where the norm on $\mathfrak{p}$ is as defined in §2.3.3. Moreover, it follows from (6.3) that $\left\|\mathtt{a}\right\|$ is a square root of an integer (see also the discussion in [21, §2]).
Lemma 6.10 ([21, Lem. 3.2](c)).
The quantity $m(v,\tau)$ can be described as follows: Consider the map $f=f_{\tau,\varrho,v}:\mathbb{G}_{m}\to V$ defined by
$f(t)=\varrho\circ\tau(t)v.$
For every functional $\varphi\in V^{*}$ we can consider the function $\varphi\circ f$ and compute its valuation at $t=0$. $m(v,\tau)$ is the minimal such valuation.
Remark 6.11.
The alternative description of $m(v,\tau)$ is the algebraic analogous for the shrinking rate of $\|\varrho(\tau(t))v\|$ as $t\to 0$.
Recall the definition of a unstable vector from §5.
Theorem 6.12 ([21, Thm. 4.2]).
Let $G$ be a semisimple $\ell$-algebraic group, $\varrho:G\to\operatorname{GL}(V)$ an algebraic $G$-representation, and $v\in V$ a unstable vector. Then, there exists $\tau\in\operatorname{Hom}_{\ell}(\mathbb{G}_{m},G)$ which maximize $\frac{m(v,\tau)}{\|\tau\|}$ and for which $m(v,\tau)>0$. This subgroup $\tau$ is unique up to taking power and conjugating by elements in $P_{\tau}$. More explicitly, if $\tau^{\prime}\in\operatorname{Hom}_{\ell}(\mathbb{G}_{m},G)$ are another such homomorphism that cannot be represented as nontrivial power, then there exists a unique element $u\in U_{\tau}$, the unipotent radical of $P_{\tau}$, such that $\tau^{\prime}=u\tau u^{-1}$. We call such $\tau$ Kempf’s homomorphism of $v$.
The following is a corollary of the uniqueness in Theorem 6.12.
Corollary 6.13.
In the setting of Theorem 6.12, suppose that $k/\ell$ is a field extension. Then the extension of scalars of the Kempf’s homomorphism of $v$ to $k$ is again a Kempf’s homomorphism of the extended $v$.
The proof of Corollary 6.13 relies on the degeneracy of a certain Galois action, and so we need the following definitions.
Definition 6.14 (Cocycle and cohomologous cocycles).
Let $\ell$ be a field, $k$ be a finite extension of $\ell$, and $U$ be an $\ell$-algebraic group. Then, there is a Galois action $\operatorname{Gal}(k/\ell)\lefttorightarrow U(k)$. A function $\alpha:\operatorname{Gal}(k/\ell)\to U(k)$ is called cocycle if for every $\sigma_{1},\sigma_{2}\in\operatorname{Gal}(k/\ell)$,
$\alpha(\sigma_{1}\sigma_{2})=\alpha(\sigma_{1})\sigma_{1}(\alpha(\sigma_{2})).$
Two cocycles $\alpha,\beta$ are said to be cohomologous if for some $u\in U$ we have
$\forall\sigma\in\operatorname{Gal}(k/\ell),\quad\beta(\sigma)=u^{-1}\alpha(\sigma)\sigma(u).$
We also need the following known result, which we prove for completion of the manuscript.
Claim 6.15.
Assuming the notation of Definition 6.14, every cocycle $\alpha:\operatorname{Gal}(k/\ell)\to\mathbb{G}_{a}$ is cohomologous to the trivial cocycle.
Proof.
Note that here the group is additive. Let $\alpha:\operatorname{Gal}(k/\ell)\to\mathbb{G}_{a}$. Let $x=\frac{1}{[k:\ell]}\sum_{\sigma\in\operatorname{Gal}(k/\ell)}\alpha(\sigma)$. Then for every $\sigma^{\prime}\in\operatorname{Gal}(k/\ell)$ we have
$x-\sigma^{\prime}(x)$ $=\frac{1}{[k:\ell]}\sum_{\sigma\in\operatorname{Gal}(k/\ell)}(\alpha(\sigma)-\sigma^{\prime}(\alpha(\sigma)))$
$=\frac{1}{[k:\ell]}\sum_{\sigma\in\operatorname{Gal}(k/\ell)}(\alpha(\sigma^{\prime}\sigma)-\sigma^{\prime}(\alpha(\sigma)))$
$=\frac{1}{[k:\ell]}\sum_{\sigma\in\operatorname{Gal}(k/\ell)}(\alpha(\sigma^{\prime}))$
$=\alpha(\sigma^{\prime}).$
The claim follows. ∎
A simple Corollary of Claim 6.15 is the following.
Corollary 6.16.
Suppose we have a pair of $\ell$ algebraic groups $U_{1}\rhd U_{2}$ such that $U_{1}/U_{2}\cong\mathbb{G}_{a}$ and $U_{1}(k)/U_{2}(k)\cong\mathbb{G}_{a}(k)$. Then every cocycle $\alpha:\operatorname{Gal}(k/\ell)\to U_{1}(k)$ is cohomologous to a cocycle with values in $U_{2}(k)$.
Proof.
The composition of $\alpha$ and the projection $U_{1}(k)\to(U_{1}/U_{2})(k)\cong\mathbb{G}_{a}(k)$ is cohomologous to the trivial cocycle by Claim 6.15. Denote by $u\in(U_{1}/U_{2})(k)$ the element such that $u^{-1}[\alpha(\sigma)]_{(U_{1}/U_{2})(k)}\sigma(u)=1$. Let $\tilde{u}\in U_{1}(k)$ be an element that projects to $u$. Then $\tilde{u}$ shows that $\alpha$ is cohomologous to a cocycle with values in $U_{2}(k)$. ∎
Proof of Corollary 6.13.
Let us assume that there is a counterexample to the claim, i.e., some $\tau_{1}\in\operatorname{Hom}_{k}(\mathbb{G}_{m},G)$ such that
$\frac{m(v,\tau_{1})}{\|\tau_{1}\|}>\frac{m(v,\tau)}{\|\tau\|}.$
Then $\tau_{1}$ is defined over a finitely generated algebra $k_{1}/\ell$. In particular, $k_{1}$ has a quotient $k_{2}$ that is a finite extension of $\ell$ and the projection of scalars $\tau_{2}$ of $\tau_{1}$ to $Hom_{k_{2}}(\mathbb{G}_{m},G)$ also has
$\frac{m(v,\tau_{2})}{\|\tau_{2}\|}>\frac{m(v,\tau)}{\|\tau\|}.$
Hence we may assume without loss of generality that $k$ is a finite field extension of $\ell$. We may replace $k$ by its Galois closure $f$ over $\ell$, applying the argument twice for $f/\ell$ and $f/k$. Thus, we may assume that $k$ is Galois over $\ell$.
For every element $\sigma\in\operatorname{Gal}(k/\ell)$, the two homomorphism $\tau_{1},\sigma(\tau_{1})$ are both Kempf’s homomorphism of $v$. The uniqueness of the parabolic in Theorem 6.12 asserts that $P_{\tau_{1}}=P_{\sigma(\tau_{1})}$. In particular, the parabolic subgroup $P_{\tau_{1}}$ is defined over $\ell$. In addition, this uniqueness also shows that for every $\sigma$ there is a unique $u_{\sigma}\in U_{\tau_{1}}(k)$ such that $u_{\sigma}^{-1}\tau_{1}u_{\sigma}=\sigma(\tau_{1})$. Applying the last for $\sigma_{1},\sigma_{2}\in\operatorname{Gal}(k/\ell)$ we obtain
$u_{\sigma_{1}\sigma_{2}}^{-1}\tau_{1}u_{\sigma_{1}\sigma_{2}}$ $=\sigma_{1}\sigma_{2}(\tau_{1})=\sigma_{1}(u_{\sigma_{2}}^{-1}\tau_{1}u_{\sigma_{2}})$
$=\sigma_{1}(u_{\sigma_{2}})^{-1}\sigma_{1}(\tau_{1})\sigma_{1}(u_{\sigma_{2}})$
$=(u_{\sigma_{1}}\sigma_{1}(u_{\sigma_{2}}))^{-1}\tau_{1}u_{\sigma_{1}}\sigma_{1}(u_{\sigma_{2}}).$
Thus, the uniqueness of $u_{\sigma_{1}\sigma_{2}}$ implies $u_{\sigma_{1}\sigma_{2}}=u_{\sigma_{1}}\sigma_{1}(u_{\sigma_{2}})$. Hence,
$u_{\bullet}:\operatorname{Gal}(k/\ell)\to U_{\tau_{1}}(k)$
is a cocycle.
Next, we claim that if $u_{\bullet}$ is cohomologous to the trivial cocycle, that is, that there is $u\in U_{\tau_{1}}(k)$ such that $u_{\sigma}=u^{-1}\sigma(u)$, then we done. Indeed, this imply that $u\tau_{1}u^{-1}$ is invariant under the Galois action,
$\sigma(u\tau_{1}u^{-1})$ $=\sigma(u)u_{\sigma}^{-1}\tau_{1}u_{\sigma}\sigma(u)^{-1}$
$=\sigma(u)(u^{-1}\sigma(u))^{-1}\tau_{1}u^{-1}\sigma(u)\sigma(u)^{-1}=u\tau_{1}u^{-1},$
and hence $u\tau_{1}u^{-1}$ is defined over $\ell$, a contradiction to the maximality of $\tau$.
Let us now show that $u_{\bullet}$ is cohomologous to the trivial cocycle. Since $U_{\tau_{1}}$ is unipotent, it has an $\ell$-algebraic composition series $U_{\tau_{1}}=U_{0}\rhd U_{1}\rhd\dots\rhd U_{s}$ such that $U_{i}/U_{i+1}\cong\mathbb{G}_{a}$ and $U_{i}(k)/U_{i+1}(k)\cong\mathbb{G}_{a}(k)$. Then, iteratively replace the cocycle $u_{\bullet}$ with cohomologous cocycles with values in $U_{i}$ using Corollary 6.16, and finally show that $u_{\bullet}$ is cohomologous to the trivial cocycle. ∎
Recall the definition of the shrink-rate function of a vector from §5, as well as the definitions of the fastest shrinking geodesic of a function and the shrinking rate of it from Definition 4.6.
The following claim identifies the fastest shrinking geodesic of the shrink-rate function of a vector and its Kempf’s homomorphism.
Lemma 6.17.
Let $\varrho:G\to\operatorname{GL}(V)$ be an $\ell$-algebraic representation, and $v\in V(\ell)$ be an unstable vector. Then, the fastest shrinking geodesic of $f_{v}$ (as defined in §5) is of the form $\gamma_{\tau}$ where $\tau\in\operatorname{Hom}_{\ell}(\mathbb{G}_{m},G)$ is Kempf’s homomorphism of $v$. In addition, the shrinking rate satisfies $a_{f_{v},M,o}=\frac{m(v,\tau)}{\|\tau\|}$.
Proof.
Let $\tau\in\operatorname{Hom}_{\ell}(\mathbb{G}_{m},G)$ be Kempf’s homomorphism of $v$. By Corollary 6.13, the extension of scalars of $\tau$ to $\mathbb{R}$ is also Kempf’s homomorphism of $v$. Thus extension of scalars to $\mathbb{R}$ does not change the claim, and we may assume that $\ell=\mathbb{R}$.
Let $\varsigma\in\operatorname{Hom}_{\mathbb{R}}(\mathbb{G}_{m},G)$ be nontrivial. Then, by Lemma 6.10 and the definitions of $f_{v}$, we have
$\frac{m(v,\varsigma)}{\|\varsigma\|}=\lim_{s\to\infty}\frac{-f_{v}(\gamma_{\varsigma}(s))}{s}.$
Using Claim 6.1, this implies that $\tau$ is the algebraic geodesic rays that shrinks $f_{v}$ the fastest. In particular, $a_{f_{v},M,o}\geq\frac{m(v,\varsigma)}{\|\varsigma\|}$ for all $\varsigma\in\operatorname{Hom}(\mathbb{G}_{m},G)$ and equality occurs if and only if the fastest shrinking geodesic is algebraic.
Without loss of generality, assume that the fastest shrinking geodesic is $\gamma(s)=\exp(s\mathtt{a})$ for some $\mathtt{a}\in\mathfrak{a}$. Note that the set of algebraic geodesic rays is dense in the set of all rays, and that by Observation 5.4 and Claim 4.12, the function
$\gamma\mapsto\lim_{s\to\infty}\frac{-f_{v}(\gamma(s))}{s},$
defined for all geodesic rays $\gamma$ in $\pi(A)$, is continuous. It follows that $a_{f_{v},M,o}$ is at most the supremum of $\lim_{s\to\infty}\frac{-f_{v}(\gamma(s))}{s}$ over all algebraic geodesic rays $\gamma$ from $o$ in $\pi(A)$. This concludes the proof. ∎
7. Alternative proof of Theorem 1.1
Let $\varrho:G\to\operatorname{GL}(V)$ be a $\mathbb{Q}$-representation, $V$ be equipped with a $K$-invariant norm, and $v$ be an unstable non-zero vector in $V(\mathbb{Q})$.
Define the shrink-rate function $f_{v}:G\to\mathbb{R}$ of $v$ as in §5, i.e.,
$f_{v}(\pi(g))=\log\|gv\|.$
Note that the function $f_{v}$ is the quantity on the left hand side of (1.1), which we wish to control. By Theorem 5.1 there exists a fastest shrinking geodesic for $f_{v}$ (see §4 for the definition of a fastest shrinking geodesic), denote it by $\gamma$, and $C\in\mathbb{R}$ such that for every $x\in M:=K\backslash G$, we have
(7.1) $f_{v}(x)\geq\tilde{a}\beta_{\gamma}(x)+C,$
where $\beta_{\gamma}$ is the Busemann function associated to $\gamma$ (defined in §3) and $\tilde{a}$ is the shrinking rate of $f_{v}$ on $\gamma$, that is, $\tilde{a}=-\lim_{t\to\infty}\frac{f_{v}(\gamma(t))}{t}$. Moreover, since $v\in V(\mathbb{Q})$, by Lemma 6.17 $\gamma$ is $\mathbb{Q}$-algebraic, and hence $\tilde{a}\beta_{\gamma}$ is a $\mathbb{Q}$-renormalized Busemann function (see Definition 6.6).
Now, Theorem 6.7 connects the above Busemann function to a $\mathbb{Q}$-highest weight representation, $\varrho^{\prime}:G\rightarrow\operatorname{GL}(W)$ implying that there exist $w\in W(\mathbb{Q})$ stabilized by a parabolic subgroup and non-negative $\tilde{a}$ and $\tilde{C}$ such that for any $g\in G$
(7.2) $\beta_{\gamma}(\pi(g))=\tilde{\alpha}\log\|\varrho^{\prime}(g)w\|-\tilde{C}$.
Note that $w$ is the highest weight vector with respect to some maximal $\ell$-split torus and a choice of a simple system by Lemma 2.23.
Combining (7.1) and (7.2), one may deduce that for some non-negative $a,c$ and all $g\in G$, we have
$f_{v}(\pi(g))\geq a\log\|\varrho^{\prime}(g)w\|-c.$
The claim now follows from the definition of $f_{v}$.
Appendix A Proof of Lemma 2.12
Here we prove Lemma 2.12. We restate it below for the convenience of the reader.
Lemma A.1 (Construction of bilinear form).
Let $\varrho:G\rightarrow\operatorname{GL}(V)$ be an $\mathbb{R}$-representation. Then, there is a $K$-invariant positive bilinear form $\left\langle\cdot,\cdot\right\rangle$ on $V$ so that the linear spaces $V_{\lambda}$ are orthogonal with respect to it (see §2.4 for the definition of $V_{\lambda}$).
For the proof we need some definitions regarding Galois group actions.
Definition A.2 (Galois Action).
Denote by $\operatorname{Gal}(\mathbb{C}/\mathbb{R})$ the Galois group, i.e., the group of automorphisms of $\mathbb{C}$ which fix $\mathbb{R}$ pointwise. Then, the only elements in $\operatorname{Gal}(\mathbb{C}/\mathbb{R})$ are the identity map and the complex conjugation, which we denote by $\operatorname{{\mathrm{conj}}}$. Moreover, there is an action of $\operatorname{Gal}(\mathbb{C}/\mathbb{R})$ on:
  1. (i)
    complexifications of $\mathbb{R}$-algebraic objects, such as $G^{\mathbb{C}}$ and $V\otimes\mathbb{C}$, in a natural way.
  2. (ii)
    the category of complex vector spaces in the following way: Given a vector space $U$, let $U^{\operatorname{{\mathrm{conj}}}}$ be a vector space with the same set of points as $U$, let the addition on $U^{\operatorname{{\mathrm{conj}}}}$ be the same as on $U$, and the multiplication by a scalar $c\in\mathbb{C}$ on $U^{\operatorname{{\mathrm{conj}}}}$ be the multiplication by the conjugate of $c$ on $U$. We denote the topological map from $U$ to $U^{\operatorname{{\mathrm{conj}}}}$ by $\operatorname{{\mathrm{conj}}}$.
  3. (iii)
    the category of complex representations of $G^{\mathbb{C}}$ as follows: For $\sigma:G^{\mathbb{C}}\to\operatorname{GL}(U)$ we can set a (canonical) representation $\sigma^{\operatorname{{\mathrm{conj}}}}:G^{\mathbb{C}}\to\operatorname{GL}(U^{\operatorname{{\mathrm{conj}}}})$ which acts by
    $\sigma^{\operatorname{{\mathrm{conj}}}}(g)(u)=\operatorname{{\mathrm{conj}}}(\sigma({\operatorname{{\mathrm{conj}}}}(g))({\operatorname{{\mathrm{conj}}}(u)})).$
  4. (iv)
    $G$ actions have the following phenomenon: Assume $G$ is a group acting on a space $X$ via $\sigma:G\to\operatorname{Aut}(X)$. An action $\operatorname{Gal}(\mathbb{C}/\mathbb{R})\lefttorightarrow\varrho$ is a tuple of actions $\operatorname{Gal}(\mathbb{C}/\mathbb{R})\lefttorightarrow G,\operatorname{Gal}(\mathbb{C}/\mathbb{R})\lefttorightarrow X$ such that
    $\operatorname{{\mathrm{conj}}}(\sigma(g)(x))=\sigma(\operatorname{{\mathrm{conj}}}(g))(\operatorname{{\mathrm{conj}}}(x)),$
    for every $g\in G,x\in X$. The above is equivalent to an extension of $\sigma$ to an action of the semidirect product $\tilde{\sigma}:\operatorname{Gal}(\mathbb{C}/\mathbb{R})\ltimes G\to\operatorname{Aut}(X)$.
The main tool we use in the proof is the unitarian trick:
Theorem A.3 (Unitarian Trick, [37, Thm. 4.11.14]).
Let $H$ be a semisimple simply-connected complex Lie group with a maximal compact subgroup $K_{H}$. The restriction of $H$-representations to $K_{H}$-representations induces a bijection from the category of complex algebraic representations of $H$ to the category of complex finite dimensional representations of $K_{H}$.
If $G$ is not algebraically simply-connected, replace it with its algebraic simply-connected cover. This preserves the conclusion of Lemma A.1.
To prove Lemma A.1, we study real representations of a real $G$ by studying the complex algebraic represenations of the complexification $G^{\mathbb{C}}$. Then, we will Unitatian trick to relate the discussion to study of complex representation of a maxiaml compact group $K^{\mathbb{C}}\subset G^{\mathbb{C}}$. The compactness of $K^{\mathbb{C}}$ will make it easy to find the desired positive definite quadratic form.
Proof of Lemma A.1.
Let $G$ be a real algebraic semisimple Lie group, and $\varrho:G\to\operatorname{GL}(V)$ a representation. Since $G$ is an algebraic group, we can consider its complexification $G^{\mathbb{C}}$, with the conjugation action $\operatorname{{\mathrm{conj}}}_{G}:G^{\mathbb{C}}\to G^{\mathbb{C}}$.
Claim A.4.
There is a maximal compact subgroup $K^{\mathbb{C}}\subset G^{\mathbb{C}}$ such that $\operatorname{{\mathrm{conj}}}_{G}(K^{\mathbb{C}})=K^{\mathbb{C}}$, and $K^{\mathbb{C}}\cap G=K$.
Proof.
To find an explicit description of $K^{\mathbb{C}}$, the maximal compact subgroup of $G^{\mathbb{C}}$, we use standard constructions, see [22, §VI,2]. Recall that $\theta:\mathfrak{g}\to\mathfrak{g}$ is the Cartan involution antihomomorphism, it acts on the Lie algebra $\mathfrak{g}=\mathfrak{k}\oplus\mathfrak{p}$ by inverting $\mathfrak{k}$ and preserving $\mathfrak{p}$. Since $\mathfrak{g}^{\mathbb{C}}=\mathfrak{g}\otimes\mathbb{C}$, we may extend $\theta$ to a real involution $\theta^{\mathbb{C}}:\mathfrak{g}^{\mathbb{C}}\to\mathfrak{g}^{\mathbb{C}}$ by defining
$\theta^{\mathbb{C}}=\theta\otimes\operatorname{{\mathrm{conj}}}.$
Since $\mathfrak{g}\otimes\mathbb{C}=\mathfrak{k}\otimes\mathbb{C}\oplus\mathfrak{p}\otimes\mathbb{C}$, the above defines a decomposition $\mathfrak{g}^{\mathbb{C}}=\mathfrak{k}^{\mathbb{C}}\oplus\mathfrak{p}^{\mathbb{C}}$, viewed as real vector spaces, for
$\mathfrak{k}^{\mathbb{C}}:=$ $\mathfrak{k}\oplus i\mathfrak{p},$
$\mathfrak{p}^{\mathbb{C}}:=$ $i\mathfrak{k}\oplus\mathfrak{p}.$
Moreover, $\mathfrak{k}^{\mathbb{C}}$ is the $1$ eigenspace of $\theta^{\mathbb{C}}$ and $\mathfrak{p}^{\mathbb{C}}$ is the 1-1 eigenspace.
Since $\theta^{\mathbb{C}}$ is a Lie algebra homomorphism, we deduce that its fixed points, i.e., $\mathfrak{k}^{\mathbb{C}}$, is also a Lie algebra. Direct computation shows that the killing form of the group $G^{\mathbb{C}}$, viewed as a real algebraic group, is positive on $\mathfrak{p}^{\mathbb{C}}$ and negative on $\mathfrak{k}^{\mathbb{C}}$. This implies that this is indeed the cartan decomposition of $G^{\mathbb{C}}$, thought of as a real Lie group. In particular, $K^{\mathbb{C}}:=\exp\mathfrak{k}^{\mathbb{C}}$ is a maximal compact subgroup of $G^{\mathbb{C}}$. Since $\mathfrak{k}^{\mathbb{C}}=\mathfrak{k}\oplus i\mathfrak{p}$ is invariant to complex conjugations, we deudce tht $\operatorname{{\mathrm{conj}}}_{G}(K^{\mathbb{C}})=K^{\mathbb{C}}$, and since $\mathfrak{k}\subseteq\mathfrak{k}^{\mathbb{C}}$ we deudce that $K\subseteq K^{\mathbb{C}}$. ∎
We assume without loss of generality that $\varrho$ is irreducible. Then, one of the following holds:
  1. (1)
    $\varrho\otimes\mathbb{C}$ is an irreducible representation of $G^{\mathbb{C}}$.
  2. (2)
    There is a decomposition $\varrho\otimes\mathbb{C}\cong\sigma\oplus\sigma^{\operatorname{{\mathrm{conj}}}}$.
We will first prove the lemma for $\varrho$ as in the case (1). In that case, $\varrho|_{K^{\mathbb{C}}}$ is also irreducible by Theorem A.3.
Claim A.5.
On every irreducible complex $K^{\mathbb{C}}$-representation there is exactly one $K^{\mathbb{C}}$ invariant, positive definite Hermitian form up to multiplication by a positive scalar.
Proof.
Let $\sigma:K^{\mathbb{C}}\to\operatorname{GL}(U)$ be an irreducible representation. There is at least one such invariant positive definite Hermitian form, as we can average a non-invariant positive definite Hermitian form along the $K^{\mathbb{C}}$ action.
If there are two such positive definite Hermitian forms, $H_{1},H_{2}$, consider the infimum
$\alpha:=\inf\{\alpha^{\prime}>0:H_{1}-\alpha^{\prime}H_{2}\text{ is not positive definite}\}.$
Then, the Hermitian form $H_{1}-\alpha H_{2}$ is non-negative definite and $K^{\mathbb{C}}$-invariant. Moreover, the set
$W=\{w\in U:(H_{1}-\alpha H_{2})(w,w)=0\}$
is nonempty, not equal to $U$, and is $K^{\mathbb{C}}$-invariant. This contradicts the irreducibility of $\sigma$ as a $K^{\mathbb{C}}$ representation. ∎
Denote the unique invariant, positive definite Hermitian form on $V^{\mathbb{C}}=V\otimes\mathbb{C}$ by $H^{\mathbb{C}}$. Since the the Galois group $\operatorname{Gal}(\mathbb{C}/\mathbb{R})$ acts on the $K^{\mathbb{C}}$, $V^{\mathbb{C}}$, and the $K^{\mathbb{C}}$-action $\varrho\otimes\mathbb{C}$ (i.e., satisfies property (iv) of Definition A.2), we deduce that $H^{\mathbb{C}}$ is Galois invariant. Hence, $H^{\mathbb{C}}$ is induced from a positive definite symmetric billinear form $H$ on $V$.
We are left to show that $H$ satisfies the desired property, that is, the linear spaces $V_{\lambda}$ are orthogonal with respect to it. Assume that $V\cong\mathbb{R}^{n}$, $V\otimes\mathbb{C}\cong\mathbb{C}^{n}$, and $H^{\mathbb{C}}$ is the standard Hermitian form on $\mathbb{C}^{n}$. Then, it is enough to show that $A$ is sent by $\varrho^{\mathbb{C}}$ to a group of Hermitian matrices.
As in Definition A.2, the representation $\varrho$ defines a map $\varrho^{\mathbb{C}}:G^{\mathbb{C}}\to\operatorname{GL}_{n}(\mathbb{C})$. Define $\theta_{n}:\operatorname{GL}_{n}(\mathbb{C})\to\operatorname{GL}_{n}(\mathbb{C})$ by
$\theta_{n}(M)=\operatorname{{\mathrm{conj}}}(M^{-t}).$
Since $\varrho^{\mathbb{C}}(A)$ is invariant under $\theta^{\mathbb{C}}$, in order to show that $\varrho^{\mathbb{C}}(A)$ is also invariant under $\theta_{n}$, it suffices to prove that $\theta_{n}\circ\varrho^{\mathbb{C}}\circ\theta^{\mathbb{C}}=\varrho^{\mathbb{C}}$. Since $K^{\mathbb{C}}$ preserves the Hermitian form $H^{\mathbb{C}}$, it follows that $\varrho^{\mathbb{C}}(K_{\mathbb{C}})\subseteq U(n)$. In particular, $\theta_{n}\circ\varrho^{\mathbb{C}}\circ\theta^{\mathbb{C}}|_{K^{\mathbb{C}}}=\rho|_{K^{\mathbb{C}}}$. Since $K^{\mathbb{C}}$ is Zariski dense in $G^{\mathbb{C}}$, it follows that $\theta_{n}\circ\varrho^{\mathbb{C}}\circ\theta^{\mathbb{C}}=\varrho^{\mathbb{C}}$, as desired.
Now, assume we are in case (2), i.e., $\varrho\otimes\mathbb{C}\cong\sigma\oplus\sigma^{\operatorname{{\mathrm{conj}}}}$, for some $\sigma:G^{\mathbb{C}}\to\operatorname{GL}(U)$. The $\mathbb{R}$-linear map $\operatorname{{\mathrm{conj}}}$ on $V\otimes\mathbb{C}\cong U\oplus U^{\operatorname{{\mathrm{conj}}}}$ defines an action on the space $\operatorname{Pos}(V\otimes\mathbb{C})$ of positive definite Hermitian forms $H$ on $V\otimes\mathbb{C}$ by
$\operatorname{{\mathrm{conj}}}(H)(v_{1},v_{2})=\operatorname{{\mathrm{conj}}}(H(\operatorname{{\mathrm{conj}}}(v_{1}),\operatorname{{\mathrm{conj}}}(v_{2}))).$
The group $G^{\mathbb{C}}$ acts on $\operatorname{Pos}(V\otimes\mathbb{C})$ as well, and the $K^{\mathbb{C}}$ invariant Hermitian forms
$\operatorname{Pos}(V\otimes\mathbb{C})^{K^{\mathbb{C}}}=\operatorname{Pos}(U)^{K^{\mathbb{C}}}\oplus\operatorname{Pos}(U^{\operatorname{{\mathrm{conj}}}})^{K^{\mathbb{C}}}.$
Each of these components is isomorphic to $\mathbb{R}^{>0}$ by Claim A.5. Note that $\operatorname{Gal}(\mathbb{C}/\mathbb{R})$ acts on the action $K^{\mathbb{C}}\lefttorightarrow\operatorname{Pos}(V\otimes\mathbb{C})$, and when restricting this action to $\operatorname{Pos}(V\otimes\mathbb{C})^{K^{\mathbb{C}}}$ we get that $\operatorname{{\mathrm{conj}}}$ replaces the two isomorphic components. Adding the conjugation to the acting group, we get a semidirect product, $\operatorname{Gal}(\mathbb{C}/\mathbb{R})\ltimes K^{\mathbb{C}}\lefttorightarrow\operatorname{Pos}(V\otimes\mathbb{C})$. The space of invariants of this action is $\mathbb{R}^{>0}$, and hence up to multiplication by positive scalar there is again only one such positive quadratic form. Since $\operatorname{{\mathrm{conj}}}$ preserves $K^{\mathbb{C}}$ in $G^{\mathbb{C}}$ we get an action of the compact group $\operatorname{Gal}(\mathbb{C}/\mathbb{R})\ltimes K^{\mathbb{C}}$ on $V^{\mathbb{C}}$. The rest of the proof is similar to case (1). ∎
The next example shows that case (2) in the proof of Lemma A.1 can occur.
Example A.6.
Let $G=\operatorname{SU}(3)$, $V=\mathbb{C}^{3}$, and $\varrho:G\rightarrow\operatorname{GL}(V)$ be the standard inclusion representation. Since $G$ acts transitively on the unit sphere in $V$, it is irreducible. Moreover, restriction of scalars $\varrho^{\prime}=\operatorname{Res}_{\mathbb{C}/\mathbb{R}}(\varrho)\in{\rm Rep}_{\mathbb{R}}(G)$ is also irreducible. However, the complexification of $\varrho^{\prime}$ is reducible, as $\varrho^{\prime}\otimes_{\mathbb{R}}\mathbb{C}=\varrho\otimes_{\mathbb{R}}\mathbb{C}$ is of complex dimension $6$ and has the $\mathbb{C}$-linear multiplication map to $\varrho=\varrho\otimes_{\mathbb{C}}\mathbb{C}$.
Funding statement
The first autor is supported by ERC grant HomDyn, ID 833423. The second author did not receive support from any organization for the submitted work.